A Complete Bibliography of Publications in Annals of Probability (1980–1989)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 March 2019 Version 1.02

Title word cross-reference

2 [Hei84]. 2^{−n} [Mit81, Mit82, Mit83b]. 2M − X [Rog81]. α [Ber89b, SS89]. B [De 81, KP80, LW89, Phi80, Phi86, RR80, dA82]. B_{p} [JM83]. d̄ [Ell80a]. β

[Nol88, Nol92]. C([0, 1]; Y′) [Mit83a]. C^{2} [Zei89]. C^{∞} [Rei86]. C(\{\frac{S_{n}}{a_{n}} \})

[dAK83]. d [Ald83, Gri83b, Gri83c]. D(A) [BP85]. D([0, 1]; Y′) [Mit83a]. E max_{1≤k≤n} S_{k}^{+}/E_{n}^{+} [Kla89]. E|S_{n}| [Kla80]. f [Aar86]. I [Csi84, Dyk85, Pin85c, Pin85b]. K [BP83b, Hal85b, Pol82]. l^{2} [IM89]. L^{2}(0, 1) [Mas84]. l^{∞} [FT80b]. L^{p} [Bic81, Bur87b, Phi80, Phi86]. L_{1} [JK83]. L_{2} [Lig89, Oss87]. L_{p} [Bro80a, CH88]. L log^{+} L [FI87]. M [Jan83b, JS87, Jan84]. M/GI/1 [BM89a]. R^{2} [Ros88]. R^{d} [Ney83, Uch82b]. R^{n} [BC83b]. Z [BG80, GG82]. Z^{d} [Cox89]. L_{∞} [Tak81]. N [BP83b]. {n^{−1/α}S_{n}} [Mij82]. N_{0n}/n [ZZ84]. \overline{d} [Ell80b]. p [CM80, CRW85, GH85, KS88, RW86, Tal88a]. Φ [HKS87, Pel85]. R^{1} [MS80a]. R^{2} [HM81]. ρ [Bra88, Pel87]. s [Jur85]. S_{n}/a_{n} [Kue81]. σ [LR86a]. t [Hal87]. t^{1/2} [BC83a]. θ_{p} [CRW85]. U [JJ86, NP88]. X [Kue85]. Z [Arr83a, DG83]. Z^{d} [Arr81]. Z_{d} [Arr83b].

-Bound [Tak81]. -Bounded [FI87]. -Cube [Ald83]. -Dependent [Jan84, Jan83b]. -Dimensional [Gri83b, Gri83c, Hal85b]. -Distance [Ell80b]. -Estimators [JS87]. -Expansions [Aar86]. -Fields [LR86a]. -Function [Pin85c, Pin85b]. -Inequality [Bur87b]. -Invariance [Phi86, Phi80]. -Means [Pol82]. -Mixing [Bra88, Pel85, Pel87, HKS87]. -Processes [NP88]. -Projection [Csi84]. -Projections [Dyk85]. -Radonifying [CRW85]. -Selfdecomposable [Jur85]. -Spaces [RR80]. -Stable [CRW85, GH85, KS88, RW86, Tal88a]. -Statistics [JJ86]. -Theory [Bic81]. -Uniformly [Hei84]. -Valued [IM89, Phi86, De 81, KP80, LW89, Phi80, dA82]. -Variation [Ber89b].

1 [Ros83].

89m [BL97].

A. [Ano89i]. Abel [Che80]. Above [CCN87]. Absence [Sch87a]. Absolute [Bar81a, Bra85, BS86, CK83b, Ric82, She82a]. Absorption [Ber89a]. Abstract [Dun80]. Access [KM87]. Accompanying [AGMZ81]. Achieve [Kie80a]. Acknowledgment [Kur84, Mit83b]. Acquiring [Whi81]. Activities [Shi89]. Adapted [Hoo87]. Addendum [Isc89]. Additif [Der83]. Additive [Ast81, Edg82, NN87a, NN87b, Ust84, Yam85]. Admit [BS87a]. Advantage [HP83]. Affine [HK81]. Age [GH81, Kuc82]. Age-Dependent [GH81]. Aging [BG84, Dal88]. Alan [Don87b]. Alarm [Lin80]. Algebraic [AGKdV89, Far86, Far87]. Algorithm [Mia88, Pit87]. Algorithmic [CGG89]. Algorithms [Ald83, Dup88, SW84]. Allocation [Kar84]. Almost [BY88b, BD88, DP82, Fol84, Gor88, Heb80, Hil82, KP80, Muk84, Rup82, Stu86, TC83, Der83]. almost-subadditive [Der83]. Alphabet [Dun80]. Alternate [Cox84]. Amarts [Edg82, Sch89]. Among [KO88a]. Analysis [CCN87, Dup88, GHR84, Pit87, Wat87]. Analytic [McC85]. Analytical [Res85]. Ann [BL97]. Ann. [Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano82g, Ano82h, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano84g, Ano84h, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano88h, Ano89a, Ano89b, Ano89c, Ano89d]. Ann. [Ano89e, Ano89f, Ano89g, Ano89h]. Annihilating [Arr81, Arr83b]. Annular [Yan86]. Answers [GFZ86]. Anticipative [Ber88b]. Appell [AT87]. Application [ARS87, ER82, ER83, Ell88b, GW82, Hu88, Hug85, KS84, Kon82, LR84, Lin80, Mic82, Mit83c, Rei81, Str86b]. Applications [AS87, BC86, Bro83a, Bro83b, De 81, De 80, Eri81, Gae88, GS80, Glo81a, GKZ81, Gra86, HR83b, JJS88, JJ86, Jan88, Kip86, LW82, Wat87, dA82, dA83b]. Approach [Bru84, BS87b, Coh85, Cra87, DP86, Li80, MN88, Oel84, Res85, Zee81]. Approaches [PT88, Str86b]. Approximate [GJP84, Kon82, Sie88]. Approximated [CCFR89]. Approximating [Bro83a, Kla81, Bro83b]. Approximation [Ari88, CS81, Dal88, DP86, Eng81, Hor84, Kla80, LR86b, Mas89b, NS82, Rei81, Ste86b]. Approximations [AGG89, BR89, Bro83c, BG84, Hal83a, Hal88b, Jen85, Mar84, Pro85a, QR84, Roo80, Rup82, Ser86, Ser88]. Arbitrary [CS85a]. Arcs [Hol81a]. Area [CS87]. Arising [AL82, BPT88, CDG87, LL86, MS89, Rei82a]. Arm [Whi81]. Arm-Acquiring [Whi81]. Armed [Man87]. Arrangement [BPT88]. Arrays [Fal84, HK81, LW82, RT81, Sam84, Ser82, dA80, dA82]. Associated [Bir88b, Bir88a, BW85, CG84, Fer80, Fla82, GG86a, Her84a, Oka85a, Pit82, Woo83, Woo85, Yan88]. Association [BW86, JDPP83]. Assumptions [Ebe86]. Asymmetric [BF87, PT83]. Asymptotic [AC88a, Bar87, Ber82a, BZ81, Car86, Chi82, Cla87, Cre80, CM86, CHM88, DPSW86, Gaw88, GGL88, GP89, GP83b, HKS87, Hij84, JP84, JS87, KKR87, Kin80, Kus85, Mal88, Mat82, MOT85, MS82b, Nai81, PY86a, PY89, Pyk80, Ros84b, Ryz86, SS89, Smi82, Swe80, Tak81, WK87]. Asymptotical [Pit85]. Asymptotically [DR84, GK80]. Asymptotics [BL88, Ney83, BL97]. Atom [MS80a]. Attaining [Ell80a]. Attraction [GH85, GJP84, HK80, HMV83, JO80, Kue85, Mal80, MO83, Pic86, Swe85]. Attractive [GG82, Gra86, Lig83a, Lig89]. Autoregressions [Fin82a]. Autoregressive [Che81a, Fin82b]. Average [BD88, BS80b, Far86, Far87, Roo86, Sol84]. Averages [DR85, GF87]. Averaging [Pin89]. Avoiding [Bar81b, Fre81, Law89, Sla89].

back [Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano82a, Ano82b, Ano82c, Ano82d, Ano83a, Ano83b, Ano83c, Ano83d, Ano84a, Ano84b, Ano84c, Ano84d, Ano85a, Ano85b, Ano85c, Ano85d, Ano86a, Ano86b, Ano86c, Ano86d, Ano87a, Ano87b, Ano87c, Ano87d, Ano88a, Ano88b, Ano88c, Ano88d, Ano89a, Ano89b, Ano89c, Ano89d]. Backward [Kal81b, Tho88]. Bad [RT84, RT89a]. Balanced [DR84]. Banach [Ale89a, Ale89b, AGMZ81, AH85, Bol84, Bur81, Ein89, Fra85, GKZ81, Got81, Got86, Hei84, Kor84, LT86, LT88, Oka85b, PR83a, Tal89]. Banach-Space [Tal89]. Banach-Valued [Fra85]. Banachique [LT86]. Bandits [Man87, Whi81]. Based [EvZ88, GF87, GP83b, Hal82b, Qui80, QR84]. Bayes [GW82]. Bayesian [Hij84]. Become [FK86]. Behavior [BM89a, CR84, Chi82, GGL88, Gre84, JP84, KZ83, MOT85, Pit81a, PT83, Pyk80, Stu82b, Stu84]. Behaviour [AB86, AS87, And85, Csö81, Hal80b, KKR87, MS82b, Ryz86]. Bell [Gaw84]. Bell-Shape [Gaw84]. Bernoulli [CCN87, CCG^{+}89, Kie80b, Pin83]. Berry [Doe82, Hel81b, QR82, Woo83]. Bessel [DeB87]. Best [Bru84, Joh85, Kla88, Pou84]. Better [MS82a]. Between [GS85, Jur85, KQ89, NZ89, Ell80b]. Beurling [Sol84]. Biased [SW84]. Big [CCFR83]. Biharmonic [Van84]. Binary [Gri83a, HB87, New87]. Binomial [KQ89]. Birth [Bra89, Kas88, Pit81b]. Bjerknes [BG81b]. Blackwell [EMO84]. Block [Dun80, GOD80]. Blocks [Rus88]. BMO [Pro80]. Bochner [Oka81, Oka85b, Sat81]. Bodies [Gor88]. Books [Gri87, Ney86, Pro86]. Bootstrap [FR89, Hal88b]. Borel [Bla80, Bla81, RR80]. Bornée [Ber89b]. Bose [Way80]. Bound [Ber83a, BK89, CR85, Chi82, Chi85, HJ88, Hen83, Iye86, QR82, Sla88, Tak81]. Boundaries [BC83a, Esc87, GP83a, Pin85a, Pin85b]. Boundary [Bur88, Bur84, Cra87, Hol83, JJS88, KV83, Lai88, Nis81, Pin89, Sie82, Tak82, Uch82a, WK87]. Bounded [FI87, GF87, Jan85, MZ84, Pin85d, Pin87, Pin90]. Boundedness [Tal88a]. Bounds [BB83, Bir88a, Bro80b, CW82, Cac82, CK83b, Dal80, Dal88, Deh82, EvZ88, Fil88, Hal82a, Kla80, Kla84a, Mas81, MSS82, dA88]. Bracketing [Oss87]. Branching [Ath88, Coh85, Coh89, CG85, DFG89, FP81, Gor83, Gor84, Gre84, Gre85, Gre80, Hol82, Hug85, Isc88, Isc89, JN85, Jan83a, Key87, Kuc82, LS87a, LS88, LS89, LM89, Uch82b, YM85]. Breiman [AC88b, Bar85]. Bridge [CW82, JK83]. Brownian [Adl80, Gor84, Adl78, BC83a, Bas88, Blu83, Bur85, Bur87a, Bur88, Bur89, CW82, Cox82, CG85, Cra87, CHM89, CCFR89, DP85, DS88a, Dur82, Dyn88a, Dyn88b, Eps89, Eri80, Esc87, Fal87, GHR84, GM83b, GG86a, Gol84, Gor83, GP83a, Gro83, Har81, HR81b, HS81, HLS85, HW87, Hug85, Imh85, Isc88, Isc89, Jen85, JK83, KS84, Kau89, Ken87, Kin80, LS87a, LS88, LS89, Le 86, LM89, Mar86, Mat88a, Mit83c, Mou89b, Pen89, Per81, PT88, PY86a, PY89, Pit89, Ros85, Ros86a, Uch82a, Wal82, Wen80a, Wen80b, Wil85, Wu82]. Brownien [Gol84, Le 86]. Burgess [Wei85]. Busy [Ros83].

C [Mic82]. C. [Mic82]. C.D.F. [JK83]. Cadlaguity [AF84]. Calculation [Ric82, Ros83]. Calculus [BC86, Ber88b, Haj82, KK85, Wat87]. Cantelli [Tal87b]. Capacitary [Kan89]. Capacities [Poo82]. Case [AA87, Cog80, DP85, DST89, MZ84, Pin89, Stu84]. Catastrophes [De 80]. Cauchy [PY86b, PT83]. Cell [Zah88]. Central [Bra85, HHK83, Wat84, AP86, Ale87a, AD87, Bar86, BP84a, Bha85, Bin81, Bir88b, Bol82b, Bol82a, Bol89, Bra88, Che88, CG84, DDP86, Deu88, Dut89, Eri81, GM83a, Got86, Hae84, HJ88, Hae88, HK81, Hal80a, HH81, Hal82a, Hal83b, Hal84, Hal88a, Han89, Han91, Her83, Her84a, Her84b, Kip86, KT87, Kot86, Kuc89, Kur81, Kur84, LR86b, LR87, Oss87, Pel87, Pol82, RT84, RT89a, Sch86, Swe80, dA83b]. Certain [Dar83, Don87a, Jan83a, JD83, NK83, NW81, Str86b, Tal88b]. Cervonenkis [Ale87a]. Cesaro [Mor83]. Chain [Ell88b, EW89, Lyo83, dA88]. Chains [Atk86, BC83c, CI78, Cog80, DF80a, Kee82, Kur81, Kur84, Mat88b, MS89, Naw82, Zac83, CI80]. Change [DPSW86, JJS88, McC85, Str86a]. Change-Point [JJS88]. Changed [Gzy80, Hug85, Le 81]. Changement [Str86a]. Changes [Glo81a, Glo81b, Kas88, Kur80b]. Changing [Imk86, MN86]. Channel [NS82]. Channels [GOD80, Kie80b]. Chaos [EK89, Kwa87]. Characterisations [GG86a]. Characteristic [Csö81, Lac89, Mar81, Yam84]. Characteristics [Goo88]. Characterization [Ale89a, Bur81, GH85, Hoo87, KO88b, LLP80, LT88, Luc84, MN86, MS80b, Per81, Ter83, Wei84]. Characterizations [Jac84, Mas82b]. Characterize [Tay89]. Characterizing [Hal80a, Rog81]. Chemical [Kot86]. Chen [AGG89]. Chernoff [Hog86, Kla85a]. Chi [Hal83a]. Choice [Bru84]. Chosen [Yam84]. Circle [Hol81a]. Circuit [Kel87, Mac81]. Circuit-Switched [Kel87]. Class [AGKdV89, BS87a, BC83a, BL88, BC83c, Bru84, BS87b, CC89, ENS87, Ell84, FI87, GOV87, Gun80, Jur82, MS82a, MS89, PV81, Pin83, Pit85, Pru83, RS88, She81, Shi83, Sil83, BL97]. Classe [RE83]. Classes [Ale87a, BS80a, BG84, Dud87, LLP80, Tal87a, Zak81, RE83]. Classical [Eng81, Kas85]. Classification [Pin85a, Sil80, Wil85]. Clock [RW86]. Close [PT88]. Closure [AL82]. Closures [Pit83]. CLT [HT83, RY89]. Clump [Hal86]. Cluster [Ale89a, Ale89b, GKR88, Gan89, Kue81, dAK83]. Clustering [BG80, CG86, Pol82, Zam84a]. Clusters [Ber83a]. CMJ [Coh85]. Coalesced [Pit87]. Coalescing [Arr81, Cox89]. Coarseness [Fil88]. Codes [GOD80]. Coding [AP89, Dun80, GOD80, Kie80a]. Coefficients [Bha85]. Coharmonic [Sil80]. Coin [Rei82a, SW84]. Coinvariant [Sil80]. Collision [KM87]. Collisions [MS80a]. Combinations [GS81, Hel81b, Joh85, VV82]. Combinatorial [VF87]. Comments [Hog86, Lyn83]. Common [KO88a]. Commute [Beq89]. Compact [GH85, Jan83a, Pru83]. Compactifications [Glo80]. Compactness [DR84]. Comparaison [Bro80a]. Compared [KS87]. Comparing [EG81]. Comparison [Arr85, LT89, Mil81, SC84, Tud89, Bro80a]. Comparisons [HK82, Kin80]. Compensators [Bro83c]. Complete [AK80, Gut85, Hsu89, Sch87b]. Completely [Cer86]. Complexes [Zah88]. Complexity [CGG89]. Components [Sha83, Spi81]. Compound [Ber80a, Ell88a, Ser86, Ser88]. Compounding [Sha88]. Compression [Dun80]. Concave [Gro83, Han88]. Concept [PR85]. Concepts [BSS82]. Concerning [Ell80b, VF87, Wei85]. Conclude [Hil82]. Condition [Beq89, Bra80, BW86]. Conditional [BM81, CL89b, Csi84, Fad85, Fal87, Gar82, Hil82, Hol81b, JJS88, KO88b, Lal84, LS87a, LR86b, LR87, Muk84, Ram81, Shi83, Stu86, Vas80, Zab80, vPvS85]. Conditioned [DeB88, GP83a, HK84, Pin85d, Rit81]. Conditioning [Car88, Jak86, UZ89]. Conditions [AK80, BY88b, Bar88, CD81, DDP86, Ell80a, Fad85, GG86a, GH81, Hal87, Hel81a, JO80, LT86, MS80b, Nis81, Pin88, Tal88a]. Cones [Bur85]. Conjecture [Alp85, Coh81, Ell80b, GR86, Mat82, She82b, VF87].

Connecting [Bur88, DS88a]. Consensus [Cox89]. Conserved [BR84b]. Constants [Joh85]. Construction [AGKdV89, Bar81a, Hu88]. Constructions [BM89b, Mas89b]. Contact [And88a, BG81a, Dur80, DG83, DL88, DS88b, DST89, GP87, Gri83a, Sch86, Sch87b]. Content [HKP80]. Contiguity [The89]. Continues [ESY83]. Continuity [BD88, Bar88, BS86, Cuz82, CD82, Ken89, Ros87, Ros88, Sam88]. Continuous [BK85, ESY83, GF87, Hae88, Kar82, Man87, Nua84, Pic86, Pin83, San88]. Continuum [Hal85a]. Contribution [Pru87]. Contributions [CGG89, Tho86a]. Control [KS84, LL86, LV83]. Controlled [Kus85]. Converge [Hil82]. Convergence [AK80, Ast81, Avr88, BY88a, BY88b, BP85, Bir88b, Blu83, Bol82a, CRW85, CL89b, CCHM86a, DDL86, Doo89, Eri81, Fal84, FR89, Fil83, Fol84, GG86a, GG86b, GH81, GK89a, Got86, Gre80, Gun80, Gut80, Gut85, Hae84, HJ88, Hae88, Hal80a, HH81, Hal82a, Hal82c, Hal83b, Hal84, Hal88a, Hal88b, Hel81a, Hol85, Hol81a, Hwa80, Jan88, Kuc82, LWZ81, Lig89, MR87, Mar81, Mas84, McC80b, Nor84, Ome88, Pin85d, Pou84, RT84, RT89a, Sam84, Sch87b, Sil83, Stu86, Swe80, TC83, Zab80, dHR89]. Converse [Fit87, GJ86, Hal81]. Convex [CHM89, Dyk85, Gor88, GP83b, Han88, Pri82]. Convexification [AH85]. Convexity [Ney84]. Convolution [ENS87, Uhr84]. Convolutions [BK82, HKP80, OR85]. Cooptional [Jac84]. Correction [Adl80, Ale87b, BE86, Ber80b, Ber84, Bha80, BL97, Bro83b, CI80, Cuz87, ER83, Far87, Gor84, Han91, Kes87a, LR82, Nol92, Phi86, Pin90, RT89a, Rob83, SP81, Ser88]. Corrections [Bro83b, Rob83]. Correlated [BR84a, Pit82]. Correlation [And88b, Bro87, CCG^{+}89, Cox84]. Correspondant [Bro80a]. correspondence [Bro80a]. Corresponding [Rus88]. Cost [DH81, Lor81]. Countable [Zac83]. Countably [Ast81]. Counterexample [Imk86]. Counts [Hal86]. Coupled [Gre84, Gre85]. Coupling [CL89a, LR86c, Tho86b]. Coverage [Hal85b, Hol81a]. Covering [Ber81, Mat88a, Mat88b]. Cox [GW82]. Crabgrass [BDS89]. Cramér [SRP85, VV82]. Creation [Get88]. Creative [Shi89]. Criteria [Bha78, DH81, Key84, Bha80]. Criterion [Dun80, GG82, Kla85b, Lyo83, Pin86]. Critical [AW85, Ath88, BM87, CR85, CG85, DP85, DFG89, Deu89, DST89, GH81, GL82, Hol82, Isc88, Isc89, Lig83b, LM89, YM85]. Croissant [Bro80a]. Crossing [JJS88, Lai88, Sie82, WK87]. Crossings [Lin85, NSK86, PY86b]. Csörgo [Sha89]. Cube [Ald83, Ber88a]. Cubic [CR85]. Cumulants [BP83a]. Curvature [Dro82]. Curves [Zei89]. Curvilinear [Tak82]. Cut [Bur89]. Cyclic [BG89].

Dade [Avr88]. Dans [LT86]. Death [Kas88]. Decision [DH81]. Decomposability [GW82]. Decomposition [HB87, Kar82, Ken82, New87]. Decoupling [KLS82, Kwa87, MT86]. Degenerate [Bez87b, Hol83, Kli87, Nis81]. Densities [Bar85, BH80, EK89, Fal84, Gaw84, GH80, Han88, Ome88, Rei86]. Density [CCG^{+}89, Gor83, Gor84, Jen85, KS84, Rei82b, Ric82, Ros84b, Stu82a]. Denumerable [Lal84]. Dependence [BF85, BSS82, Ebe86, FT85, JD83, Joe87, Lee85a, Lee85b, McC87]. Dependency [BR89]. Dependent [AGKdV89, CS87, Dav83, DP82, Dup88, GKR88, GH81, Her84b, Jan83b, Jan84, Jan88, Kle89a, Kuc82, NW81, Rup82, Sil83, Tak81]. Derivation [AK84]. Derivatives [Gaw88, Zak85]. Design [LL86]. Detect [KM87]. Detection [Lin80, Ste86a]. Determination [Yan86]. Determine [Hal83b]. Deux [Bro80a]. Development [Kie84]. Deviation [Arr85, CS85a, Chi85, EW89, GS85, MS89, OP88, RT89b, Rob77, Rob83, Str86b]. Deviations [Bez87a, Bol84, CS87, Deu89, Dup88, Ell84, Ell88b, FO88, Fre85, Gri83c, GS81, Gut80, IM89, Lee88, Lee89, LS87b, Ney83, Ney84, Ney86, NN87b, SRP85, Sie82, Sla88, VV82, dA83b, dA88]. DFR [Sha88]. Dichotomies [The89]. Dichotomy [Eag81]. Difference [Bur81, Kle89b]. Differences [Elt81]. Differentiability [Oka85a]. Differentiable [Bor85, Pic86]. Differential [Heu86, Pro85a, Rei83]. Differentials [Pro80]. Differentiation [Ell86]. Diffusion [Ber88c, Bro86, CL89a, Deu88, EG87, Fre85, Hel81a, Hij84, Kar84, Ken82, Kot86, LR84, MR85, MNS89, MOT85, Oel88, Pin85d, Pin85a, Pin85b, Pin86, Pin88, Sch85a, Str86b, Zei89]. Diffusions [BE85, BE86, Bez87b, Bha78, Bha85, Cla87, DeB88, GGL88, HP86, Kli87, Le 81, LR86c, Nis81, Oka85a, Pin87, Pin89, Pin90, Rog81, Sha80, Bha80]. Diffusive [CG86]. Dilatantes [RE83]. dilatory [RE83]. Dimension [Adl78, Adl80, Gra86, Hen83, Mou89b]. Dimensional [And88c, Bol89, Bro86, CR85, Cox82, CG86, Dur80, Dur82, Eri80, GP87, GKR88, GG86a, Gri83b, Gri83c, Hal85b, Hol85, Kau89, Kel87, Kuc82, Lal86, MOT85, Oka85a, Ros80, Sil85]. Dimensions [And88a, Dur84, Jan87, Law89, Mal88, Sch87a, Sch87b, Sla89]. Dini [Che80]. d’Intégrabilité [LT86]. Directed [Kur80a]. Dirichlet [HHL81, Sie83, Van84, Yam84]. Discontinue [CMEM80]. Discontinuous [JS87, CMEM80]. Discrepancy [Wri81]. Discrete [GS85, GOD80, Hae88, KV83, KKR87, OR85, Van84]. Discretized [BF85]. Discussion [ADG^{+}84, Bar81b]. Disjoint [VF87]. Dispersion [BG80]. Distance [Ell80b, KQ89]. Distances [BB83, JJ86]. Distributed [Elt81, Gri86, Gut83, Sch85b, Ulb81, vZ82]. Distribution [Ber88a, BK89, BC83c, CW82, Che81b, Cog85, Cre80, CM86, CCHM86a, CHM88, Deh84, Don87a, EvZ88, EG81, GP83b, HM87, Hal83a, HLS85, Hoo87, HM81, Kie80a, Kuc82, LS82, LWZ81, MZ84, Mas81, Mit81, Nai81, NSK86, Sch87a, Sil83, Smi82, Stu86, Tuc80, vZ82, Mit83b]. Distributional [HHL81, JS87, Tho86b]. Distributions [AL82, BR89, BS80a, BS80b, Bro83a, Bro83b, BG84, BK85, BM81, Che85, Che81a, Coo85, CH88, Dal88, Ell88a, Gar82, Gru87, HW87, Joe87, KQ89, Key87, Kle89b, KO88b, LLP80, Lia85, Lia88, MS82a, MO83, Nor84, OR85, Pic86, Pit83, Rei82a, Roo80, RS86b, Rüs81, Sch85b, Tch80, Ter83, WRL82, Yam84, Yam82]. Divergence [Oel88].

Divergence-Free [Oel88]. Divergent [GF87]. Divisible [BH80, BR88, EG81, Han88, Yam82]. Divisors [Llo84]. Do [Rei82b]. Doleans [Avr88]. Domain [HK80, HMV83, Kue85]. Domains [Fal87, GH85, GJP84, Hu88, JO80, Mal80, MO83, Pic86, Swe85]. Dominating [Ney83]. Domination [Lia85]. Donné [ESY83]. Donsker [Dud87, Pin85c, Tal87a, Wat89]. Doob [DeB88, MP80, Ram81]. Double [KW87, LW82, RT81, RW86]. Drift [Kla83, Wen80b, Zha89]. Drifts [Tud89]. Driven [Pro85a, Pro85b]. Dual [FP81, Glo80, Oka85b, Sat81]. Duality [CS85b, Gra86, Kas85, RN80]. Dynamic [BM89a, Kar84]. Dynamical [GOV87]. Dynamics [Fri87]. Dyson [BM87].

Early [Doo89]. Edge [GP87, Kuc89, Sch87a]. Edges [Ste88]. Edgeworth [BR82, Got81, Got89, Hal87]. Effect [Hal88a, MS80a]. Eigenfunctions [Tay89]. Eigenvalue [BY88b, Sil85, NN87a]. Einstein [Way80]. Elementary [MOT85, vE86]. Elements [TC83]. Elliptical [HHM82, Luc84]. Empiric [Kuc82]. Empirical [AB86, AS87, Ale84, ADG^{+}84, Ale87a, Ale87b, Csö81, CCHM86a, CCHM86b, CH88, Deu89, EM88a, EvZ88, Ell88b, EW89, FO88, GZ84, GZ86, GKZ81, GP83b, KD80, Lac89, LT89, Mar81, MZ84, Mas81, Mas84, MV87, Mas89b, NSK86, SW82, Sil83, Stu82b, Stu84, Stu86, Tal87c, Wri81, vZ82, Gae88]. Empty [Kue81]. Entropy [Bar86, BF85, Dud87, Dun80, KV83, NS82, Oss87]. Environment [Coh89, Gre85, JN85, Kal81a, Key84, Key87, Mad86]. Environments [And88c, Cog80, Naw82]. Equal [Smi82]. Equation [AK84, Far86, Far87, Hij84, Kla88]. Equations [Bez87a, CN88, Fre85, Haj82, Heu86, Kle89b, MP80, MOT85, Pro85a, Pro85b, Rei83, Tud89]. Equiconvergence [Muk84]. Equilibrium [BR84b, Hol85]. Equipartition [AC88a]. Equivalence [DPSW86, The89]. Equivalence/Contiguity [The89]. Erdos [AW89, CS81, DDL86, DD87, Lyn83, Mas89a, Ste86b]. Ergodic [Bar85, Bec81, CK83a, GW82, Jaj87, KR81, Lig85, McC87, Sin89, Der83, Sin88]. Ergodique [Der83]. Error [GOD80, Kla80, Mar82, Roo80]. Escape [Cox82, Eri80, Gri82, Gri83b]. Esseen [Doe82, Hel81b, QR82, Woo83]. Estimate [Eng81, Rei83]. Estimates [Bas88, Car83, FR89, GGL88, Gri83c, JP87a, LR87, Ros84b, dHR89]. Estimation [Hij84]. Estimator [Smy80]. Estimators [GW82, Hal82b, JS87, Stu82a]. Euclidean [Gol84, Ste81, Ste86a, Ste88]. Euclidienne [Gol84]. Evaluating [Pin85c]. Evaluation [SS89]. Events [RS80, VF87]. Every [Pen89]. Everywhere [RS86a]. Evolution [Bez87a]. Evolutions [Wat84, Wat85]. Exact [Bol82a, DDL86]. Example [Her84a, KW88, Mah82]. Examples [Rei82b, Sch85a, Wei85]. Exchangeability [Zam84b, Zam86]. Exchangeable [DF80b, HLS87, Joh85, Kal88, KT87]. Exclusion [And88b, Arr83a, Arr85, BF87, DPSW86, Fer86, Kip86, Lig80, Saa87]. Excursion [Blu83, Kas85]. Excursions [Bur87a]. Exhaustible [Fal87]. Existence [Bha78, Bha80, Cox82, Fad85, GJ86, Jaj87, Nor84, Pen89, SP81].

Exit [BC83b, Jen85, Lia85]. Expansion [HW82, Hal87]. Expansions [Aar86, Bar87, BR82, EK89, Got81, Got89, Swe80, WK87]. Expectation [CL89b]. Expectations [HK82, Kla81, Muk84, Zab80]. Expected [HW82, MO82, RS86a]. Experiments [LL86, Li80]. Explicit [JK83]. Exponential [Bro83a, Bro83b, BG84, CK83a, Dal88, HW87, Hol81b, Lig89, Mar84, Sha83, Web88, dA80]. Exponentials [Avr88]. Exponents [BM87, HHM82, HM81, HJV86, Key88, Lig83b]. Extended [Eag81, Hor84, Mas89a]. Extension [ARS87, BM82, BF88, Hu88, McC80a, Pro80]. Extensions [Kla80]. Extinction [FP81, LM89]. Extrapolation [Sol84]. Extrema [Don87a]. Extremal [LR88]. Extreme [Ber88c, CM86, HM87, MO83, Mas82a, O’B87, Ome88, Pic86, Roo86, Ser82, Swe85]. Extremes [Ber74, Ber80b, Ber82b, Ber84, DR84, Deh86, FK86, Hal84, Nai81, dHR82].

F [Don87b]. Factorisation [Don87a]. Failure [BS80b]. Families [Hol81b]. Faster [BC83a]. Feller [Bai89, CZ86]. Feynman [Pin86]. Fidelity [Dun80]. Field [Adl84, FO88, Oel88, Yan88]. Fields [Bol82b, DPSW86, Fol84, HB87, Hol83, Ken89, LR86a, New87, Nol88, Nol92, Ros84a, Sie88, Sol84, Zac83]. Filtering [KK85, KO88b]. Finding [Mia88]. Finetti [DF80a, Res85, RS88, Zam86]. Finetti-type [Res85]. Finite [AP86, Ari88, DF80b, DL88, DS88b, DST89, Ell86, GG86a, Kla84a, Lal84, Lig83b, Naw82, Pit83, Rob77, Rob83, Zam83, Zam86]. Finite-Dimensional [GG86a]. Finitely [Ast81]. Finiteness [Pin86]. First [Bro83a, Bro83b, BC83c, Cog85, Fin82a, Hij84, Jen85, Kes87b, Sha88, Tak82, ZZ84]. First-Exit [Jen85]. First-Passage [Kes87b, Sha88, ZZ84]. Fisher [Ste86a, The89]. Fit [BB83, GS81]. Fixation [BG89]. Fixed [RS82]. FKG [FDS88, She82b]. Fleming [DH82]. Flows [BH86]. Fluctuating [Mad86]. Fluctuation [DPSW86, Le 88]. Fluctuations [BR84b, DFG89, GP87]. Flux [BG89]. Fold [Alp85]. Form [Zam86]. Forms [CRW85, FT80a, FT85, Gaw88, KS86, KW87, MT86]. Formula [Dro82, Esc87, JK83, Pin88, Ros86a]. Formulae [Ell86]. Foundation [Doo89]. Free [Oel88, Yan88]. French [Ber89b, Bro80a, CMEM80, Der83, ESY83, Gol84, Le 86, LT86, RE83, Str86a]. front [Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano82e, Ano82f, Ano82g, Ano82h, Ano83e, Ano83f, Ano83g, Ano83h, Ano84e, Ano84f, Ano84g, Ano84h, Ano85e, Ano85f, Ano85g, Ano85h, Ano86e, Ano86f, Ano86g, Ano86h, Ano87e, Ano87f, Ano87g, Ano87h, Ano88e, Ano88f, Ano88g, Ano88h, Ano89e, Ano89f, Ano89g, Ano89h]. Frontier [LS87a]. Function [Cac82, Car83, Csö81, CCHM86a, Dal80, JS89, Lac89, Mar81, McC85, Mia88, Mil81, Pin85c, Pin85b, vZ82]. Functional [BP84a, GK89a, Got89, Han89, Han91, Her84b, NP88, PV81, Ser82, Zee81, Zei89, dA83b]. Functionals [BP83a, Eps89, HHL81, Kin80, Oka81, Pin86, Ste81, Wat87, Yam85, Zak85, dA88]. Functions [BB83, BPT88, Cox82, CCHM86a, DR84, EvZ88, GZ86, GP83b, JS87, Kla81, Kon82, LS82, Mar86, Mas81, Mas82b, NSK86, New87, Pin87, Pin90, RP81, Rüs81, Sha83, Sie83, Sil83, Sil80, Stu86, Uhr84, Van80, Van84, Yam84]. Fundamental [Ber83b]. Further [Bro81, Ete80, PY89]. Fuzzy [PR85].

G [Ros83]. Galen [Gae88]. Galton [Kle89a, Kus85, VN82a, VN82b]. Gambler [KS87]. Games [Bla81, Now85]. Gamma [Kle89b]. Gauge [Dyn88b]. Gaussian [Ber80b, Ber84, ER83, Adl84, AS87, Ale86, Ber74, Ber80a, Ber80c, BK89, BK85, BH83, Cuz78, Cuz82, CD82, Cuz87, Dar83, De 80, Eag81, ER82, FT80b, Gor88, Heb80, Hwa87, JM83, KLS82, Kon82, Kue85, Lin85, McC80a, McC80b, MP87, Oka81, Sam88, Sat81, Tal88b, Yan88]. GB [Tal88b]. GB-Sets [Tal88b]. General [Alp85, Ber82a, BL88, Car86, Deh84, Ell84, Gra86, HHK83, Hil87, MP80, Pru81a, BL97]. Generalised [Zah88]. Generalization [Eme82, Hu88, Mic82]. Generalizations [GM83b, Hil82]. Generalized [Bar85, Csi84, Fer80, Gru87, HLS80, Hoo84, Kal81a, Kle89b, MOT85, Ros81, SRP85]. Generated [OV85]. Generating [Mia88]. Generation [CDG87]. Generators [Gzy80, Yan88]. Genetics [MS89]. Geodesic [Lia88]. Geometric [BPT88, Coh81, FP81, Kle89a, Mat82, Sha88, Ste81]. Geometrical [Bur81]. Geometry [GM83b, LT89, Tal87a]. Getoor [GR86]. Gibbs [Ber89a, FO88, Gan89, Lal86]. Gine [ADG^{+}84]. Gittins [Kar84]. Given [AW89, Bar81a, ESY83, Kie80a, Pri82, Tch80]. Giving [Pin88]. Glivenko [Tal87b]. Global [Per81]. Goodness [BB83, GS81]. Gradient [Fri87]. Graphs [BR89, Gri87, Jan88, NK83]. Group [FOW85, HJV86]. Groups [Bec81, Jan83a, KV83, Wat89]. Grow [BC83a]. Growth [BG81b, Coh89, Dur80, DL81, KKR87, Kla84b, Kle89a, Kus85, Mil81, Pit85, Pru81b, Rit81, Ste81, Ste88, Uch82b, Bro80a].

H [GR86, Pit89]. Harmonic [Mar86, Pin87, Pin90, Sie83]. Harris [Cox84]. Hartman [dA83a]. Hashing [Pit87]. Hausdorff [Hen83, LR86a]. Hausdorff-Metric [LR86a]. Having [FT85, HW87, Sch85a]. Heat [Hsu89, Wat87]. Heavy [HK84]. Heyde [Mic82]. Hierarchical [BM87]. High [And88a, BK82, CCG^{+}89, Gor83, Law89, OR85, Sla89, Gor84]. High-Density [CCG^{+}89]. Higher [Jan88, Mal88]. Highly [KW88]. Hilbert [Ber80c, CM89, Jaj87, LR84, RT84, RT89a]. Hitting [CK83a, Ken82, Lia88, Pit89, Tak82, Wen80a]. Hoeffding [BF88]. Homeomorphisms [Har81]. Homogeneity [Kal87]. Homogeneous [EK89, Jan83a, Lin81, Str86b]. Homogenization [Oel88]. Hopf [Don87a]. Hull [CHM89]. Hungarian [BM89b]. Hunt [Maz88, GR86]. Hydrodynamic [DFG89]. Hydrodynamical [AK84, BF87, Gre84]. Hydrodynamics [PS83]. Hyperbolic [Haj82]. Hypotheses [Tom80]. Hypothesis [Doe82, GR86].

I.I.D [CK83b, DH84, Deh86, HR83b, HK82, Kla89, MZ84, Sla88]. Ideal [MR87, RY89]. Identically [Elt81, Gri86, Gut83, Ulb81, vZ82]. Identifiability [BK85, LS82]. IFRA [BS81, ENS87]. II [Ald89, DS88b, LS89]. II. [NN87b]. III [DST89]. Image [Hwa87, Pri82]. Immeasurable [Coh86]. Immigration [Key87, VN82a]. Improved [CS81, Lig85, Ste86b]. Improvement [Kla88]. IMRL [Bro83b, Bro83a]. Increasing [BS80b, BPT88, Kas82]. Increments [Ber83a, CCFR83, CF86, CS81, DS89, GFZ86, HR83a, HR83b, HR89, JP87b, Lin81, LS87b, OV85, Rev82, Ste86b, Ver85]. Independence [Ber82a, FR88, GS85, JD83, Pit89, UZ89, Wen80b, vPvS85]. Independent [Bro87, Ein87, FK86, Gri86, Gut83, Hal80b, Hal83a, Joh85, JS89, Ken85, Kla81, LW82, Lin81, LS87b, MT86, O’B80, Rei82b, Rei86, RS88, SC84, Ser82, Ste86a, Tal89, Tom80, Ulb81, vZ82]. Indeterminate [Ber88a]. Index [BD88, Kon82, Mit81, Mit82, Mit83b, Nol88, Nol92]. Index- [Nol88, Nol92]. Indexed [Ale86, AP86, BP84a, BP84b, GZ86, GZ87, GG86a, GG86b, Gut83, Hur85, Kur80a, LV83, SW82, WW81, BP85]. indexes [Bro80a]. Indices [Bro80a, Gut80, Kar84, LW89]. Inequalities [Ale84, Ale87b, BPT88, Bro80b, BR88, Bur84, CW82, Che88, De 81, Ell88a, HKP80, Joh85, KLS82, Kwa87, MT86, MS80b, SW82, Tch80]. Inequality [And88b, Arr85, Bur87b, Che81b, Cox84, JDPP83, Kla85a, Kla89, KS87, LR86a, MV87, MP80, RT89b, She82b]. INF [HR89]. Infection [Sch87b]. Inference [Don87b]. Infinite [AA87, Bra88, CR84, Cox82, Deu88, Dut89, Eri80, Fri87, GKR88, Gan89, GH81, Gor83, Gor84, Kip86, Law89, Lee88, Spi81]. Infinitely [BH80, BR88, EG81, EG87, Han88, Yam82]. Infinitely-Many-Sites [EG87]. Infiniteness [Pin86]. Infinitesimal [Gzy80]. Influence [Deh86, Hal84, Kes86, Kes87a]. Information [Ano80m, Ano81m, Ano82i, Ano83i, Ano84i, Ano85i, Ano86k, Ano87i, Ano88i, Ano89l, CGG89, Ste86a, The89]. Inhomogeneous [CC89, LS88, LS89, Mou89a]. Inner [RW86]. Instant [ESY83]. Integer [Llo84]. Integrability [GJ86, RW86, Tal89, LT86]. Integral [ARS87, BM82, CCHM86a, Gri83b, KS88, Ree80, Ric82, SS89, She82a, Wat85, Ber89b]. Intégrale [Ber89b]. Integrals [Avr88, Bur87b, ER82, ER83, EG81, GM83a, KW87, Lin81, NZ89, Roo80, RW86, Str86b]. Integration [Bic81, EK89, Eme82, MNS89, Pro86, Ric82, RW86]. Inter [JJ86]. Inter-Point [JJ86]. Interacting [BG80, Deu88, LR84, Oel84, Spi81]. Interaction [AK84, Way80]. Intermediate [Coo85, WRL82]. Interpolator [Pou84]. Intersection [Dyk85, Dyn88a, Dyn88b, Ros85, Ros87, Ros88]. Intersections [GHR84, Pen89, Ros84a, Ros86a, RS82]. Interval [Pyk80]. Intervals [Ber81, BD86, SW82]. Intrinsic [Per81]. Invariance [CHS87, DP82, Deu89, Ebe86, Ein87, Ein89, KP80, NW81, Pel82, Pel85, Phi80, Phi86, Wat89, dA82, vPvS85]. Invariant [And82, And85, Bha78, Bha80, JS89, Key88, Kli87, Pin85a, Tal88b, Wil85]. Invasion [CCN87]. Inverse [Imh85]. Inverses [Hal82b]. Investigating [NK83]. Investment [AC88a]. Involves [Get88]. Involving [Che81b]. Invulnerable [GOD80]. Irreducible [BL88, BL97]. Ising [Hol85]. Isomorphism [Kie84]. Isoperimetry [Tal89]. Isotropic [BH86, Ber80c]. Issue [Ano89j]. Itô [RW86]. Iterated [Ale84, Ale87b, Avr88, BP84a, BE85, BE86, Dev81, EM88a, Fer80, FT80a, GK89b, HM87, Hug85, Lac89, LW82, LT88, LW89, MZ84, PV81, Pru81a, Sta84, dA83a, dA83b]. Iterative [Dyk85].

Joining [Ell80a]. Joint [CD82, Ros87, Tuc80]. Jon [Gae88]. Jump [BC86]. Jumps [Le 81, MR85].

Kac [Ano86i, Ano86j, Kes86, Kes87a, Pin86, Tho86a]. Kahane [Hwa87]. Kakutani [Pyk80]. Kamae [Ell80b]. Karr [Don87b]. Kazamaki [Pro80]. Kernel [Stu82a]. Kernels [Jan85, MPvW83, Wat87, Wei84]. Kiefer [Rup82]. Killing [Get88]. Klass [Ree80]. KMT [MV87, Mas89b]. Kolmogorov [Ano89j, Ano89i, Ano89k, CW82, CGG89, Doo89, Dyn89, Hu88, Shi89, Sin89]. Krengel [Sin88].

Lag [HR83b]. Laplace [BS80a, ER82, Hwa80, Ros83, Zee81, ER83]. Large [Arr85, AK80, BP84b, Bez87a, Bol84, CS85a, CS87, Chi85, CZ86, CN84, De 81, Deu89, Dup88, Ell84, Ell88b, EW89, Elt81, Fil83, FO88, Fre85, Gem86, GZ87, GS81, Gut85, Hal82c, HR81a, Hei84, IM89, Kot86, Lee88, Lee89, LS87b, Mas82a, Mas89a, MS89, Ney83, Ney84, Ney86, Oel84, OP88, Pin85d, PT83, PR83a, PR83b, Rob77, Sch87b, SRP85, Sie82, Sil85, Sla88, Str86b, VV82, Vas80, Wri81, dA88, NN87b, Rob83]. Largest [BY88b, Hol81a]. Last [Lia85]. Lattice [CR85, Deu88, Sch89]. Law [Ale84, Ale87b, And88c, AK80, BY88a, BP84a, BP84b, Ber86, CZ86, De 81, Dev82, Fil83, HM87, Hal82c, HR81a, HMV83, Kon82, Kot86, Kue85, LW82, LT88, LW89, MZ84, Mas89a, PV81, PR83a, PR83b, Sch85a, Sta84, Stu82a, Vas80, Way80, Wri81, dA83a, AW89, Elt81, GZ87, Gut85, Hei84, HLS80, Lyn83, McC80a, Oel84, Van80]. Laws [AGMZ81, AH85, BE85, BE86, CS81, DH84, DD87, Dev81, EM88a, Fer80, FT80a, Gaw88, GJP84, GK89b, HHM82, Hug85, Kas85, KD80, Lac89, Mas82a, Mit81, Mit82, Mit83b, PY86a, PY89, Pru81a, Pru83, Ste86b, Tuc80, dA83b, Gut83, Hil82, Mas82b, Rus88]. Layers [Wri81]. Lemma [BF88]. Length [CCG^{+}89, KO88a, Ros83]. Less [BD88]. Letter [KO88a]. Level [Adl78, Adl80, Lin85, PY86b]. Lévy [Bar88, EG81, Fit87, GM83a, Hen83, JP88, Jan85, LRS89, McC85, Pru81b, Sil80]. Lie [Wat89]. Life [BS80a, Coh86, LLP80, LS82, Sha83, Shi89]. Lifetime [GH81]. Lifetimes [DeB88]. Lifted [Car88]. LIL [Ale89a, Ale89b, GKZ81, GK89a, Kla84a, KZ83, Kue85, dAK83]. LIM [HR89]. Limit [ADG^{+}84, AP86, Ale87a, AD87, AH85, AT87, Bar87, Bar86, BP84a, BD80, BF87, Ber80a, Bha85, BB83, Bin81, Bir88b, Bol82b, Bol82a, Bol89, Bor85, Bra85, Bra88, CS85a, CS87, Che88, Che81a, CC89, CD81, CG83, CG84, Cox88, Csi84, Csö81, DH84, DR85, DFG89, Deh84, DDL86, DD87, DDP86, Deu88, DL81, Dut89, EM88b, Eps89, Eri81, Ete80, Fer80, Fla82, FT85, Fre85, Gem80, GM83a, GZ84, Gor83, Gor84, Got86, Got89, GP83a, GJP84, Hae84, HJ88, Hae88, HK81, HHK83, Hal80a, HH81, Hal82a, Hal82b, Hal83b, Hal84, Hal88a, Han89, Han91, Heb80, Her83, Her84a, Her84b, Hol81b, HK84, Jak86, JJ86, Jan83a, Kas82, Kip86, KT87, Kot86, Kuc89, Kur81, Kur84, LS87a, LR86b, LR87, LT89, Mij82, NN87a, NP88, Oss87]. Limit [Pel87, Pol82, Pru83, Qui80, RT84, RT89a, Roo80, Roo87, RT81, Saa87, Sch86, Ser82, Shi83, SW82, Sla89, Str86b, Swe80, Tal85, Teu81, Tom80, Tuc80, VN82a, VN82b, Wat84, Woo85, Wu82, Yam85, ZZ84, Zha86, dA83b, dHR82, RE83]. Limite [RE83]. limited [Ber89b]. Limiting [Arr81, Che85, Coo85, Deh82, Hal80b, Key87, Pin85d, Pit81a, WRL82]. Limits [BW85, Dav83, GP89, Tho88, vE86]. Line [Esc87, Hel81a, KM88, Kee82]. Linear [Blo85, Bou87, Doe82, GS81, HW82, Hel81b, Joh85, Kot86, Mas82b, Oka81, Poo82, Pou84, Van80, VV82]. Lipschitz [Eri81, GZ86]. Little [GM83b]. Living [Uch82b]. Load [Smi82]. Load-Sharing [Smi82]. Local [BS87a, Bar88, BE85, BE86, BC86, Bro80a, CS85a, CCFR83, CF86, CCFR89, Cuz78, Cuz82, CD82, Cuz87, Dyn88a, ESY83, GHR84, GP83a, GJP84, Gun80, JP84, JP87b, KS84, Mit83c, Per81, Ros85, Ros87, Ros88, San88, Swe85, Woo85, dHR82, RE83, Bro80a]. Locale [RE83]. Locales [ESY83]. Locally [Spi81]. Log [AC88a, Dev82, Han88, Kon82, KD80]. Log-Concave [Han88]. Log-Convex [Han88]. Log-Optimum [AC88a]. Logarithm [Ale84, Ale87b, BP84a, BE85, BE86, Dev81, EM88a, Fer80, FT80a, GK89b, HM87, Hug85, Lac89, LW82, LT88, LW89, MZ84, PV81, Pru81a, Sta84, Stu82a, dA83a, dA83b]. Long [And85, BG81a, FT85, Lig80]. Long-Range [BG81a, FT85]. Longest [NK83]. Look [Gre80]. Low [Bra89]. Lower [BC83a, CW82, Cac82, Chi82, Chi85, GF87, Hen83, Iye86, JP87a, PF86, Wri81, Zha86, dA88]. Lyapunov [Key88].

M [Ree80, Ros83]. Machlup [Zei89]. Magnitude [BP83a]. Majorant [Gro83]. Majorization [Joe87]. Malliavin [BC86, Ber88b, Wat87, Zak85]. Malliavin-Type [Ber88b]. Manifold [Hsu89, Ken87]. Manifolds [Mar86]. Many [EG87]. Mappings [Bor85, Han89, Han91, Pit83]. Marcinkiewicz [Kor84]. Marginal [Che81a, Kie80a]. Marginals [AP89, Tch80]. Mark [Ano86i, Ano86j, Tho86a, Kes86, Kes87a]. Markov [BL97, CI80, SP81, Atk86, BS87a, Beq89, BL88, BC83c, CK83a, Chi82, CC89, CI78, CS85b, Cog80, DF80a, Dyn89, Ell86, Ell80b, Ell88b, EW89, Fol84, Get88, Glo81a, Glo81b, HB87, Kal87, Kas85, Kee82, Kur80b, Kur81, Kur84, Lal84, LV83, Lyo83, Mat88b, MS89, Mou89a, Naw82, NN87a, NN87b, O’B87, Pit81b, RP81, Ros87, Ros80, Sie83, Yan88, Zac83, Zam84b, Zam86, dA88]. Markov-Renewal [SP81]. Markovian [Cog80, Ell80a]. Martingale [Bar81a, Bol82a, Bur81, Bur84, Che80, Coh85, Elt81, Hae84, HJ88, HH81, Li80, MN88, Oel84, RS86a]. Martingales [Ast81, BM89a, BD80, Bro80a, ESY83, FT80a, Fol84, FI87, Gun80, Hae88, Hoo84, Imk86, Kur80a, Nua84, Pri82, Ros83, San88, Tal85, Zak81]. Maruyama [Wri80]. Masani [Mia88]. Match [FDS88]. Matching [AW85, AW89]. Mathematical [Tho86a]. Matrices [Alp83, Ber83b, Bou87, Bro87, CN84, Gem80, Gem86, Hai85, Key88, Pin83]. Matrix [BY88b, Coh81, Gri86, HK80, Kar82, Sil85]. Mauldin [Wei85]. Max [De 84, Pin88, dHR89]. Max-Stable [dHR89, De 84]. Maxima [HT83, Heb80, Kla84b, Kla85b, McC80a, McC80b, Sie88, Smy80, Tom86]. Maximal [Deh82, Deh84, Deh86, Dev82, JP87b, Jan87, KO88a, McC87, MS80b, Tho86b, Bro80a]. Maximale [Bro80a]. Maximizing [Ken85, Kla89]. Maximum [Ath88, BK89, BF85, CW82, Che81a, DH84, Fin82a, GS85, Kla83, MSS82, Pru87, SC84]. McKean [Pit89]. McMillan [AC88b, Bar85]. Mean [AA87, CR84, Deu89, GK80, Jen85, Kla84a, Kla89, Pou84, She81]. Means [Cra87, Lai88, Mal88, Mor83, Pol82, Yam84]. Measurability [Tal87c]. Measurable [HP83, Now85, Oka81, Oka85b, Sat81]. Measure [Bec81, BH83, Ell88b, EG81, EG87, FO88, FT80b, Get88, GS85, Isc88, Isc89, KW88, KS88, LM89, Oka81, PR83b, Wil85]. Measure-Preserving [Bec81]. Measure-Valued [EG87, Isc88, Isc89, LM89]. Measured [McC85]. Measures [AF84, And82, And85, Ber89a, Bha78, Bha80, BH80, BK83, BM81, BW85, BZ81, DH82, Eag81, Gar82, GK80, Gru87, Hil87, HJV86, Hwa80, JM83, Jur82, Jur85, KR83, Kan89, Kli87, KD80, Luc84, Mah82, Pin83, Pin85a, Rei81, Ryz86, SP81, Tal88b]. Measuring [PT88]. Mechanics [BDS89, ER82, ER83]. Median [Hal80b, LR81, LR82]. Medium [Bro86, DFG89]. Metastability [GOV87]. Method [AGG89, ER82, ER83, Hwa80, Kla81, NK83]. Methods [CL89a]. Metric [CM89, Dud87, LR86a, MR87, Oss87]. Metrics [RY89]. Migration [YM85]. Mills [Iye86]. Mind [Coh86]. Mini [Pin88]. Mini-Max [Pin88]. Minimal [Bur87a, DDP86, Hal87, Hel81a, Kla84b, Ste88, Tay89]. Minimax [Poo82]. Minimization [Ald83]. Minorants [GP83b]. Mins [ENS87]. Miscellaneous [Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano82g, Ano82h, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano84g, Ano84h, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano88h, Ano89a, Ano89b, Ano89c, Ano89d]. Miscellaneous [Ano89e, Ano89f, Ano89g, Ano89h]. Mises [BP83a]. Mismatches [AW89]. Mittal [McC80a]. Mixed [MP87]. Mixing [Alp85, Bol82b, Bra80, Bra88, DDP86, GG86b, HKS87, Her83, Hol88, KP80, Pel82, Pel85, Pel87, Ros84b, Sam84]. Mixtures [BK85]. Mode [Hal80b]. Model [BM87, BG81b, CG83, CG84, CG86, Cox88, Cox89, DH82, DL81, EG87, Lin80, PS83, Pyk80, Sen82, Sol80, Zee81]. Models [Asm80, Dut89, Fla82, GF87, Hol85, LW82, Ros81, Zam84b]. Moderate [Gut80]. Moduli [Ale86, Yan86]. Modulus [Pru87]. Moment [BB83, Bir88a, BPT88, Hal87, Joh85, MSS82, Tom80]. Moments [AGG89, CK83a, CK83b, GJ86, Hal82a, KT87, Mar82, Nai81, Rei83, Tuc80, dA80]. Monotone [CS85b, SS81, Smy80, RE83]. Monotones [RE83]. Monotonicity [Bro80b, Bro81, Ree80]. Monro [Rup82]. Morceaux [RE83]. Mosaic [Hal86]. Motion [Arr83a, BC83a, Cra87, CHM89, Dur82, Dyn88a, Dyn88b, Eri80, Esc87, Fal87, GM83b, GG86a, Gro83, HR81b, HS81, HLS85, Hug85, Imh85, Isc88, Isc89, KS84, Kau89, Ken87, Kin80, LS87a, LM89, MS80a, Mar86, Mat88a, PY86a, PY89, Ros85, Ros86a, RW86, Wen80a, Wen80b, Wil85, Wu82]. Motions [Cox82, CG85, Gor83, Har81, HW87, LS88, LS89, Gor84]. Mouvement [Gol84, Le 86]. movement [Gol84, Le 86]. Moving [BD88, BC83a, DR85, Roo86, Rus88, Sol84, Uch82a]. MR [BL97]. Multi [Man87]. Multi-Armed [Man87]. Multidimensional [Bha78, Bha80, BS81, CL89a, Cla87, Gun80, Gut80, HT83, HK81, HW87, LW89, LR86c, Pin88]. Multilinear [KS86, MT86]. Multiparameter [Bas88, Bec81, Fol84, Kur80b, Man87]. Multiple [HHM82, HM81, KS88, LRS89, Lin81, RW86, SS89, Le 86]. Multiples [Le 86]. Multiplicateurs [LT86]. Multiplicative [Kar82]. multipliers [LT86]. Multitype [Coh89, Hol82, Key87, Kle89a]. Multivariate [BS80b, BF88, BR88, Coo85, Ell88a, Ell88b, HK80, Iye86, Joe87, LS82, MS82a, MO83, Mas89b, Mia88, Pou84, Sch85b, Stu84]. Mutual [BF85].

N [Ano89i, Ano89k]. Nash [GGL88]. Natural [Cox82, Hol83, LR81, LR82]. NBU [Sha83]. Near [Kle89a, Ryz86]. Near-Supercritical [Kle89a]. Nearest [BB83, Bra89, GG82, GL82, Lig83a, Lig83b, Lig89]. Nearest-Particle [Bra89]. Necessary [AK80, BY88b, Bar88, CD81, Fad85, GH81, Pin88, Tal88a]. Negative [BSS82, Kla83]. Neighbor [BB83, GG82]. neighborhood [Gol84]. Networks [Kel87]. Neumann [Nis81]. no [BL97, Bla80]. Noise [DM84, Dup88, KK85, Rup82]. Noisy [GOD80]. Non [HW82, HP83, Kar82, MZ84, Zam84a, Zei89, vZ82]. Non- [vZ82]. Non-Clustering [Zam84a]. Non-I.I.D [MZ84]. Non-Linear [HW82]. Non-Measurable [HP83]. Non-Singular [Kar82]. Nonasymptotic [BM89b]. Noncentral [AT87, FT85]. Noncharacteristic [Pin89]. Nonconstant [BK89]. Nondecreasing [JP87a]. Nondeterminism [Cuz78, Cuz87]. Nonhomogeneous [YM85]. Noninteracting [Lee88, Lee89]. Nonlinear [CN88, KK85, KO88b, Lin80, Sch85a, Ste81, Zha88]. Nonnegative [SC84, Yam85]. Nonparametric [LLP80]. Nonsmoothed [FR89]. Nonuniform [HJ88, LR87]. Norm [Gem80, JK83]. Normal [BR89, Ber88a, Ber86, BH83, Che81b, CCHM86a, Eng81, Goo88, HK80, HMV83, Iye86, Jan88, JDPP83, Nai81, Pit82, QR84]. Normality [Car86, GP89, GP83b, HKS87, Mal88, Mat82, PR85, Ros84b, Tak81]. Normalization [HK80, McC80b]. Normalized [Gri86, GK89b, Zha86]. Normed [GH81, HT83, HK81, Pru83]. Normes [Bro80a]. Norming [Hal88a]. Norms [CH88, Bro80a]. Note [Bha80, Bor85, Bra80, BG81a, CR84, Che81b, CZ86, Gar82, Hae84, Kan89, Mal80, PF86, Ram81, SP81, Sta84, Yu81]. Notes [Adl80, Ber80b, CI80, CM89]. Nuclear [Ust84]. Null [Doe82]. Null- [Doe82]. Number [Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano82g, Ano82h, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano84g, Ano84h, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano88h, Ano89a, Ano89b, Ano89c, Ano89d]. Number [Ano89e, Ano89f, Ano89g, Ano89h, Bru84, BS87b, CDG87, Far86, Far87, KM87, NSK86, She81, dlC89]. Numbers [AK80, Hal82c, HR81a, Kot86, Mas82a, Qui80, QR84, Vas80, Wri81, BP84b, CZ86, De 81, Elt81, Fil83, GZ87, Gut85, Hei84, Mas89a, Oel84, PR83a, PR83b]. Numerical [Ric82].

Obtaining [Dyk85]. Occupancy [Eng81, Qui80, QR82, QR84]. Occupation [CG83, CG85, Cox88, GH80, KS84, Kas82]. Occupied [GKR88]. Occurrence [Li80]. Occurrences [VF87]. Occurring [Zee81]. Ogawa [NZ89]. One [AGKdV89, AP89, And88c, BD88, Bro86, Dur80, GP87, Gaw88, Gra86, Hol85, Kau89, Kel87, Kuc82, Lal86, MOT85, Oka85a, Pru81a, Ros80, She81]. One-Dependent [AGKdV89]. One-Dimensional [Bro86, Kau89, Kel87, Lal86, MOT85, Oka85a, Ros80]. One-Sided [Gaw88, Pru81a]. Onsager [Zei89]. onto [Dyk85]. Operational [Jac84]. Operations [AL82]. Operator [HHM82, HM81, HMV83, HJV86, Jur82, Luc84]. Operator-Selfdecomposable [Jur82]. Operator-Stable [HHM82, HM81, HMV83, Luc84]. Operators [CRW85, Swe80]. Optimal [BS87b, cCS80, De 80, Fin82b, Irl81, Ken85, LV83, Lin85, Lor81, Mar84, Mil85, SS81, Ste82]. Optimality [AC88a]. Optimum [AC88a]. Optional [Hur85, Kur80a, WW81]. Options [Bru84, BS87b]. Order [Ast81, BP83a, CM80, Che85, Coo85, Dev81, FR88, GS81, Hal82b, Hel81b, Hij84, Jen85, JS87, LR86b, LWZ81, Mas82b, Rei81, RS80, Teu81, Van80, VV82, WRL82]. Ordered [Ber81, GR89, Hur85, LV83, Llo84, WW81]. Ordering [Rüs81].

Oriented [Dur84, Kuc89, Sch87a]. Origin [Ryz86]. Orlicz [KW87, Web88]. Ornstein [Deu89, IM89, Tay89]. Orthant [HR81b]. Orthogonal [BM81, Gar82, Mah82, MPvW83, Mor83, Wei84]. Oscillation [Gol84, Stu82b, Stu84]. Other [Bro87].

P [Pit89]. P. [Fit87]. Packets [KM87]. pages [Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano82g, Ano82h, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano84g, Ano84h, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano88h, Ano89a, Ano89b, Ano89c, Ano89d]. pages [Ano89e, Ano89f, Ano89g, Ano89h]. Pairs [Eag81]. Pairwise [O’B80, RS88]. Paper [ADG^{+}84]. Parabolic [KW88]. Parallel [Smi82]. Parameter [CW82, Ete80, Haj82, Imk86, Maz88, McC87, Nua84, Pin89, San88, Sch87b, Zak81, Zak85, FI87]. Parameters [Sol80]. Part [Ken87]. Partial [AP86, BP84a, BP84b, CS81, Dav83, Ein87, Ein89, GZ87, Hil82, JO80, Kla84b, Kla85b, KP80, Mal80, MSS82, Smy80]. Partial-Sum [AP86, BP84a, BP84b]. Partially [GR89, Hur85, LV83, WW81]. Particle [Arr83a, BG80, BG89, Bra89, BR84b, DFG89, Fer86, GL82, Lee88, Lig83a, Lig83b, Lig89, Saa87, Way80]. Particles [Lee89, Uch82b]. Particular [Adl84]. Partition [New87]. Partitioning [Hil87]. Partitions [KR81]. Parts [EK89, MNS89]. Passage [Bro83a, Bro83b, BC83c, Cog85, Fin82a, Kes87b, Ros80, Sha88, ZZ84]. Path [CM80, CS85b, Gri83a, Nol88, Nol92, Pen89, PT88, Ver85]. Paths [Bur85, Bur88, Bur89, DS88a, GHR84, NK83, RW86]. Pattern [AW89]. Patterns [Li80]. Percolation [CR85, CCN87, CCG^{+}89, CD81, CG84, Dur84, GKR88, Hal85a, Kes87b, Kuc89, Sch87a, Wie82, ZZ84]. Period [Ros83]. Periodic [Bha85, Sch85a]. Periodically [JN85]. Permutation [FDS88]. Perturbation [Ste86a]. Perturbations [Bez87a, Bez87b, GOV87]. Petkau [Hog86]. Phase [AL82, CCG^{+}89, Pem88, RN80]. Phenomena [AW85, BM87, GL82]. Photograph [Ano86i, Ano89i]. Physics [BM87, Tho86a]. Picard [GM83b]. piecewise [RE83]. piecewise-monotone [RE83]. Pinned [Ric82, She82a]. Places [Pit89]. Planar [CHM89, Dyn88a, HLS85, PY86a, PY89, Ros86a]. Plane [GHR84, Har81, MN88, Mil85, Pin87, Pin90, Rei83, Ste82]. Poincaré [Che88]. Poincaré-Type [Che88]. Point [Arr81, Ber83a, Fri87, JJS88, JJ86, MN88, OV85, Zee81, Don87b]. Points [Ber81, Bur88, Bur89, DP85, Fer80, Heb80, Le 86, LRS89, Mij82, Mou89a, Ney83]. Pointwise [HHK83]. Poisson [AGG89, Bar87, Ber80a, BK83, Bro83c, DP86, Ell88a, KQ89, MN86, Ros81, Ser86, Ser88]. Polarity [Mou89a]. Polymers [Bar81b, Fre81]. Polynomial [Kwa87, MS82b]. Polynomials [AT87, She81, vE86]. Population [Asm80, CR84, Kle89a, MS89, Rob77, Rob83]. Position [Saa87]. Positive [Kon82, Pin85d, Rit81, Way80]. Positively [Pit82]. Positivity [Lee85b]. Possible [Kla88]. Potential [Maz88]. Pour [Ber89b, Bro80a, LT86, RE83]. Power [Hol83, Ste88]. Pramarts [Tal85]. Precision [Kla80]. Predictable [Kal88, Str86a]. Prediction [Ari88, De 80, Lin80, Lin85]. Predictors [Blo85]. Preface [Ano89j]. Prescribed [AP89]. Preservation [Sha88]. Preserving [Bec81, Oka85a]. Presque [Der83]. Prevalence [Heu86]. Previsible [Str86a]. Prime [Llo84]. Primitive [NS82]. Principle [Bez87a, Deu89, Jak86, NW81, Pel85, Pin89, Wat89]. Principles [CHS87, DP82, Ebe86, Ein87, Ein89, KP80, OP88, Pel82, Phi80, Phi86, dA82]. Priority [Kur84, Mit83b]. Prises [ESY83]. Probab [Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano82g, Ano82h, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano84g, Ano84h, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano88h, Ano89a, Ano89b, Ano89c, Ano89d]. Probab [Ano89e, Ano89f, Ano89g, Ano89h, BL97]. Probabilistic [Dro82, Pit87, Van84]. Probabilities [Ber89a, DR85, Fad85, Gut80, JJS88, Mit83a, PF86, Ram81, Rob77, Rob83, Sie82, Sie88]. Probability [Ale84, Ale87b, Bas88, BPT88, Bol84, CR85, FP81, GS80, Gri83c, GF87, HKP80, Hil87, HJV86, Hwa80, JP87a, Jur82, Kes86, Kes87a, Mat82, MSS82, PR83b, Ric82, Ste81, dA82]. Probability-Based [GF87]. Problem [FP81, HHK83, Hog86, Hol83, Hwa87, Kar84, LL86, Lor81, Nis81, Pit89, QR82, RT89b, Sha89, Shi83, Sie83, Tal87b, Van84, Wat84]. Problems [Bru84, BS87b, Bur84, Irl81, JJS88, Lai88, LV83, Mat88a, Mat88b, Sal87, Tal87c, WK87]. Procedure [Dyk85]. Process [AP89, And82, AK84, And85, And88b, And88a, Ari88, Ath88, BF87, BK89, BG81a, Bro86, BW86, Coh89, CCFR83, CF86, Dar83, Deu89, DL88, DS88b, DST89, Ell86, EW89, Fer86, GP87, Get88, Gre84, Gre85, Gri83a, GFZ86, HHL81, HR83a, HR83b, HR89, Hel81a, Imh85, Kie80a, Kon82, Lin81, Mad86, MR87, MZ84, Mas84, MV87, MN86, Oel88, PY86b, Pit81a, Ric82, Saa87, Sch86, Sch87a, She82a, SW82, Sil80, Ste86a, Tak82, Tay89, Uch82b, VN82a, VN82b, Yam84, Zak85, Ber89b, Bro80a]. Processes [AGKdV89, AB86, AS87, Ale84, ADG^{+}84, Ale86, AP86, Ale87a, Ale87b, AD87, Arr81, Arr85, BD88, BS87a, Bar88, BC83b, BP84a, BP84b, BP85, BC86, Bas88, Ber74, Ber80a, Ber80c, Ber82b, Ber83c, Ber88c, BL88, Bir88b, BS81, Bro80b, Bro81, BR88, CM80, CK83a, CL89a, Che81a, Chi82, CC89, CS85b, Coh85, CD81, CCHM86b, CHS87, CH88, Cuz78, Cuz87, DeB87, De 84, De 80, DPSW86, DS89, DH81, Deu88, Don87a, Don87b, Dur80, DG83, Dyn89, EM88a, EM88b, Ell80b, Eri81, Fal84, FP81, Fra85, Gae88, GS80, GM83a, GZ84, GZ86, GZ87, GW82, Glo80, Glo81a, Glo81b, GG86b, GH81, GKZ81, Gor88, Gre80, GF87, Gzy80, Hen83, HLS80, HLS87, Hol82, Hor84, Hu88, Hug85, Hur85, IM89, JP88, JN85, Jan85, Kal88, Kar84, Kas85, Kas88, KKR87]. Processes [Key87, Kle89a, KLS82, KS87, Kuc82, Kur80b, Kus85, LRS89, LR88, LT89, LR84, Lig80, Lin80, Lin85, LS87b, Mac81, Man87, Mas89b, Maz88, McC80a, McC80b, MR85, MN88, Mia88, MNS89, MP87, Mou89a, NN87a, NN87b, NP88, OV85, Oel84, OP88, Pin85d, Pin85a, Pin85b, Pin86, Pin88, Pit81b, Pou84, Pru81b, PT83, Rev82, RS88, Roo86, Ros87, Ros88, Ros80, RS80, Rup82, Sal87, Sam88, Sch85a, Sch87b, Sie83, Ste86b, Stu82b, Stu84, Tal87c, Tal88a, Ust84, Ver85, Web88, Yam85, YM85, Zee81, Zei89, dHR89, BL97, Ber80b, Ber84, SP81]. Processus [Ber89b, Bro80a]. Product [Car88, Rei81]. Products [Bou87, CN84]. Professors [ADG^{+}84]. Programming [FP81]. Projection [Csi84]. Projections [Dyk85, Sch85b]. Proof [AC88b, Bec81, CS85b, Cox84, Dur82, Sch87b, dA83a]. Propagation [CN88, Wal82]. Properties [AC88a, Bro80b, Bro81, BZ81, CM80, Cla87, Hol88, Kau89, Ken89, NN87a, Nol88, Nol92, Oka85a, Ver85, vPvS85, vZ82]. Property [CL89b, Csi84, EW89, FDS88, Kal87, Ken87, Rog81, Sch89, Sha88, Wen80b, Zam84a]. Prophet [Kla89, KS87]. Prophets [Ken85]. Proportion [AW89, Uch82a]. Publications [Ano86j, Ano89k]. Pure [BC86]. Purely [Ast81].

Quadratic [CRW85, FT80a, FT85, FI87, KW87, Nua84, San88]. Quantile [CCHM86b, CH88, EM88b, FR89, Mas84]. Quantiles [Rus88]. Quantities [BR84b]. Quasi [MSS82, Pen89]. Quasi-Every [Pen89]. Quasi-Superadditive [MSS82]. Question [Bra85]. Questions [GFZ86]. Queue [BM89a, Ros83]. Queues [Kip86, RN80].

R [Gae88]. Radial [Ken87]. Radius [Gem86]. Radon [HHK83]. Radonifying [CRW85]. Ramsey [BK83]. Random [Aar86, AF84, Ald83, And88c, Arr81, Arr83b, AH85, Bél89, BR84a, Ber83a, BH89, BF85, Bez87a, Bol82b, Bol89, Bou87, BK83, BW85, Cac82, CRW85, CN88, Cer86, CS85a, CS87, Che80, Chi85, CDG87, Cog80, CN84, Coh89, CK83b, CG84, Cox89, Dal80, DH84, Dav83, DR85, DH82, DFG89, De 81, DP82, DDP86, Dyn88b, Ein87, Ein89, Ell84, Ete80, Far86, Far87, FT80a, Fil88, FDS88, Fla82, FOW85, Fol84, FT85, GOV87, Gem80, Gem86, GS85, GH85, GG86b, GP83a, Gre85, Gri82, Gri83b, Gri83c, Gri86, Gri87, Gut80, Gut83, GJ86, HK80, HKS87, HB87, Hal80b, Hal83a, Hal88a, HHL81, Han89, Han91, HR83b, Her84b, HK82, Hol83, Hol88, Hol81a, Hu88, JP84, JP87a, JP87b, Jak86, Jan83a, Jan88, JDPP83, Joh85, JS89, KV83, Kal81a, KO88a]. Random [Kas88, KM87, Ken85, Ken89, Key84, Key87, Kin80, Kla81, Kla83, KT87, KS86, KS88, KP80, KW87, LW82, LT89, LW89, Llo84, Mar84, Mat82, MT86, McC80b, Mor83, Naw82, NK83, New87, Ney83, Nor84, O’B80, Oel88, Pel82, Pel87, Pem88, Phi80, Phi86, Pin83, Pit83, Pit85, Pru81b, PR83a, PR85, Rei82b, Rei86, Rit81, Ros84a, Ros86b, RS86b, Rus88, RS82, SC84, SS81, Sam84, Ser82, Ser86, Ser88, She81, Sie88, Sil83, Sla88, Sla89, Sol84, Sta84, Ste82, Tal87a, Tal89, TC83, Tom80, Ulb81, Wat84, Wat85, Woo83, Zac83, Zah88, Zha86, dA80, dA82, vZ82]. Randomized [Mil85]. Randomly [Ber81, DM84, Mad86, Sol80, TC83]. Randomness [Joe87]. Range [And82, AK84, BG81a, CI78, CI80, Ete80, FT85, Imh85, Lig80, Way80]. Rank [Doe82, SRP85]. Ranks [Che85]. Rapid [Hol85]. Rapidly [Fal87]. Rare [Ber83c]. Rate [Bir88b, BS80b, Bra89, Cox82, DDL86, FK86, Got86, Gri83b, Hae84, HJ88, Hae88, Hal80a, Hal82a, Hal82c, Hal83b, Hal84, Hal88a, Hal88b, Kla84b, MR87, Str86b, dHR89]. Rates [Bol82a, BG80, DS89, Eri80, Fil83, GK89a, Gri82, Gut80, HH81, Mar82, Ome88, RY89, RT84, RT89a, Ste88, Zab80]. Ratio [Iye86, MO82, Roo87]. Raw [Glo81a, Glo81b]. Reaction [Fre85]. Reaction-Diffusion [Fre85]. Reactions [Kot86]. Real [Far86, Far87, KM88, Kee82, She81]. Rearrangement [JS89, Rüs81]. Records [GR89]. Rectangular [Ter83]. Recurrence [Arr83b, Bha78, Bha80, Cer86, Key84, Kli87, Pin87, Pin88, Pin90, Wil85]. Recurrent [CI78, CI80, JP84, Jan85, Lee89, Pin85d, Way80]. Recursive [Dup88]. Refinement [MV87]. Reflected [HR81b, HLS85, HW87, MR85, Wil85]. Reflection [LR86c, CMEM80]. Réflexion [CMEM80]. Regarding [Mat82]. Regeneration [Key87]. Regenerative [FT88, Kal81b, KM88, Lal86]. Region [HLS85, Pin85d]. Regions [Yan86]. Register [Zam83]. Regression [LW82, Smy80]. Regular [Fad85, Lee85a, Pit81b, Ram81, Tom86, Bro80a]. Regularity [Bra85, Fra85]. Regularized [Dyn88a]. Regularly [CM86, DR85, HM87]. Regulières [Bro80a]. Reinforced [Pem88]. Related [BP83a, Fil83, HKP80, Mas89b, Pit83, Rev82, Shi83, Wri81]. Relation [NZ89, vPvS85]. Relations [Jur85]. Relative [Kie84, Kla80]. Reliability [AL82, BP83b]. Remain [Pin85d]. Remainder [Car83, Eng81]. Remark [Ber83b]. Remarks [Uhr84]. Renewal [AA87, Bro80b, Bro81, BW86, Car83, CHS87, Dal80, DS89, EMO84, Gru87, HW82, Hor84, Jan83b, Kee82, Lal84, Ney82, SP81, Sen82, Sol80, Ste86b, Zha88, Zha89]. Renormalization [Ros86a]. Rényi [DD87, Lyn83, Mas89a, AW89, CS81, DDL86, Ste86b]. Repeated [Li80]. Replacement [Ney82]. Representation [ARS87, Coh81, De 84, JS87, Lal86, Ros85, Sol84, Str86a, Wat85]. Representations [Alp83, Blo85, FOW85, Hai85, Kur80b]. Rescalings [Arr81]. Residence [Gol84]. Respect [Eme82, GM83a, KS88, PR83b, RW86, San88, Zak85]. Result [Arr85, Ber82a, Dur82, McC80a, Mic82, Tak81]. Results [Ein89, GKZ81, GJ86, HHL81, HR83a, HR83b, HR89, KZ83, Le 88, MS89, Mou89b, Pro80, Rei82a, Rei82b, Tal85, dAK83]. Return [Cog85]. Reversal [HP86, JP88, MNS89, Mit83c, Sha80]. Reverse [Ell86, Lee85a]. Reversed [Wu82]. Reversible [DPSW86, GL82, Lig83b, Lig89, Lyo83]. Revesz [Sha89]. Review [Don87b, Gae88, Sin88]. Revisited [Bai89, Hwa80]. Revival [Sie83]. Richardson [DL81]. Riemannian [Hsu89]. Right [Kuc89, Sal87]. Robbins [Kla85b, Rup82]. Role [BR84b]. Root [GP83a]. Roots [Far86, Far87, She81]. Rosen [Hal81]. Rotated [HKP80]. Rotational [Alp83, Hai85]. Rotationally [Mar86]. Rows [Sam84]. Ruin [Shi83]. Rule [HK82, Lee85a]. Rules [HP83, Mar82, Mar84, SC84]. Runs [Jan84, Pit81a]. RWRE [Pem88].

Salesman [RT89b]. Sample [Ale86, CR84, CS85b, DR84, FK86, HW82, Lai88, Mil81, RW86, SS81, Tal88a, Tuc80, Ver85, dHR82]. Samples [Goo88, Rob77, Rob83, Ste82]. Sampling [Hur85, Kal88, Kur80a, Lor81, Ney82, WW81]. Sandwich [AC88b]. Sanov [Csi84]. Satisfying [Her83]. Saucisse [Le 86]. Sausage [Le 88, Le 86]. Scale [Zak85]. Scaled [ENS87]. Scaled-Mins [ENS87]. Scaling [BW85, Sla89]. Scan [Cre80]. Sceneries [Bol89]. Scenery [Hol88]. Scheme [JJ86]. Schemes [Fin82b, KM87]. Score [JS87]. Scores [Qui80, QR84]. Second [Jen85, JS87, LR86b]. Second-Order [Jen85, JS87, LR86b]. Secretary [Lor81]. Sections [Gor88]. Sector [Gut83]. Seen [Fer86]. Séjour [Gol84]. Selected [Sol80]. Selecting [KR81]. Selection [BS87b, SS81]. Self [Bar81b, Dyn88a, Dyn88b, Fre81, GK89b, Law89, MR87, OV85, Pen89, Ros84a, Sla89, Ver85]. Self-Avoiding [Bar81b, Fre81, Law89, Sla89]. Self-Intersection [Dyn88a, Dyn88b]. Self-Intersections [Pen89, Ros84a]. Self-Normalized [GK89b]. Self-Similar [MR87, OV85, Ver85]. Selfdecomposable [Jur82, Jur85]. Semi [Luc84]. Semi-stable [Luc84]. Semicircle [BY88a]. Semigroup [DP86, Hsu89]. Semigroups [Beq89, BZ81, Cer86, LR84, Oka85a]. Semiinvariants [Jan88]. Semimartingale [Ken87]. Semimartingales [Avr88, Bic81, Eme82, Kar82, Pro85a, Pro85b]. Seminorms [BS86]. Semipolar [Kan89]. Sense [Ram81]. Separability [Ros86b]. Sequence [Ale89a, Ale89b, AW85, Ath88, Car86, Deh86, Li80, SS81]. Sequences [Avr88, Bir88a, Bra88, Bur81, CS85a, Che85, DDP86, DF80b, Eag81, Han88, Heb80, Her83, Her84a, Her84b, Jan84, Kal88, KO88a, Lin85, Mor83, NW81, O’B87, Pel82, Pel85, Pel87, WRL82, Woo85, Zam83, Zam84a]. Sequential [LL86, SS81]. Series [Blo85, Che80, Hol83, Hwa87, Kip86, Kla85b, KW87, LW82, Smi82, dA80, dlC89]. Series-Parallel [Smi82]. Set [Ale86, AP86, Ale89a, BP85, DL81, DL88, DS88b, DST89, FDS88, GG86a, GG86b, GR89, Hur85, Kue81, LV83, Nor84, PR83b, dAK83]. Set-Indexed [Ale86, AP86, GG86a, GG86b, BP85]. Set-Valued [PR83b]. Sets [Adl78, Adl80, Ale89b, AH85, BP84a, BP84b, Ber83c, Bla81, BN81, Cog85, Dyk85, FT88, GH85, GZ87, Hal83b, Jan85, Kal81b, Kan89, Kur80a, PR83a, Ros86b, RS82, Tal88b, WW81, Zah88]. Several [Jan87, Sch87b]. Sex [Asm80]. Shannon [AC88b, Bar85]. Shape [DL81, Gaw84]. Sharing [Smi82]. Sharp [BK89, Bur84, Bur87b, CK83b, DS89, Kla84a, RT89b].

Sheet [Adl78, CCFR89, Eps89, Mou89b, Wal82, Adl80]. Shepp [DDL86, DD87, Lyn83]. Shift [Zam83]. Shifted [DH84]. Shock [Ros81]. Shorack [Gae88]. Shuffles [FOW85]. Sided [Atk86, Gaw88, Mar82, Pru81a]. Siegmund [Kla85b]. Sign [Tal88b]. Sign-Invariant [Tal88b]. Signals [DM84]. Similar [MR87, OV85, Ver85]. Simple [Arr83a, Bec81, BF87, Cer86, Doe82, FP81, Fer86, Kie84, Kip86, Lyo83, Saa87, VN82b]. Simplex [Mat82]. Single [Ber81, Mou89a, Way80]. Singular [Bez87b, Kar82, KW88, Oka85a, Pin83]. Singularities [CN88, Wal82]. Singularity [Ros88]. Site [Arr83b, GKR88]. Sites [EG87]. Size [Ber83a, HW82, Kle89a, McC85]. Sizes [RS80]. Skew [HS81]. Slepian [JDPP83]. Sliding [Dun80, GOD80]. Sliding-Block [Dun80, GOD80]. SLLN [Kor84]. Sloping [Esc87]. Slow [DP85]. Slowly [Kas82]. Small [Bez87a, Cog85, CF86, GOV87, Gri83c, Hij84, Kla83, Lia88, Pin89, dA83b]. Smallest [Sil85]. Smooth [EK89, Hei84]. Smoothed [FR89]. Smoothing [And85, Ell86, Poo82]. Smoothness [BR82, CHM89, Eri81]. Sojourn [Uch82a]. Sojourns [Ber74, Ber80a, Ber82b, Ber83c, Ber88c, Ber80b, Ber84]. Solution [Van84]. Solutions [Heu86, MOT85, Pro85a, Rei83]. Some [ADG^{+}84, Asm80, BSS82, Bol82a, BG80, Bro80b, Bro83c, CM80, CS87, Chi82, Cla87, CD81, Cox88, Dal88, Dup88, Eps89, Fla82, GZ84, GKZ81, GFZ86, HR83a, HR83b, HR89, HLS80, Hol81b, KZ83, LT89, Lyn83, Mas82b, Rei82a, Rei82b, Sch85a, Sha80, Sie88, Tal85, Uhr84, VN82b, Wu82, Zak81, dA82, dAK83]. Sources [GOD80, Kie80b]. Sous [Der83]. Sous-Additif [Der83]. Space [Ale89a, Ale89b, BP85, Car88, Ein89, Ete80, GKZ81, Hal85b, Lal84, Oka85b, Pen89, PR83a, RT84, RT89a, Ros85, Tal89, UZ89, Van84, Zac83]. Spaces [AGMZ81, AH85, Bol84, Bur81, BZ81, CM89, Got81, Got86, Hei84, Jan83a, JS89, Kor84, KW87, LT88, LR84, RR80, Ust84, Web88]. Spacing [Hol81a]. Spacings [Deh82, Deh84, Deh86, Dev81, Dev82, EvZ88, Hal82b, Jan87, Pyk80]. Spanning [Ste88]. Spatial [MN86, Uch82b]. Specialized [Bro80b, Bro81]. Spectral [De 84, Fal84, Gem86, Ken82, Pin86, Ros84b]. Speed [DPSW86]. Speeds [Swe80, Way80]. Sphere [Ber80c, Gol84, Gol84]. Spheres [Hal85b, Lia88, Mat88a, Wen80a]. Spherical [Gor88]. Sphericity [Ber86]. Spin [GG82, Gra86]. Spitzer [ARS87, Dur82, GL82, Hal81]. Splicing [KR83]. SPLIFs [Bla80]. Splitting [BD86, Kal81b]. Spreading [Kal88]. Square [GP83a, HKP80, MO82]. Squared [Hal83a]. Squares [MO82]. Stability [Bou87, CN84, Ein89, FT80a, GG82, Tom86]. Stable [BD88, BC83b, BS86, CM80, CRW85, CCHM86a, Dav83, DFG89, Don87a, Gaw84, Gaw88, GH85, HHM82, HM81, HMV83, HJV86, KS88, LWZ81, Luc84, Mit81, Mit82, Mit83b, Nol88, Nol92, Ros88, RW86, Ryz86, SS89, Tal88a, dHR89, De 84, Mit83b]. Standard [RR80]. Started [DM84]. State [Dup88, Ell86, Ell80b, Lal84, Zac83]. States [Gan89, Lal86]. Stationarity [Zam83]. Stationary [AP89, Ari88, BM89a, Ber83a, BH89, Ber80a, Ber82b, Ber83c, Bol82b, Bra88, Car86, Che85, Fal84, FT88, Gan89, Get88, GK80, GF87, HLS85, HW87, Heb80, Her83, Kie80a, Kie80b, KLS82, McC80a, McC80b, Naw82, OV85, O’B87, OP88, Pin83, Pou84, RS88, RS80, Sam84, SP81, Sch87a, Sol84, Ste86a, Ver85, WRL82, Zam84a, Zam86]. Statistic [Car86, Cre80, Hal87]. Statistical [BM87, BDS89, Don87b, ER82, ER83, RS80]. Statistics [Ber82a, BP83a, CR84, Che85, Coo85, Dev81, Doe82, FR88, GS81, GP83b, Hal82b, Hel81b, JJ86, LWZ81, Mas82b, MS82b, Rei81, SRP85, Teu81, Van80, VV82, WRL82, Gae88]. Stay [Rit81]. Stein [AGG89, Tak81]. Stochastic [Alp83, AD87, BH86, BM82, Ber83b, Ber88b, Bic81, Bou87, BR84b, Bur87b, Car88, Coh81, DR84, Eme82, GM83a, Hai85, Haj82, Heu86, Hol85, Hu88, KS84, KKR87, Kle89b, KS88, KW87, LR88, LR84, MP80, Mia88, Now85, Oel84, Pou84, Pro85a, Pro86, Rei83, RT89b, Roo80, RW86, Str86b, Tud89, Wat85, Web88, Yan86, CMEM80]. Stochastically [CS85b, Pru83]. Stochastiques [CMEM80]. Stop [HK82, HP83, SC84]. Stopped [Esc87, GJ86, MP80, Pin86, RS86a]. Stopping [Ald89, cCS80, DeB87, Fin82b, Irl81, Ken85, Lor81, Mar82, Mar84, Mil85]. Storage [Yam85]. Strassen [Kon82]. Strategies [LV83, Now85]. Stratonovich [Haj82, NZ89]. Strength [Smi82]. Strict [Bur87b]. Strictly [KS88]. Strings [Zam83]. Strong [Bar85, CS81, Deh82, Deh84, Ebe86, Ein87, EvZ88, EM88b, Ein89, Heu86, Hor84, Imk86, Kal87, Mas89a, Mas89b, Ros84b, Ros80, Ste86b, AW89, BP84b, CZ86, De 81, Fil83, Gut83, HLS80, Hil82, Mas82b, McC80a, PR83a, PR83b, Rus88, Van80]. Strongly [Her83, Yam82]. Structure [Jur82, MSS82, Tal85, Tal88b]. Structures [JD83]. Student [Hal87]. Study [Li80]. Subadditive [GS80, Lig85, Ste81, Der83]. Subcritical [Sch87a, VN82a]. Subdivision [Pyk80]. Subfiltration [Hoo84]. Subgroups [BH83]. Subject [GOV87]. Submartingales [WW81]. Subordinated [Gru87]. Subordinators [FT88, JP87a, PV81]. Subsequences [Gut85]. Subsequential [GP89]. Subsets [Bro87]. Subspaces [Key88]. Suffice [AGG89]. Sufficient [AK80, BY88b, Bar88, Beq89, BW86, CD81, Fad85, GH81, Pin88]. Sum [AP86, BP84a, BP84b, CK83b, GZ87, Hal80b, Hal83a, MO82, MSS82, Now85, Pru87, Tal89]. Summand [Pru87]. Summands [Tak81]. Summation [Eri81]. Sums [BH89, Ber80a, BB83, CS81, CM86, CHM88, Dav83, Ein87, Ein89, Fil88, GG86b, Gri86, GP89, Gut80, HM87, HT83, HK80, HKS87, HR83b, Hil82, Jak86, Jan88, JS89, Kla81, Kla89, KP80, Mas82a, Phi80, Phi86, Pru83, QR84, Rei82b, Rei86, Roo87, RS86b, Sam84, Ser86, Ser88, Sla88, Smy80, Sta84, TC83, Tom80, Ulb81]. Superadditive [MSS82]. Supercritical [Coh85, Coh89, DG83, GP87, Kle89a, Kuc82, Kuc89, VN82b]. Supports [Isc88, Isc89]. Suprema [AB86, AS87]. Supremum [Adl84, Dar83, HK82]. Sure [BY88b, BD88, DP82, Fol84, Heb80, KP80, Muk84, Rup82, Stu86, TC83]. Surely [Hil82]. Survival [BG81a, Bra89]. Switched [Kel87]. Symmetric [And88b, Arr83a, Arr85, BC83b, CM80, Hen83, Mar86, MT86, Mit81, Mit83b, MS82b, RS86b, SS89, Tuc80, vE86]. Symmetry [HHM82, HJV86, Luc84].

Synchronization [GOD80]. Synonymity [Hoo84]. System [Arr83a, Car88, DFG89, Deu88, Smi82]. Systèmes [CMEM80]. Systems [BP83b, Bou87, BG80, BG89, Bra89, BR84b, Fri87, GOV87, Gor83, Gor84, GG82, Gra86, GL82, KM88, Lee88, Lee89, LR84, Lig83a, Lig83b, Lig89, Spi81, CMEM80].

Tactics [Mil85]. Tagged [Arr83a, Fer86, Saa87]. Tail [AB86, AS87, BH89, BK89, DR85, EG81, HM87, Hol83, JP87a, SS89, Sie88]. Tails [EM88a, Roo87]. Taken [ESY83]. Tanaka [Ros86a]. Tauberian [Bin81]. Taylor [Hwa87]. Temps [Gol84, Str86a]. Term [Car83, Eng81, Fin82a]. Terms [Ast81, BR89, MS80b]. Test [BB83, Gri83b]. th [CM80]. Their [CN84, LR84, DeB88, Don87b, Pin85a]. Théorême [RE83]. Theorem [AC88b, ARS87, AA87, AGMZ81, Bai89, Bar87, Bar85, Bec81, Bor85, BK83, Che81a, CC89, Coh81, Csi84, DF80a, Eag81, EMO84, Fit87, Gem80, GM83b, GP83a, HB87, Hal81, Hel81b, Hu88, Hur85, Kas82, Kur80a, LS87a, Lig85, Lyn83, McC87, Oka81, Oka85b, RN80, RS88, Roo87, RT81, Saa87, Sat81, Sch87b, Sen82, Sol84, Tud89, VN82a, Wei85, Woo83, Woo85, Wri80, Yam85, Yu81, Zam86, ZZ84, dlC89, Bro80a, Der83, RE83, AP86, Ale87a, AD87, Bar86, BP84a, Bha85, Bin81, Bir88b, Bol82b, Bol89, Bra88, Che88, Deu88, Eri81, GM83a, Got86, Hae84, HJ88, Hae88, HK81, Hal80a, HH81, Hal82a, Hal83b, Hal84, Hal88a, Han89, Han91, Her83, Her84a, Her84b, KT87, Kot86, Kuc89, Kur81, Kur84, LR86b, LR87, Oss87, Pel87]. Theorem [Pol82, RT84, RT89a, Sch86, Swe80, dA83b]. Theorème [Der83, Bro80a]. Theorems [ADG^{+}84, AT87, BB83, Bin81, CS85a, CS87, CD81, CG83, Cox88, Deh84, DDL86, Doe82, EM88b, Eps89, Ete80, Fla82, FT85, Fre85, GZ84, Gor83, Gor84, Got89, GJP84, Hal82b, Hol81b, HK84, Jak86, JJ86, Jan83a, LT89, Mil81, NN87a, NP88, Qui80, Res85, Ser82, Shi83, SW82, Teu81, Tom80, VN82b, Wu82, dHR82, Bol82a, CG84, DDP86, Dut89, Kip86, Sin88]. Theory [AL82, Bic81, CGG89, DR85, Doo89, Dyn89, HW82, Jan83b, KK85, Kee82, Kes86, Kes87a, Kes87b, KR81, Kie84, Lal84, LR88, Maz88, Ney82, Ome88, Roo86, Sal87, Sin89, Str86b, Swe85, Wie82, Zha88, Zha89, dlC89]. There [Bla80]. Thin [BN81]. Thinness [Bur87a]. Thouvenot [Kie84]. Three [CR85, Pro86, Qui80, Sal87, dlC89]. Three- [CR85]. Threshold [CCN87, SC84]. Tight [Dal80, Dal88]. Tightness [Ald89, Bou87, JO80, Mit83a]. Time [And85, Bar88, BC86, BC83c, CG83, CCFR83, CF86, CCFR89, Ell86, GHR84, Glo81a, Glo81b, GP83a, GF87, Gun80, Gzy80, Hae88, HP86, Hu88, Hug85, Imk86, JP88, JP84, JP87b, Kas88, KKR87, Ken82, Kur80b, LW82, Le 81, McC85, MN86, MNS89, Mit83c, Mou89a, Per81, Pin85d, PT83, Ros85, Ros88, San88, Sch85a, Sha80, Str86a, Tak82, Gol84]. Time-Changes [Glo81a]. Time-Periodic [Sch85a]. Times [Ald89, BS87a, BC83b, Bro83a, CK83a, Cog85, CG85, Cox88, Cox89, Cuz82, CD82, DeB87, Dyn88a, Fin82a, Jac84, Jen85, Kal81b, KS84, Kas82, Key87, Pit89, Pit81b, Ros87, Ros80, Sha88, Bro83b]. TLC [LT86]. Tori [Yan86]. Torus [Cox89]. Tossing [Rei82a, SW84]. Total [DH81, KQ89, Lee85b]. Total-Cost [DH81]. Traces [RS82]. Traffic [HK84, Zee81]. Transform [Che80, HHK83, Jaj87, Ros83]. Transformation [Alp85]. Transformations [Bec81, HK81, RE83, Sha80]. Transforms [BS80a, Bur84, KS87]. Transience [Key84, Lyo83, Pin87, Pin88, Pin90]. Transient [BM89a, Hen83, Jan85]. Transition [MPvW83, Pem88, Wei84]. Transitive [Pit83]. Transitivity [Irl81]. Translation [HHK83]. Transmission [Kie80b]. Transmitted [KM87]. Traveling [LS88, RT89b]. Travelling [LS89]. Treatment [Swe80]. Tree [SW84]. Trees [Pem88, Pit85, Ste88, Zac83]. Triangular [JJ86, Sam84, dA80, dA82]. Triangulation [Ste82]. Trimmed [CHM88, GP89, HKS87, Mal88]. Trivariate [KS84]. Triviality [BH89]. Tumour [BG81b]. Two [Alp85, AGG89, Asm80, Atk86, Beq89, Bol89, BG84, CW82, CK83b, CG86, Dur82, Dur84, Ell80b, Esc87, Ete80, FI87, GKR88, Gri87, Haj82, Imk86, Jac84, Lig83b, Mar82, Maz88, McC87, Ney86, Nua84, San88, Sch87a, Wei85, Zak81, Bro80a]. Two- [Ell80b, GKR88]. Two-Dimensional [Bol89]. Two-Fold [Alp85]. Two-Parameter [CW82, Ete80, Haj82, Imk86, Maz88, McC87, Nua84, San88, Zak81, FI87]. Two-Sex [Asm80]. Two-Sided [Atk86, Mar82]. Type [AL82, Ber88b, Che88, DD87, EMO84, Haj82, Oka85b, RN80, SRP85, VV82, Web88, Res85].

Uhlenbeck [Deu89, IM89, Tay89]. Ulrich [Sin88]. Unavoidable [BN81]. Unbiased [SW84]. Unconditional [Bur81]. Uncorrelatedness [JD83]. Undominated [PF86]. Unified [Bru84, BS87b]. Uniform [Adl78, Adl80, AP86, BP84a, Che81a, CH88, Deh82, Dev81, Dev82, EM88a, EvZ88, EW89, GJ86, Hen83, MV87, Mou89b, Sch89, SW82, Swe85]. Uniformity [Cre80]. Uniformly [CK83a, Hei84, Sch85b]. Unimodal [Uhr84, Yam82]. Unimodality [BK82, OR85, Ros80]. Unique [Heu86, KO88b]. Uniqueness [Ber89a, GKR88, Gan89, SP81]. Univariate [Deh84]. Universal [BM87, Dud87, KR81]. Universally [Now85]. Unknown [BK85, Bru84, BS87b]. Unscaled [Gor83, Gor84]. Unusual [Ale89b]. Upper [CW82, Cac82, CR85, HM87, Kon82, Sla88]. Urn [cCS80, Dut89, Fla82, HLS80, HLS87, Sen82, Zam84b]. Used [MS82a]. Using [Bro83c, HP83, McC80b, Tak81].

Valeurs [ESY83]. Value [Bar81a, Bur84, MO83, Ome88, Pic86, RS86a, Ric82, Roo86, She82a, Swe85, Web88]. Valued [De 81, DP82, Ein89, EG87, Fra85, GS80, Isc88, Isc89, IM89, Kar82, KP80, LW89, LM89, Phi80, Phi86, PR83a, PR83b, Tal89, dA80, dA82, dA88]. Values [CM86, HM87, Mas82a, O’B87, Ser82, ESY83]. Vapnik [Ale87a]. Varadhan [Pin85c]. Variable [Bro80a, Cac82, Che85]. Variables [BH89, BF85, CRW85, CS85a, CS87, CK83b, CG84, DH84, Dav83, DR85, DP82, DDP86, Ein89, FT80a, Fla82, FT85, GS85, GG86b, Gut80, Gut83, HKS87, Hal80b, Hal83a, HR83b, Her84b, HK82, Jak86, Jan83b, Jan88, JDPP83, Joh85, JS89, Ken85, Kla81, KT87, KP80, LW82, LW89, Mar84, MT86, Mor83, O’B80, Pel82, Pel87, Phi80, Pit82, PR85, Rei82b, Rei86, RS86b, Rus88, SC84, Ser82, Ser86, Ser88, Sil83, Sta84, Tal89, Tom80, Ulb81, Woo83, Phi86]. Variance [AP86, BK89, Bra88, Cac82, GG86b, GH81, Jen85]. Variates [Kla89]. Variation [Ber89b, FI87, KQ89, Nua84, San88, Tom86]. Variational [Pin88]. Variations [Avr88]. Varying [CM86, DR85, HM87, JN85, Zha89]. Vector [BZ81, DP82, GS80, dA80, dA88]. Vector-Valued [DP82, dA88]. Vectors [De 81, Ein87, Ell84, Gri86, HK80, RS86b, Sam84, Sla88, dA82, vZ82]. Venttsel [Nis81]. Version [Dro82, Kon82, Mas89a]. Via [BM89a, Bla81, JD83, LWZ81, Mas89b, RY89, Ros83]. Viewpoint [BM89b]. Viot [DH82]. Vitali [MS80b]. Voisinage [Gol84]. Volterra [Pro85b, Tud89]. Volume [Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano80m, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano82g, Ano82h, Ano82i, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano83i, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano85i, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano86k, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano87i, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e]. Volume [Ano88f, Ano88g, Ano88h, Ano88i, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Ano89l]. Volumes [MP87]. Voter [CG83, CG86, Cox88, Cox89, PS83].

Wald [Kla88]. Walk [Ald83, Ber83a, Dal80, Ete80, GP83a, Gri83b, Gri83c, Hol88, JP84, JP87b, Kal81a, Key84, Kin80, Kla83, Law89, Ney83, Pem88, Sla89, Zha86]. Walks [And88c, Arr81, Arr83b, Bar81b, Bél89, BR84a, Bol89, Cer86, Chi85, CDG87, Cox89, Dyn88b, Fre81, Gri82, GJ86, JP87a, Jan83a, KV83, Pru81b, Rit81, RS82]. Wandering [DH82]. Wasserstein [CM89]. Watson [Kle89a, Kus85, VN82a, VN82b]. Wave [CN88, Yan88]. Waves [LS88, LS89]. Weak [Avr88, BP85, Blu83, FR89, GG86b, Hal82c, Hel81a, Hwa80, Mar81, Mas84, McC80b, Phi80, vE86, Phi86]. Weakly [Ari88, DP82, Fal84, Her84b, Oel84, Tak81]. Wedge [Wil85]. Weighted [CCHM86b, CH88, DH84, EM88a, EvZ88, EM88b, GKZ81, MZ84, Mas81, Mas84, Rei82b, Rei86, Roo87, Sta84, Ste88, TC83, Ulb81]. Weiner [Mia88]. Wellner [Gae88]. Which [BC83a, BL88, Bur81, Hal83b, Hil82, Str86b, BS87a, BL97, FK86]. White [DM84, KK85]. Wiener [Bec81, Bro86, CCFR83, CF86, CCFR89, Don87a, GFZ86, HR83a, HR83b, HR89, Le 86, Le 88, Rev82, Ric82, She82a, Tak82, UZ89, Wat87, Zak85]. Wigner [BY88b]. Williams [BG81b]. Winding [Dur82]. Windings [Bél89].

Wintner [dA83a]. Wishart [Sil85]. Without [KT87, Ney82, Ros86b, Tom80]. Wolfowitz [Rup82]. Woodroofe [Yu81]. Words [KO88a]. Work [Doo89, Sin89].

XYZ [She82b].

Ylvisaker [McC80a].

Zero [And82, AK84, And88c, GOD80, Kla89, Now85, Way80]. Zero-Range [AK84, Way80]. Zero-Sum [Now85]. Zeros [BH80, Cuz78, Cuz87, New87]. Zinn [ADG^{+}84].

References

Anderson:1987:RTI

[AA87] Kevin K. Anderson and Krishna B. Athreya. A renewal theorem in the infinite mean case. Annals of Probability, 15(1):388–393, January 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992277.

Aaronson:1986:RE

[Aar86] Jon Aaronson. Random f-expansions. Annals of Probability, 14 (3):1037–1057, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992457.

Adler:1986:TBS

[AB86] Robert J. Adler and Lawrence D. Brown. Tail behaviour for suprema of empirical processes. Annals of Probability, 14(1):1–30, January 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992616.

Algoet:1988:AOA

[AC88a] Paul H. Algoet and Thomas M. Cover. Asymptotic optimality and asymptotic equipartition properties of log-optimum investment. Annals of Probability, 16(2):876–898, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991793.

Algoet:1988:SPS

[AC88b] Paul H. Algoet and Thomas M. Cover. A sandwich proof of the Shannon–McMillan–Breiman theorem. Annals of Probability, 16 (2):899–909, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991794.

Andersen:1987:CLT

[AD87] N. T. Andersen and V. Dobric. The Central Limit Theorem for stochastic processes. Annals of Probability, 15(1):164–177, January 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992262.

Alexander:1984:SLT

[ADG^{+}84] Kenneth S. Alexander, R. M. Dudley, Peter Gaenssler, Walter Philipp, David Pollard, Ronald Pyke, and Winfried Stute. Some limit theorems for empirical processes: Discussion of the paper of Professors Gine and Zinn. Annals of Probability, 12(4):990–998, November 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993139.

Adler:1978:UDL

[Adl78] Robert J. Adler. The uniform dimension of the level sets of a Brownian sheet. Annals of Probability, 6(3):509–515, June 1978. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995535. See

correction [Adl80].

Adler:1980:CNC

[Adl80] Robert J. Adler. Correction notes: Correction to “The Uniform Dimension of the Level Sets of a Brownian Sheet”. Annals of Probability, 8(5):1001–1002, October 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994630. See [Adl78].

Adler:1984:SPG

[Adl84] Robert J. Adler. The supremum of a particular Gaussian field. Annals of Probability, 12(2):436–444, May 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993299.

Adler:1984:CRM

[AF84] Robert J. Adler and Paul D. Feigin. On the cadlaguity of random measures. Annals of Probability, 12(2):615–630, May 1984. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993309.

Arratia:1989:TMS

[AGG89] R. Arratia, L. Goldstein, and L. Gordon. Two moments suffice for Poisson approximations: The Chen–Stein method. Annals of Probability, 17(1):9–25, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991491.

Aaronson:1989:ACC

[AGKdV89] Jon Aaronson, David Gilat, Michael Keane, and Vincent de Valk. An algebraic construction of a class of one-dependent processes. Annals of Probability, 17(1):128–143, January 1989. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991499.

Araujo:1981:ALT

[AGMZ81] Aloisio Araujo, Evarist Gine, V. Mandrekar, and Joel Zinn. On the accompanying laws theorem in Banach spaces. Annals of Probability, 9(2):202–210, April 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994462.

Artstein:1985:CLL

[AH85] Zvi Artstein and Jennie C. Hansen. Convexification in limit laws of random sets in Banach spaces. Annals of Probability, 13(1): 307–309, February 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993083.

Asmussen:1980:NSC

[AK80] Soren Asmussen and Thomas G. Kurtz. Necessary and sufficient conditions for complete convergence in the law of large numbers. Annals of Probability, 8(1):176–182, February 1980. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994835.

Andjel:1984:DHE

[AK84] Enrique D. Andjel and Claude Kipnis. Derivation of the hydrodynamical equation for the zero-range interaction process. Annals of Probability, 12(2):325–334, May 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993293.

Assaf:1982:CPT

[AL82] David Assaf and Benny Levikson. Closure of phase type distributions under operations arising in reliability theory. Annals of Probability, 10(1):265–269, February 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993932.

Aldous:1983:MAR

[Ald83] David Aldous. Minimization algorithms and random walk on the d-cube. Annals of Probability, 11(2):403–413, May 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993605.

Aldous:1989:STT

[Ald89] David Aldous. Stopping times and tightness. II. Annals of Probability, 17(2):586–595, April 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991417.

Alexander:1984:PIE

[Ale84] Kenneth S. Alexander. Probability inequalities for empirical processes and a law of the iterated logarithm. Annals of Probability, 12(4):1041–1067, November 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:/

/projecteuclid.org/euclid.aop/1176993141. See correction

[Ale87b].

Alexander:1986:SMS

[Ale86] Kenneth S. Alexander. Sample moduli for set-indexed Gaussian processes. Annals of Probability, 14(2):598–611, April 1986. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992533.

Alexander:1987:CLT

[Ale87a] Kenneth S. Alexander. The Central Limit Theorem for empirical processes on Vapnik–Cervonenkis classes. Annals of Probability, 15(1):178–203, January 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992263.

Alexander:1987:CPI

[Ale87b] Kenneth S. Alexander. Correction: Probability inequalities for empirical processes and a law of the iterated logarithm. Annals of Probability, 15(1):428–430, January 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992281. See [Ale84].

Alexander:1989:CCS

[Ale89a] Kenneth S. Alexander. Characterization of the cluster set of the LIL sequence in Banach space. Annals of Probability, 17(2):737– 759, April 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991424.

Alexander:1989:UCS

[Ale89b] Kenneth S. Alexander. Unusual cluster sets for the LIL sequence in Banach space. Annals of Probability, 17(3):1170–1185, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991263.

Alpern:1983:RRS

[Alp83] Steve Alpern. Rotational representations of stochastic matrices. Annals of Probability, 11(3):789–794, August 1983. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993523.

Alpern:1985:CGM

[Alp85] Steven Alpern. Conjecture: In general a mixing transformation is not two-fold mixing. Annals of Probability, 13(1):310–313, February 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993084.

Andjel:1982:IMZ

[And82] Enrique Daniel Andjel. Invariant measures for the zero range process. Annals of Probability, 10(3):525–547, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993765.

Andjel:1985:IML

[And85] Enrique D. Andjel. Invariant measures and long time behaviour of the smoothing process. Annals of Probability, 13(1):62–71, February 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993066.

Andjel:1988:CPH

[And88a] Enrique D. Andjel. The contact process in high dimensions. Annals of Probability, 16(3):1174–1183, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991683.

Andjel:1988:CIS

[And88b] Enrique D. Andjel. A correlation inequality for the symmetric exclusion process. Annals of Probability, 16(2):717–721, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991782.

Andjel:1988:ZOL

[And88c] Enrique D. Andjel. A zero or one law for one dimensional random walks in random environments. Annals of Probability, 16 (2):722–729, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991783.

Anonymous:1980:MBPa

[Ano80a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 8, Number 1 (1980). Annals of Probability, 8(1):??, February 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994836.

Anonymous:1980:MBPb

[Ano80b] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 8, Number 2 (1980). Annals of Probability, 8(2):??, April 1980. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994790.

Anonymous:1980:MBPc DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994736.

Anonymous:1980:MBPd

[Ano80d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 8, Number 4 (1980). Annals of Probability, 8(4):??, August 1980. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994673.

Anonymous:1980:MBPe

[Ano80e] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 8, Number 5 (1980). Annals of Probability, 8(5):??, October 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994631.

Anonymous:1980:MBPf

[Ano80f] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 8, Number 6 (1980). Annals of Probability, 8(6):??, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994585.

Anonymous:1980:MFPa

[Ano80g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 8, Number 1 (1980). Annals of Probability, 8(1):??, February 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994823.

Anonymous:1980:MFPb

[Ano80h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 8, Number 2 (1980). Annals of Probability, 8(2):??, April 1980. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994770.

Anonymous:1980:MFPc

[Ano80i] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 8, Number 3 (1980). Annals of Probability, 8(3):??, June 1980. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994715.

Anonymous:1980:MFPd

[Ano80j] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 8, Number 4 (1980). Annals of Probability, 8(4):??, August 1980. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994657.

Anonymous:1980:MFPe

[Ano80k] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 8, Number 5 (1980). Annals of Probability, 8(5):??, October 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994617.

Anonymous:1980:MFPf

[Ano80l] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 8, Number 6 (1980). Annals of Probability, 8(6):??, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994564.

Anonymous:1980:VI

[Ano80m] Anonymous. Volume information. Annals of Probability, 8(6):??, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994563.

Anonymous:1981:MBPa

[Ano81a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 9, Number 1 (1981). Annals of Probability, 9(1):??, February 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994520.

Anonymous:1981:MBPb

[Ano81b] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 9, Number 2 (1981). Annals of Probability, 9(2):??, April 1981. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994480.

Anonymous:1981:MBPc DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994429.

Anonymous:1981:MBPd

[Ano81d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 9, Number 4 (1981). Annals of Probability, 9(4):??, August 1981. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994381.

Anonymous:1981:MBPe

[Ano81e] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 9, Number 5 (1981). Annals of Probability, 9(5):??, October 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994322.

Anonymous:1981:MBPf

[Ano81f] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 9, Number 6 (1981). Annals of Probability, 9(6):??, December 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994277.

Anonymous:1981:MFPa

[Ano81g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 9, Number 1 (1981). Annals of Probability, 9(1):??, February 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994507.

Anonymous:1981:MFPb

[Ano81h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 9, Number 2 (1981). Annals of Probability, 9(2):??, April 1981. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994458.

Anonymous:1981:MFPc

[Ano81i] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 9, Number 3 (1981). Annals of Probability, 9(3):??, June 1981. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994409.

Anonymous:1981:MFPd

[Ano81j] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 9, Number 4 (1981). Annals of Probability, 9(4):??, August 1981. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994358.

Anonymous:1981:MFPe

[Ano81k] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 9, Number 5 (1981). Annals of Probability, 9(5):??, October 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994304.

Anonymous:1981:MFPf

[Ano81l] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 9, Number 6 (1981). Annals of Probability, 9(6):??, December 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994263.

Anonymous:1981:VI

[Ano81m] Anonymous. Volume information. Annals of Probability, 9(6):??, December 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994262.

Anonymous:1982:MBPa

[Ano82a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 10, Number 1 (1982). Annals of Probability, 10(1):??, February 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993934.

Anonymous:1982:MBPb

[Ano82b] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 10, Number 2 (1982). Annals of Probability, 10(2):??, May 1982. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993875.

Anonymous:1982:MBPc

[Ano82c] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 10, Number 3 (1982). Annals of Probability, 10(3):??, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993804.

Anonymous:1982:MBPd

[Ano82d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 10, Number 4 (1982). Annals of Probability, 10(4):??, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993736.

Anonymous:1982:MFPa

[Ano82e] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 10, Number 1 (1982). Annals of Probability, 10(1):??, February 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993911.

Anonymous:1982:MFPb

[Ano82f] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 10, Number 2 (1982). Annals of Probability, 10(2):??, May 1982. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993856.

Anonymous:1982:MFPc

[Ano82g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 10, Number 3 (1982). Annals of Probability, 10(3):??, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993763.

Anonymous:1982:MFPd

[Ano82h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 10, Number 4 (1982). Annals of Probability, 10(4):??, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993710.

Anonymous:1982:VI

[Ano82i] Anonymous. Volume information. Annals of Probability, 10(4):??, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993709.

Anonymous:1983:MBPa

[Ano83a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 11, Number 1 (1983). Annals of Probability, 11(1):??, February 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993673.

Anonymous:1983:MBPb

[Ano83b] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 11, Number 2 (1983). Annals of Probability, 11(2):??, May 1983. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993613.

Anonymous:1983:MBPc

[Ano83c] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 11, Number 3 (1983). Annals of Probability, 11(3):??, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993535.

Anonymous:1983:MBPd

[Ano83d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 11, Number 4 (1983). Annals of Probability, 11(4):??, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993458.

Anonymous:1983:MFPa

[Ano83e] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 11, Number 1 (1983). Annals of Probability, 11(1):??, February 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993654.

Anonymous:1983:MFPb

[Ano83f] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 11, Number 2 (1983). Annals of Probability, 11(2):??, May 1983. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993593.

Anonymous:1983:MFPc

[Ano83g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 11, Number 3 (1983). Annals of Probability, 11(3):??, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993496.

Anonymous:1983:MFPd

[Ano83h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 11, Number 4 (1983). Annals of Probability, 11(4):??, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993434.

Anonymous:1983:VI

[Ano83i] Anonymous. Volume information. Annals of Probability, 11(4):??, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993433.

Anonymous:1984:MBPa

[Ano84a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 12, Number 1 (1984). Annals of Probability, 12(1):??, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993396.

Anonymous:1984:MBPb

[Ano84b] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 12, Number 2 (1984). Annals of Probability, 12(2):??, May 1984. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993313.

Anonymous:1984:MBPc

[Ano84c] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 12, Number 3 (1984). Annals of Probability, 12(3):??, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993245.

Anonymous:1984:MBPd

[Ano84d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 12, Number 4 (1984). Annals of Probability, 12(4):??, November 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993153.

Anonymous:1984:MFPa

[Ano84e] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 12, Number 1 (1984). Annals of Probability, 12(1):??, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993369.

Anonymous:1984:MFPb

[Ano84f] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 12, Number 2 (1984). Annals of Probability, 12(2):??, May 1984. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993290.

Anonymous:1984:MFPc

[Ano84g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 12, Number 3 (1984). Annals of Probability, 12(3):??, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993219.

Anonymous:1984:MFPd

[Ano84h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 12, Number 4 (1984). Annals of Probability, 12(4):??, November 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993137.

Anonymous:1984:VI

[Ano84i] Anonymous. Volume information. Annals of Probability, 12(4):??, November 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993136.

Anonymous:1985:MBPa

[Ano85a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 13, Number 1 (1985). Annals of Probability, 13(1):??, February 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993086.

Anonymous:1985:MBPb

[Ano85b] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 13, Number 2 (1985). Annals of Probability, 13(2):??, May 1985. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993017.

Anonymous:1985:MBPc

[Ano85c] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 13, Number 3 (1985). Annals of Probability, 13(3):??, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992928.

Anonymous:1985:MBPd

[Ano85d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 13, Number 4 (1985). Annals of Probability, 13(4):??, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992823.

Anonymous:1985:MFPa

[Ano85e] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 13, Number 1 (1985). Annals of Probability, 13(1):??, February 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993062.

Anonymous:1985:MFPb

[Ano85f] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 13, Number 2 (1985). Annals of Probability, 13(2):??, May 1985. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992993.

Anonymous:1985:MFPc

[Ano85g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 13, Number 3 (1985). Annals of Probability, 13(3):??, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992900.

Anonymous:1985:MFPd

[Ano85h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 13, Number 4 (1985). Annals of Probability, 13(4):??, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992797.

Anonymous:1985:VI

[Ano85i] Anonymous. Volume information. Annals of Probability, 13(4):??, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992796.

Anonymous:1986:MBPa

[Ano86a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 14, Number 1 (1986). Annals of Probability, 14(1):??, January 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992634.

Anonymous:1986:MBPb

[Ano86b] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 14, Number 2 (1986). Annals of Probability, 14(2):??, April 1986. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992543.

Anonymous:1986:MBPc

[Ano86c] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 14, Number 3 (1986). Annals of Probability, 14(3):??, July 1986. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992466.

Anonymous:1986:MBPd

[Ano86d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 14, Number 4 (1986). Annals of Probability, 14(4):??, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992385.

Anonymous:1986:MFPa

[Ano86e] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 14, Number 1 (1986). Annals of Probability, 14(1):??, January 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992615.

Anonymous:1986:MFPb

[Ano86f] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 14, Number 2 (1986). Annals of Probability, 14(2):??, April 1986. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992520.

Anonymous:1986:MFPc

[Ano86g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 14, Number 3 (1986). Annals of Probability, 14(3):??, July 1986. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992435.

Anonymous:1986:MFPd

[Ano86h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 14, Number 4 (1986). Annals of Probability, 14(4):??, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992354.

Anonymous:1986:PMKa

[Ano86i] Anonymous. [photograph]: Mark Kac. Annals of Probability, 14(4):1102, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992355.

Anonymous:1986:PMKb

[Ano86j] Anonymous. Publications of mark Kac. Annals of Probability, 14 (4):1149–1154, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992359.

Anonymous:1986:VI

[Ano86k] Anonymous. Volume information. Annals of Probability, 14(4):??, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992353.

Anonymous:1987:MBPa

[Ano87a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 15, Number 1 (1987). Annals of Probability, 15(1):??, January 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992282.

Anonymous:1987:MBPb

[Ano87b] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 15, Number 2 (1987). Annals of Probability, 15(2):??, April 1987. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992177.

Anonymous:1987:MBPc

[Ano87c] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 15, Number 3 (1987). Annals of Probability, 15(3):??, July 1987. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992098.

Anonymous:1987:MBPd

[Ano87d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 15, Number 4 (1987). Annals of Probability, 15(4):??, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991999.

Anonymous:1987:MFPa

[Ano87e] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 15, Number 1 (1987). Annals of Probability, 15(1):??, January 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992254.

Anonymous:1987:MFPb

[Ano87f] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 15, Number 2 (1987). Annals of Probability, 15(2):??, April 1987. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992154.

Anonymous:1987:MFPc

[Ano87g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 15, Number 3 (1987). Annals of Probability, 15(3):??, July 1987. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992068.

Anonymous:1987:MFPd

[Ano87h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 15, Number 4 (1987). Annals of Probability, 15(4):??, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991974.

Anonymous:1987:VI

[Ano87i] Anonymous. Volume information. Annals of Probability, 15(4):??, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991973.

Anonymous:1988:MBPa

[Ano88a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 16, Number 1 (1988). Annals of Probability, 16(1):??, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991915.

Anonymous:1988:MBPb

[Ano88b] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 16, Number 2 (1988). Annals of Probability, 16(2):??, April 1988. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991797.

Anonymous:1988:MBPc

[Ano88c] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 16, Number 3 (1988). Annals of Probability, 16(3):??, July 1988. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991698.

Anonymous:1988:MBPd

[Ano88d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 16, Number 4 (1988). Annals of Probability, 16(4):??, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991603.

Anonymous:1988:MFPa

[Ano88e] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 16, Number 1 (1988). Annals of Probability, 16(1):??, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991883.

Anonymous:1988:MFPb

[Ano88f] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 16, Number 2 (1988). Annals of Probability, 16(2):??, April 1988. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991766.

Anonymous:1988:MFPc

[Ano88g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 16, Number 3 (1988). Annals of Probability, 16(3):??, July 1988. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991669.

Anonymous:1988:MFPd

[Ano88h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 16, Number 4 (1988). Annals of Probability, 16(4):??, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991575.

Anonymous:1988:VI

[Ano88i] Anonymous. Volume information. Annals of Probability, 16(4):??, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991574.

Anonymous:1989:MBPa

[Ano89a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 17, Number 1 (1989). Annals of Probability, 17(1):??, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991521.

Anonymous:1989:MBPb

[Ano89b] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 17, Number 2 (1989). Annals of Probability, 17(2):??, April 1989. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991430.

Anonymous:1989:MBPc

[Ano89c] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 17, Number 3 (1989). Annals of Probability, 17(3):??, July 1989. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991270.

Anonymous:1989:MBPd

[Ano89d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 17, Number 4 (1989). Annals of Probability, 17(4):??, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991183.

Anonymous:1989:MFPa

[Ano89e] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 17, Number 1 (1989). Annals of Probability, 17(1):??, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991489.

Anonymous:1989:MFPb

[Ano89f] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 17, Number 2 (1989). Annals of Probability, 17(2):??, April 1989. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991407.

Anonymous:1989:MFPc

[Ano89g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 17, Number 3 (1989). Annals of Probability, 17(3):??, July 1989. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991244.

Anonymous:1989:MFPd

[Ano89h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 17, Number 4 (1989). Annals of Probability, 17(4):??, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991154.

Anonymous:1989:PKa

[Ano89i] Anonymous. [photograph]: A. N. Kolmogorov. Annals of Probability, 17(3):814, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991246.

Anonymous:1989:PKI

[Ano89j] Anonymous. Preface to the Kolmogorov issue. Annals of Probability, 17(3):??, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991245.

Anonymous:1989:PKb

[Ano89k] Anonymous. Publications of a. n. Kolmogorov. Annals of Probability, 17(3):945–964, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991252.

Anonymous:1989:VI

[Ano89l] Anonymous. Volume information. Annals of Probability, 17(4):??, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991153.

Alexander:1986:UCL

[AP86] Kenneth S. Alexander and Ronald Pyke. A uniform Central Limit Theorem for set-indexed partial-sum processes with finite variance. Annals of Probability, 14(2):582–597, April 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992532.

Alpern:1989:CSP

[AP89] S. Alpern and V. S. Prasad. Coding a stationary process to one with prescribed marginals. Annals of Probability, 17(4):1658–1663, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991180.

Arimoto:1988:AFP

[Ari88] Akio Arimoto. Approximation of the finite prediction for a weakly stationary process. Annals of Probability, 16(1):355–360, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991907.

Arratia:1981:LPP

[Arr81] Richard Arratia. Limiting point processes for rescalings of coalescing and annihilating random walks on Z^{d}. Annals of Probability, 9 (6):909–936, December 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994264.

Arratia:1983:MTP

[Arr83a] Richard Arratia. The motion of a tagged particle in the simple symmetric exclusion system on Z. Annals of Probability, 11 (2):362–373, May 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993602.

Arratia:1983:SRA

[Arr83b] Richard Arratia. Site recurrence for annihilating random walks on Z_{d}. Annals of Probability, 11(3):706–713, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993515.

Arratia:1985:SEP

[Arr85] Richard Arratia. Symmetric exclusion processes: A comparison inequality and a large deviation result. Annals of Probability, 13 (1):53–61, February 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993065.

Alzaid:1987:ESI

[ARS87] Abdulhamid A. Alzaid, C. Radhakrishna Rao, and D. N. Shanbhag. An extension of Spitzer’s integral representation theorem with an application. Annals of Probability, 15(3):1210–1216, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992092.

Adler:1987:TBS

[AS87] Robert J. Adler and Gennady Samorodnitsky. Tail behaviour for the suprema of Gaussian processes with applications to empirical processes. Annals of Probability, 15(4):1339–1351, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991980.

Asmussen:1980:STS

[Asm80] Soren Asmussen. On some two-sex population models. Annals of Probability, 8(4):727–744, August 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994662.

Astbury:1981:OCM

[Ast81] Kenneth A. Astbury. Order convergence of martingales in terms of countably additive and purely finitely additive martingales. Annals of Probability, 9(2):266–275, April 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994467.

Avram:1987:NLT

[AT87] Florin Avram and Murad S. Taqqu. Noncentral limit theorems and Appell polynomials. Annals of Probability, 15(2):767–775, April 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992170.

Athreya:1988:MSC

[Ath88] K. B. Athreya. On the maximum sequence in a critical branching process. Annals of Probability, 16(2):502–507, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991770.

Atkinson:1986:TSM

[Atk86] Bruce W. Atkinson. Two-sided Markov chains. Annals of Probability, 14(2):459–479, April 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992526.

Avram:1988:WCV

[Avr88] Florin Avram. Weak convergence of the variations, iterated integrals and Doleans–Dade exponentials of sequences of semimartingales. Annals of Probability, 16(1):246–250, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991898.

Arratia:1985:CPS

[AW85] Richard Arratia and Michael S. Waterman. Critical phenomena in sequence matching. Annals of Probability, 13(4):1236–1249, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992808.

Arratia:1989:ERS

[AW89] R. Arratia and M. S. Waterman. The Erdős–Rényi Strong Law for pattern matching with a given proportion of mismatches. Annals of Probability, 17(3):1152–1169, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991262.

Bai:1989:TFR

[Bai89] Z. D. Bai. A theorem of Feller revisited. Annals of Probability, 17 (1):385–395, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991518.

Barlow:1981:CMG

[Bar81a] M. T. Barlow. Construction of a martingale with given absolute value. Annals of Probability, 9(2):314–320, April 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994473.

Barrett:1981:PSA

[Bar81b] Wayne W. Barrett. [polymers as self-avoiding walks]: Discussion. Annals of Probability, 9(4):555–556, August 1981. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994360.

Barron:1985:SET

[Bar85] Andrew R. Barron. The strong ergodic theorem for densities: Generalized Shannon–McMillan–Breiman theorem. Annals of Probability, 13(4):1292–1303, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992813.

Barron:1986:ECL

[Bar86] Andrew R. Barron. Entropy and the Central Limit Theorem. Annals of Probability, 14(1):336–342, January 1986. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992632.

Barbour:1987:AEP

[Bar87] A. D. Barbour. Asymptotic expansions in the Poisson limit theorem. Annals of Probability, 15(2):748–766, April 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992169.

Barlow:1988:NSC

[Bar88] M. T. Barlow. Necessary and sufficient conditions for the continuity of local time of Lévy processes. Annals of Probability, 16(4): 1389–1427, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991576.

Bass:1988:PEM

[Bas88] Richard F. Bass. Probability estimates for multiparameter Brownian processes. Annals of Probability, 16(1):251–264, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991899.

Bickel:1983:SFN

[BB83] Peter J. Bickel and Leo Breiman. Sums of functions of nearest neighbor distances, moment bounds, limit theorems and a goodness of fit test. Annals of Probability, 11(1):185–214, February 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993668.

Bass:1983:BML

[BC83a] R. F. Bass and M. Cranston. Brownian motion with lower class moving boundaries which grow faster than t^{1/2}. Annals of Probability, 11(1):34–39, February 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993657.

Bass:1983:ETS

[BC83b] R. F. Bass and M. Cranston. Exit times for symmetric stable processes in R^{n}. Annals of Probability, 11(3):578–588, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993502.

Brown:1983:FPT

[BC83c] Mark Brown and Narasinga R. Chaganty. On the first passage time distribution for a class of Markov chains. Annals of Probability, 11(4):1000–1008, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993448.

Bass:1986:MCP

[BC86] R. F. Bass and M. Cranston. The Malliavin calculus for pure jump processes and applications to local time. Annals of Probability, 14(2):490–532, April 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992528.

Bellow:1980:ML

[BD80] A. Bellow and A. Dvoretzky. On martingales in the limit. Annals of Probability, 8(3):602–606, June 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994730.

Brennan:1986:SI

[BD86] Michael D. Brennan and Richard Durrett. Splitting intervals. Annals of Probability, 14(3):1024–1036, July 1986. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992456.

Balkema:1988:ASC

[BD88] A. A. Balkema and L. De Haan. Almost sure continuity of stable moving average processes with index less than one. Annals of Probability, 16(1):333–343, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991905.

Bramson:1989:SMC

[BDS89] M. Bramson, R. Durrett, and G. Swindle. Statistical mechanics of crabgrass. Annals of Probability, 17(2):444–481, April 1989. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991410.

Bass:1985:LLI

[BE85] R. F. Bass and K. B. Erickson. Local laws of the iterated logarithm for diffusions. Annals of Probability, 13(2):616–624, May 1985. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993014. See

correction [BE86].

Bass:1986:CLL

[BE86] R. F. Bass and K. B. Erickson. Correction: Local laws of the iterated logarithm for diffusions. Annals of Probability, 14(2):731, April 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992542. See [BE85].

Becker:1981:MGM

[Bec81] M. E. Becker. Multiparameter groups of measure-preserving transformations: A simple proof of Wiener’s ergodic theorem. Annals of Probability, 9(3):504–509, June 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994423.

Belisle:1989:WRW

[Bél89] Claude Bélisle. Windings of random walks. Annals of Probability, 17(4):1377–1402, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991160.

Bequillard:1989:SCT

[Beq89] A. L. Bequillard. A sufficient condition for two Markov semigroups to commute. Annals of Probability, 17(4):1478–1482, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991168.

Berman:1974:SEG

[Ber74] Simeon M. Berman. Sojourns and extremes of Gaussian processes. Annals of Probability, 2(6):999–1026, December 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996495. See cor-

rection [Ber80b].

Berman:1980:CPL

[Ber80a] Simeon M. Berman. A compound Poisson limit for stationary sums, and sojourns of Gaussian processes. Annals of Probability, 8(3):511–538, June 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994725.

Berman:1980:CNC

[Ber80b] Simeon M. Berman. Correction notes: Correction to “Sojourns and Extremes of Gaussian Processes”. Annals of Probability, 8(5):999, October 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994628. See [Ber74].

Berman:1980:IGP

[Ber80c] Simeon M. Berman. Isotropic Gaussian processes on the Hilbert sphere. Annals of Probability, 8(6):1093–1106, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994571.

Berbee:1981:CSP

[Ber81] Henry Berbee. On covering single points by randomly ordered intervals. Annals of Probability, 9(3):520–528, June 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994426.

Bergman:1982:GAI

[Ber82a] Bo Bergman. On a general asymptotic independence result in statistics. Annals of Probability, 10(3):831–837, August 1982. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993793.

Berman:1982:SES

[Ber82b] Simeon M. Berman. Sojourns and extremes of stationary processes. Annals of Probability, 10(1):1–46, February 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993912. See cor-

rection [Ber84].

Berbee:1983:BSP

[Ber83a] Henry Berbee. A bound on the size of point clusters of a random walk with stationary increments. Annals of Probability, 11 (2):414–418, May 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993606.

Berger:1983:RSF

[Ber83b] Marc A. Berger. A remark on stochastic fundamental matrices. Annals of Probability, 11(1):215–216, February 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993669.

Berman:1983:SSP

[Ber83c] Simeon M. Berman. Sojourns of stationary processes in rare sets. Annals of Probability, 11(4):847–866, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993436.

Berman:1984:CCB

[Ber84] Simeon M. Berman. Correction: Correction to “Sojourns and Extremes of Gaussian Processes”. Annals of Probability, 12(1):281, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993394. See [Ber82b].

Berk:1986:SNL

[Ber86] Robert H. Berk. Sphericity and the normal law. Annals of Probability, 14(2):696–701, April 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992538.

Berg:1988:CND

[Ber88a] Christian Berg. The cube of a normal distribution is indeterminate. Annals of Probability, 16(2):910–913, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991795.

Berger:1988:MTA

[Ber88b] Marc A. Berger. A Malliavin-type anticipative stochastic calculus. Annals of Probability, 16(1):231–245, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991897.

Berman:1988:ESD

[Ber88c] Simeon M. Berman. Extreme sojourns of diffusion processes. Annals of Probability, 16(1):361–374, January 1988. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991908.

Berbee:1989:UGM

[Ber89a] Henry Berbee. Uniqueness of Gibbs measures and absorption probabilities. Annals of Probability, 17(4):1416–1431, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991162.

Bertoin:1989:IPP

[Ber89b] Jean Bertoin. Sur une intégrale pour les processus à α-variation bornée. (French) [On an integral for the process of limited α variation]. Annals of Probability, 17(4):1521–1535, October 1989. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991171.

Bezuidenhout:1987:LDP

[Bez87a] Carol Bezuidenhout. A large deviations principle for small perturbations of random evolution equations. Annals of Probability, 15(2):646–658, April 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992163.

Bezuidenhout:1987:SPD

[Bez87b] Carol Bezuidenhout. Singular perturbations of degenerate diffusions. Annals of Probability, 15(3):1014–1043, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992078.

Bertoluzza:1985:MDR

[BF85] Carlo Bertoluzza and Bruno Forte. Mutual dependence of random variables and maximum discretized entropy. Annals of Probability, 13(2):630–637, May 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993016.

Benassi:1987:HLA

[BF87] Albert Benassi and Jean-Pierre Fouque. Hydrodynamical limit for the asymmetric simple exclusion process. Annals of Probability, 15(2):546–560, April 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992158.

Block:1988:MEH

[BF88] Henry W. Block and Zhaoben Fang. A multivariate extension of Hoeffding’s lemma. Annals of Probability, 16(4):1803–1820, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991598.

Bramson:1980:CDR

[BG80] Maury Bramson and David Griffeath. Clustering and dispersion rates for some interacting particle systems on Z. Annals of Probability, 8(2):183–213, April 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994771.

Bramson:1981:NSL

[BG81a] Maury Bramson and Lawrence Gray. A note on the survival of the long-range contact process. Annals of Probability, 9(5): 885–890, October 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994316.

Bramson:1981:WBT

[BG81b] Maury Bramson and David Griffeath. On the Williams–Bjerknes tumour growth model I. Annals of Probability, 9(2):173–185, April 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994459.

Brown:1984:EAT

[BG84] Mark Brown and Guangping Ge. Exponential approximations for two classes of aging distributions. Annals of Probability, 12 (3):869–875, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993235.

Bramson:1989:FFC

[BG89] Maury Bramson and David Griffeath. Flux and fixation in cyclic particle systems. Annals of Probability, 17(1):26–45, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991492.

Brockett:1980:ZDI

[BH80] Patrick L. Brockett and William N. Hudson. Zeros of the densities of infinitely divisible measures. Annals of Probability, 8 (2):400–403, April 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994789.

Byczkowski:1983:GMN

[BH83] T. Byczkowski and A. Hulanicki. Gaussian measure of normal subgroups. Annals of Probability, 11(3):685–691, August 1983. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993513.

Baxendale:1986:ISF

[BH86] Peter Baxendale and Theodore E. Harris. Isotropic stochastic flows. Annals of Probability, 14(4):1155–1179, October 1986. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992360.

Berbee:1989:TTS

[BH89] H. C. P. Berbee and W. Th. F. Den Hollander. Tail triviality for sums of stationary random variables. Annals of Probability, 17(4): 1635–1645, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991177.

Bhattacharya:1978:CRE

[Bha78] R. N. Bhattacharya. Criteria for recurrence and existence of invariant measures for multidimensional diffusions. Annals of Probability, 6(4):541–553, August 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:/

/projecteuclid.org/euclid.aop/1176995476. See correction

[Bha80].

Bhattacharya:1980:CNC

[Bha80] R. N. Bhattacharya. Correction note: Correction to “Criteria for Recurrence and Existence of Invariant Measures for Multidimensional Diffusions”. Annals of Probability, 8(6):1194–1195, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994584. See [Bha78].

Bhattacharya:1985:CLT

[Bha85] Rabi Bhattacharya. A Central Limit Theorem for diffusions with periodic coefficients. Annals of Probability, 13(2):385–396, May 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992998.

Bichteler:1981:SIT

[Bic81] Klaus Bichteler. Stochastic integration and L^{p}-theory of semimartingales. Annals of Probability, 9(1):49–89, February 1981. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994509.

Bingham:1981:TTC

[Bin81] N. H. Bingham. Tauberian theorems and the Central Limit Theorem. Annals of Probability, 9(2):221–231, April 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994464.

Birkel:1988:MBA

[Bir88a] Thomas Birkel. Moment bounds for associated sequences. Annals of Probability, 16(3):1184–1193, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991684.

Birkel:1988:CRC

[Bir88b] Thomas Birkel. On the convergence rate in the Central Limit Theorem for associated processes. Annals of Probability, 16(4): 1685–1698, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991591.

Brockett:1982:UHC

[BK82] Patrick L. Brockett and J. H. B. Kemperman. On the unimodality of high convolutions. Annals of Probability, 10(1):270–277, February 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993933.

Brown:1983:RTP

[BK83] Timothy C. Brown and Joseph Kupka. Ramsey’s theorem and Poisson random measures. Annals of Probability, 11(4):904–908, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993440.

Bruni:1985:ICM

[BK85] C. Bruni and G. Koch. Identifiability of continuous mixtures of unknown Gaussian distributions. Annals of Probability, 13(4):

1341–1357, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992817.

Berman:1989:MGP

[BK89] Simeon M. Berman and Norio Kono. The maximum of a Gaussian process with nonconstant variance: A sharp bound for the distribution tail. Annals of Probability, 17(2):632–650, April 1989. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991419.

Bhattacharya:1988:ACM

[BL88] Rabi N. Bhattacharya and Oesook Lee. Asymptotics of a class of Markov processes which are not in general irreducible. Annals of Probability, 16(3):1333–1347, July 1988. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991694. See cor-

rection [BL97].

Bhattacharya:1997:CBC

[BL97] Rabi N. Bhattacharya and Oesook Lee. Correction: “Asymptotics of a class of Markov processes which are not in general irreducible” [Ann. Probab. 16 (1988), no. 3, 1333–1347; MR 89m:60148]. Annals of Probability, 25(3):1541–1543, July 1997. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1024404526. See

[BL88].

Blackwell:1980:TNB

[Bla80] D. Blackwell. There are no Borel SPLIFs. Annals of Probability, 8(6):1189–1190, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994581.

Blackwell:1981:BSG

[Bla81] D. Blackwell. Borel sets via games. Annals of Probability, 9 (2):321–322, April 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994474.

Bloomfield:1985:SRL

[Blo85] Peter Bloomfield. On series representations for linear predictors. Annals of Probability, 13(1):226–233, February 1985. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993077.

Blumenthal:1983:WCB

[Blu83] R. M. Blumenthal. Weak convergence to Brownian excursion. Annals of Probability, 11(3):798–800, August 1983. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993525.

Burgess:1981:CDO

[BM81] John P. Burgess and R. Daniel Mauldin. Conditional distributions and orthogonal measures. Annals of Probability, 9(5):902–906, October 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994320.

Berger:1982:ESI

[BM82] Marc A. Berger and Victor J. Mizel. An extension of the stochastic integral. Annals of Probability, 10(2):435–450, May 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993868.

Bleher:1987:CPU

[BM87] P. M. Bleher and P. Major. Critical phenomena and universal exponents in statistical physics. On Dyson’s hierarchical model. Annals of Probability, 15(2):431–477, April 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992155;   http://www.
jstor.org/stable/2244058.

Baccelli:1989:DTS

[BM89a] François Baccelli and Armand M. Makowski. Dynamic, transient and stationary behavior of the M/GI/1 queue via martingales. Annals of Probability, 17(4):1691–1699, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991182.

Bretagnolle:1989:HCN

[BM89b] J. Bretagnolle and P. Massart. Hungarian constructions from the nonasymptotic viewpoint. Annals of Probability, 17(1):239–256, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991506.

Bramson:1981:TUS

[BN81] Maury Bramson and Arnold Neidhardt. Thin but unavoidable sets. Annals of Probability, 9(1):154–156, February 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994516.

Bolthausen:1982:ECR

[Bol82a] E. Bolthausen. Exact convergence rates in some martingale Central Limit Theorems. Annals of Probability, 10(3):672–688, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993776.

Bolthausen:1982:CLT

[Bol82b] E. Bolthausen. On the Central Limit Theorem for stationary mixing random fields. Annals of Probability, 10(4):1047–1050, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993726.

Bolthausen:1984:PLD

[Bol84] E. Bolthausen. On the probability of large deviations in Banach spaces. Annals of Probability, 12(2):427–435, May 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993298.

Bolthausen:1989:CLT

[Bol89] Erwin Bolthausen. A Central Limit Theorem for two-dimensional random walks in random sceneries. Annals of Probability, 17(1): 108–115, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991497.

Borovkov:1985:NLT

[Bor85] K. A. Borovkov. A note on a limit theorem for differentiable mappings. Annals of Probability, 13(3):1018–1021, August 1985. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992924.

Bougerol:1987:TPR

[Bou87] Philippe Bougerol. Tightness of products of random matrices and stability of linear stochastic systems. Annals of Probability, 15 (1):40–74, January 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992256.

Bhattacharya:1983:OMC

[BP83a] R. N. Bhattacharya and M. L. Puri. On the order of magnitude of cumulants of von Mises functionals and related statistics. Annals of Probability, 11(2):346–354, May 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993600.

Boland:1983:RS

[BP83b] Philip J. Boland and Frank Proschan. The reliability of K out of N systems. Annals of Probability, 11(3):760–764, August 1983. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993520.

Bass:1984:FLI

[BP84a] Richard F. Bass and Ronald Pyke. Functional law of the iterated logarithm and uniform Central Limit Theorem for partialsum processes indexed by sets. Annals of Probability, 12(1):13–34, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993371.

Bass:1984:SLL

[BP84b] Richard F. Bass and Ronald Pyke. A Strong Law of Large Numbers for partial-sum processes indexed by sets. Annals of Probability, 12 (1):268–271, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993390.

Bass:1985:SWC

[BP85] Richard F. Bass and Ronald Pyke. The space D(A) and weak convergence for set-indexed processes. Annals of Probability, 13 (3):860–884, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992911.

Boland:1988:MGP

[BPT88] Philip J. Boland, Frank Proschan, and Y. L. Tong. Moment and geometric probability inequalities arising from arrangement increasing functions. Annals of Probability, 16(1):407–413, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991911.

Bickel:1982:EES

[BR82] P. J. Bickel and J. Robinson. Edgeworth expansions and smoothness. Annals of Probability, 10(2):500–503, May 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993873.

Bender:1984:CRW

[BR84a] Edward A. Bender and L. Bruce Richmond. Correlated random walks. Annals of Probability, 12(1):274–278, February 1984. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993392.

Brox:1984:EFS

[BR84b] Th. Brox and H. Rost. Equilibrium fluctuations of stochastic particle systems: The role of conserved quantities. Annals of Probability, 12(3):742–759, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993225.

Brown:1988:IMI

[BR88] Lawrence D. Brown and Yosef Rinott. Inequalities for multivariate infinitely divisible processes. Annals of Probability, 16 (2):642–657, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991777.

Baldi:1989:NAD

[BR89] Pierre Baldi and Yosef Rinott. On normal approximations of distributions in terms of dependency graphs. Annals of Probability, 17 (4):1646–1650, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991178.

Bradley:1980:NMC

[Bra80] Richard C. Bradley. A note on a mixing condition. Annals of Probability, 8(3):636–638, June 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994735.

Bradley:1985:CLQ

[Bra85] Richard C. Bradley. On the central limit question under absolute regularity. Annals of Probability, 13(4):1314–1325, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992815.

Bradley:1988:CLT

[Bra88] Richard C. Bradley. A Central Limit Theorem for stationary ρmixing sequences with infinite variance. Annals of Probability, 16 (1):313–332, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991904.

Bramson:1989:SNP

[Bra89] Maury Bramson. Survival of nearest-particle systems with low birth rate. Annals of Probability, 17(2):433–443, April 1989. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991409.

Brossard:1980:CND

[Bro80a] Jean Brossard. Comparaison des “normes” L_{p} du processus croissant et de la variable maximale pour les martingales regulières à deux indices. Theorème local correspondant. (French) [Comparison of L_{p} norms of the growth process and the maximal variable for regular martingales with two indexes. Local correspondence theorem]. Annals of Probability, 8(6):1183–1188, December 1980. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994580.

Brown:1980:BIM

[Bro80b] Mark Brown. Bounds, inequalities, and monotonicity properties for some specialized renewal processes. Annals of Probability, 8 (2):227–240, April 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994773.

Brown:1981:FMP

[Bro81] Mark Brown. Further monotonicity properties for specialized renewal processes. Annals of Probability, 9(5):891–895, October 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994317.

Brown:1983:AID

[Bro83a] Mark Brown. Approximating IMRL distributions by exponential distributions, with applications to first passage times. Annals of Probability, 11(2):419–427, May 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:/

/projecteuclid.org/euclid.aop/1176993607. See corrections

[Bro83b].

Brown:1983:CCB

[Bro83b] Mark Brown. Corrections: Correction to “Approximating IMRL Distributions by Exponential Distributions, with Applications to First Passage Times”. Annals of Probability, 11(4):1055, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993457. See [Bro83a].

Brown:1983:SPA

[Bro83c] Timothy C. Brown. Some Poisson approximations using compensators. Annals of Probability, 11(3):726–744, August 1983. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993517.

Brox:1986:ODD

[Bro86] Th. Brox. A one-dimensional diffusion process in a Wiener medium. Annals of Probability, 14(4):1206–1218, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL http://projecteuclid.org/euclid.aop/
1176992363.

Brown:1987:ISC

[Bro87] Timothy C. Brown. Independent subsets of correlation and other matrices. Annals of Probability, 15(1):416–422, January 1987. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992279.

Bruss:1984:UAC

[Bru84] F. Thomas Bruss. A unified approach to a class of best choice problems with an unknown number of options. Annals of Probability, 12(3):882–889, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993237.

Block:1980:LTC

[BS80a] Henry W. Block and Thomas H. Savits. Laplace transforms for classes of life distributions. Annals of Probability, 8(3):465–474, June 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994721.

Block:1980:MIF

[BS80b] Henry W. Block and Thomas H. Savits. Multivariate increasing failure rate average distributions. Annals of Probability, 8 (4):793–801, August 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994666.

Block:1981:MIP

[BS81] Henry W. Block and Thomas H. Savits. Multidimensional IFRA processes. Annals of Probability, 9(1):162–166, February 1981. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994518.

Byczkowski:1986:ACS

[BS86] T. Byczkowski and K. Samotij. Absolute continuity of stable seminorms. Annals of Probability, 14(1):299–312, January 1986. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992629.

Bally:1987:CMP

[BS87a] Vlad Bally and Lucretiu Stoica. A class of Markov processes which admit local times. Annals of Probability, 15(1):241–262, January 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992266.

Bruss:1987:UAC

[BS87b] F. Thomas Bruss and Stephen M. Samuels. A unified approach to a class of optimal selection problems with an unknown number of options. Annals of Probability, 15(2):824–830, April 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992175.

Block:1982:SCN

[BSS82] Henry W. Block, Thomas H. Savits, and Moshe Shaked. Some concepts of negative dependence. Annals of Probability, 10(3):

765–772, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993784.

Burkholder:1981:GCB

[Bur81] D. L. Burkholder. A geometrical characterization of Banach spaces in which martingale difference sequences are unconditional. Annals of Probability, 9(6):997–1011, December 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994270.

Burkholder:1984:BVP

[Bur84] D. L. Burkholder. Boundary value problems and sharp inequalities for martingale transforms. Annals of Probability, 12(3):647–702, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993220.

Burdzy:1985:BPC

[Bur85] Krzysztof Burdzy. Brownian paths and cones. Annals of Probability, 13(3):1006–1010, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992922.

Burdzy:1987:BEM

[Bur87a] Krzysztof Burdzy. Brownian excursions and minimal thinness. I. Annals of Probability, 15(2):676–689, April 1987. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992165.

Burkholder:1987:SSI

[Bur87b] D. L. Burkholder. A sharp and strict L^{p}-inequality for stochastic integrals. Annals of Probability, 15(1):268–273, January 1987. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992268.

Burdzy:1988:BPC

[Bur88] Krzysztof Burdzy. On Brownian paths connecting boundary points. Annals of Probability, 16(3):1034–1038, July 1988. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991675.

Burdzy:1989:CPB

[Bur89] Krzysztof Burdzy. Cut points on Brownian paths. Annals of Probability, 17(3):1012–1036, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991254.

Burton:1985:SLA

[BW85] Robert Burton and Ed Waymire. Scaling limits for associated random measures. Annals of Probability, 13(4):1267–1278, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992810.

Burton:1986:SCA

[BW86] Robert M. Burton and Ed Waymire. A sufficient condition for association of a renewal process. Annals of Probability, 14(4): 1272–1276, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992368.

Bai:1988:CSL

[BY88a] Z. D. Bai and Y. Q. Yin. Convergence to the semicircle law. Annals of Probability, 16(2):863–875, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991792.

Bai:1988:NSC

[BY88b] Z. D. Bai and Y. Q. Yin. Necessary and sufficient conditions for almost sure convergence of the largest eigenvalue of a Wigner matrix. Annals of Probability, 16(4):1729–1741, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991594;   http://www.jstor.org/stable/2243989.

Byczkowski:1981:APS

[BZ81] T. Byczkowski and T. Zak. Asymptotic properties of semigroups of measures on vector spaces. Annals of Probability, 9(2):211–220, April 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994463.

Cacoullos:1982:ULB

[Cac82] Theophilos Cacoullos. On upper and lower bounds for the variance of a function of a random variable. Annals of Probability, 10 (3):799–809, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993788.

Carlsson:1983:RTE

[Car83] Hasse Carlsson. Remainder term estimates of the renewal function. Annals of Probability, 11(1):143–157, February 1983. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993664.

Carlstein:1986:ANG

[Car86] Edward Carlstein. Asymptotic normality for a general statistic from a stationary sequence. Annals of Probability, 14(4): 1371–1379, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992377.

Carverhill:1988:CLS

[Car88] Andrew Carverhill. Conditioning a lifted stochastic system in a product space. Annals of Probability, 16(4):1840–1853, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991601.

Chiang:1989:LTC

[CC89] Tzuu-Shuh Chiang and Yunshyong Chow. A limit theorem for a class of inhomogeneous Markov processes. Annals of Probability, 17 (4):1483–1502, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991169.

Csaki:1983:HBI

[CCFR83] E. Csaki, M. Csörgő, A. Foldes, and P. Revesz. How big are the increments of the local time of a Wiener process? Annals of Probability, 11(3):593–608, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993504.

Csaki:1989:BLT

[CCFR89] E. Csaki, M. Csörgő, A. Foldes, and P. Revesz. Brownian local time approximated by a Wiener sheet. Annals of Probability, 17(2):516–537, April 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991413.

Chayes:1989:CLH

[CCG^{+}89] J. T. Chayes, L. Chayes, G. R. Grimmett, H. Kesten, and R. H. Schonmann. The correlation length for the high-density phase of Bernoulli percolation. Annals of Probability, 17(4):1277–1302, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991155.

Csorgo:1986:NSC

[CCHM86a] Miklós Csörgő, Sandor Csörgő, Lajos Horváth, and David M. Mason. Normal and stable convergence of integral functions of the empirical distribution function. Annals of Probability, 14(1):86–118, January 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992618.

Csorgo:1986:WEQ

[CCHM86b] Miklós Csörgő, Sandor Csörgő, Lajos Horváth, and David M. Mason. Weighted empirical and quantile processes. Annals of Probability, 14(1):31–85, January 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992617.

Chayes:1987:BPA

[CCN87] J. T. Chayes, L. Chayes, and C. M. Newman. Bernoulli percolation above threshold: An invasion percolation analysis. Annals of Probability, 15(4):1272–1287, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991976.

Chen:1980:OSU

[cCS80] Wen chen Chen and Norman Starr. Optimal stopping in an urn. Annals of Probability, 8(3):451–464, June 1980. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994720.

Cox:1981:SLT

[CD81] J. Theodore Cox and Richard Durrett. Some limit theorems for percolation processes with necessary and sufficient conditions. Annals of Probability, 9(4):583–603, August 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994364.

Cuzick:1982:JCG

[CD82] Jack Cuzick and Johannes P. DuPreez. Joint continuity of Gaussian local times. Annals of Probability, 10(3):810–817, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993789.

Chung:1987:RWA

[CDG87] F. R. K. Chung, Persi Diaconis, and R. L. Graham. Random walks arising in random number generation. Annals of Probability, 15(3):1148–1165, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://www.jstor.org/

stable/2244046;   http://projecteuclid.org/euclid.aop/
1176992088.

Cerrito:1986:RRW

[Cer86] P. B. Cerrito. Recurrence of random walks on completely simple semigroups. Annals of Probability, 14(4):1411–1417, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992382.

Csaki:1986:HSI

[CF86] E. Csaki and A. Foldes. How small are the increments of the local time of a Wiener process? Annals of Probability, 14(2):533–546, April 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992529.

Cox:1983:OTL

[CG83] J. Theodore Cox and David Griffeath. Occupation time limit theorems for the voter model. Annals of Probability, 11(4):876–893, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993438.

Cox:1984:CLT

[CG84] J. Theodore Cox and Geoffrey Grimmett. Central Limit Theorems for associated random variables and the percolation model. Annals of Probability, 12(2):514–528, May 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993303.

Cox:1985:OTC

[CG85] J. Theodore Cox and David Griffeath. Occupation times for critical branching Brownian motions. Annals of Probability, 13(4): 1108–1132, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992799.

Cox:1986:DCT

[CG86] J. Theodore Cox and David Griffeath. Diffusive clustering in the two dimensional voter model. Annals of Probability, 14(2):347–370, April 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992521.

Cover:1989:KCI

[CGG89] Thomas M. Cover, Peter Gacs, and Robert M. Gray. Kolmogorov’s contributions to information theory and algorithmic complexity. Annals of Probability, 17(3):840–865, July 1989. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991250.

Csorgo:1988:DNW

[CH88] Miklós Csörgő and Lajos Horváth. On the distributions of L_{p} norms of weighted uniform empirical and quantile processes. Annals of Probability, 16(1):142–161, January 1988. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991890.

Chen:1980:MTR

[Che80] Louis H. Y. Chen. Martingale transform and random Abel–Dini series. Annals of Probability, 8(3):475–482, June 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994722.

Chernick:1981:LTM

[Che81a] Michael R. Chernick. A limit theorem for the maximum of autoregressive processes with uniform marginal distributions. Annals of Probability, 9(1):145–149, February 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994514.

Chernoff:1981:NII

[Che81b] Herman Chernoff. A note on an inequality involving the normal distribution. Annals of Probability, 9(3):533–535, June 1981. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994428.

Cheng:1985:LDO

[Che85] Shihong Cheng. On limiting distributions of order statistics with variable ranks from stationary sequences. Annals of Probability, 13(4):1326–1340, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992816.

Chen:1988:CLT

[Che88] Louis H. Y. Chen. The Central Limit Theorem and Poincarétype inequalities. Annals of Probability, 16(1):300–304, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991902.

Chiang:1982:LBA

[Chi82] Tzuu-Shuh Chiang. A lower bound of the asymptotic behavior of some Markov processes. Annals of Probability, 10(4):955–967, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993717.

Chiang:1985:LBL

[Chi85] Tzuu-Shuh Chiang. On the lower bound of large deviation of random walks. Annals of Probability, 13(1):90–96, February 1985. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993068.

Csorgo:1988:ADT

[CHM88] Sandor Csörgő, Erich Haeusler, and David M. Mason. The asymptotic distribution of trimmed sums. Annals of Probability, 16 (2):672–699, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991780.

Cranston:1989:SCH

[CHM89] M. Cranston, P. Hsu, and P. March. Smoothness of the convex Hull of planar Brownian motion. Annals of Probability, 17 (1):144–150, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991500.

Csorgo:1987:IPR

[CHS87] Miklós Csörgő, Lajos Horváth, and Josef Steinebach. Invariance principles for renewal processes. Annals of Probability, 15(4): 1441–1460, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991986.

Chosid:1978:RRM

[CI78] Leo Chosid and Richard Isaac. On the range of recurrent Markov chains. Annals of Probability, 6(4):680–687, August 1978. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995489. See

correction [CI80].

Chosid:1980:CNC

[CI80] Leo Chosid and Richard Isaac. Correction notes: Correction to “On the Range of Recurrent Markov Chains”. Annals of Probability, 8(5):1000, October 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994629. See [CI78].

Carmona:1983:EMH

[CK83a] Rene Carmona and Abel Klein. Exponential moments for hitting times of uniformly ergodic Markov processes. Annals of Probability, 11(3):648–655, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993509.

Cox:1983:SBA

[CK83b] David C. Cox and J. H. B. Kemperman. Sharp bounds on the absolute moments of a sum of two i.i.d. random variables. Annals of Probability, 11(3):765–771, August 1983. CODEN APBYAE.

ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993521.

Chen:1989:CMM

[CL89a] Mu-Fa Chen and Shao-Fu Li. Coupling methods for multidimensional diffusion processes. Annals of Probability, 17(1):151–177, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991501.

Cornea:1989:CPC

[CL89b] Aurel Cornea and Peter A. Loeb. A convergence property for conditional expectation. Annals of Probability, 17(1):353–356, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991513.

Clark:1987:APS

[Cla87] Charles R. Clark. Asymptotic properties of some multidimensional diffusions. Annals of Probability, 15(3):985–1008, July 1987. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992076.

Cambanis:1980:SPP

[CM80] Stamatis Cambanis and Grady Miller. Some path properties of p th order and symmetric stable processes. Annals of Probability, 8(6): 1148–1156, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994575.

Csorgo:1986:ADS

[CM86] Sandor Csörgő and David M. Mason. The asymptotic distribution of sums of extreme values from a regularly varying distribution. Annals of Probability, 14(3):974–983, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992451.

Cuesta:1989:NWM

[CM89] Juan Antonio Cuesta and Carlos Matran. Notes on the Wasserstein metric in Hilbert spaces. Annals of Probability, 17(3):1264–1276, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991269.

Chaleyat-Maurel:1980:RDS

[CMEM80] M. Chaleyat-Maurel, N. El Karoui, and B. Marchal. Réflexion discontinue et systèmes stochastiques. (French) [Discontinuous reflection and stochastic systems]. Annals of Probability, 8(6): 1049–1067, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994567.

Cohen:1984:SLR

[CN84] Joel E. Cohen and Charles M. Newman. The stability of large random matrices and their products. Annals of Probability, 12 (2):283–310, May 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993291.

Carmona:1988:RNW

[CN88] Rene Carmona and David Nualart. Random nonlinear wave equations: Propagation of singularities. Annals of Probability, 16 (2):730–751, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991784.

Cogburn:1980:MCR

[Cog80] Robert Cogburn. Markov chains in random environments: The case of Markovian environments. Annals of Probability, 8(5): 908–916, October 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994620.

Cogburn:1985:DFP

[Cog85] Robert Cogburn. On the distribution of first passage and return times for small sets. Annals of Probability, 13(4):1219–1223, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992806.

Cohen:1981:GRS

[Coh81] Joel E. Cohen. A geometric representation of a stochastic matrix: Theorem and conjecture. Annals of Probability, 9(5):899–901, October 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994319.

Cohn:1985:MAS

[Coh85] Harry Cohn. A martingale approach to supercritical (CMJ) branching processes. Annals of Probability, 13(4):1179–1191, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992803.

Cohen:1986:LIM

[Coh86] Joel E. Cohen. A life of the immeasurable mind. Annals of Probability, 14(4):1139–1148, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992358.

Cohn:1989:GMS

[Coh89] Harry Cohn. On the growth of the multitype supercritical branching process in a random environment. Annals of Probability, 17 (3):1118–1123, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991259.

Cooil:1985:LMD

[Coo85] Bruce Cooil. Limiting multivariate distributions of intermediate order statistics. Annals of Probability, 13(2):469–477, May 1985. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993003.

Cox:1982:ENR

[Cox82] Dennis D. Cox. On the existence of natural rate of escape functions for infinite dimensional Brownian motions. Annals of Probability, 10(3):623–638, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993772.

Cox:1984:APC

[Cox84] J. Theodore Cox. An alternate proof of a correlation inequality of Harris. Annals of Probability, 12(1):272–273, February 1984. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993391.

Cox:1988:SLT

[Cox88] J. T. Cox. Some limit theorems for voter model occupation times. Annals of Probability, 16(4):1559–1569, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991583.

Cox:1989:CRW

[Cox89] J. T. Cox. Coalescing random walks and voter model consensus times on the torus in Z^{d}. Annals of Probability, 17(4):1333–1366, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991158.

Chen:1984:NBS

[CR84] Jeesen Chen and Herman Rubin. A note on the behavior of sample statistics when the population mean is infinite. Annals of Probability, 12(1):256–261, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993388.

Campanino:1985:UBC

[CR85] M. Campanino and L. Russo. An upper bound on the critical percolation probability for the three- dimensional cubic lattice. Annals of Probability, 13(2):478–491, May 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993004.

Cranston:1987:MAB

[Cra87] M. Cranston. On the means of approach to the boundary of Brownian motion. Annals of Probability, 15(3):1009–1013, July 1987. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992077.

Cressie:1980:ADS

[Cre80] Noel Cressie. The asymptotic distribution of the scan statistic under uniformity. Annals of Probability, 8(4):828–840, August 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994669.

Cambanis:1985:CQF

[CRW85] Stamatis Cambanis, Jan Rosinski, and Wojbor A. Woyczynski. Convergence of quadratic forms in p-stable random variables and θ_{p}-Radonifying operators. Annals of Probability, 13(3):885–897, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992912.

Csorgo:1981:IER

[CS81] M. Csörgő and J. Steinebach. Improved Erdős–Rényi and strong approximation laws for increments of partial sums. Annals of Probability, 9(6):988–996, December 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994269.

Chaganty:1985:LDL

[CS85a] Narasinga R. Chaganty and J. Sethuraman. Large deviation local limit theorems for arbitrary sequences of random variables. Annals of Probability, 13(1):97–114, February 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993069.

Clifford:1985:SPP

[CS85b] Peter Clifford and Aidan Sudbury. A sample path proof of the duality for stochastically monotone Markov processes. Annals of Probability, 13(2):558–565, May 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993008.

Chaganty:1987:LTA

[CS87] Narasinga Rao Chaganty and Jayaram Sethuraman. Limit theorems in the area of large deviations for some dependent random variables. Annals of Probability, 15(2):628–645, April 1987. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992162.

Csiszar:1984:SPG

[Csi84] Imre Csiszar. Sanov property, generalized I-projection and a conditional limit theorem. Annals of Probability, 12(3):768–793, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993227.

Csorgo:1981:LBE

[Csö81] Sandor Csörgő. Limit behaviour of the empirical characteristic function. Annals of Probability, 9(1):130–144, February 1981. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994513.

Cuzick:1978:LNZ

[Cuz78] Jack Cuzick. Local nondeterminism and the zeros of Gaussian processes. Annals of Probability, 6(1):72–84, February 1978. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995611. See

correction [Cuz87].

Cuzick:1982:CGL

[Cuz82] Jack Cuzick. Continuity of Gaussian local times. Annals of Probability, 10(3):818–823, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993790.

Cuzick:1987:CLN

[Cuz87] Jack Cuzick. Correction: Local nondeterminism and the zeros of Gaussian processes. Annals of Probability, 15(3):1229, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992096. See [Cuz78].

Cabana:1982:TPB

[CW82] Enrique M. Cabana and Mario Wschebor. The two-parameter Brownian bridge: Kolmogorov inequalities and upper and lower bounds for the distribution of the maximum. Annals of Probability, 10(2):289–302, May 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993858.

Chow:1986:NFS

[CZ86] Yuan Shih Chow and Cun-Hui Zhang. A note on Feller’s Strong Law of Large Numbers. Annals of Probability, 14(3):1088–1094, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992464.

deAcosta:1980:EMV

[dA80] Alejandro de Acosta. Exponential moments of vector valued random series and triangular arrays. Annals of Probability, 8(2):381– 389, April 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994786.

deAcosta:1982:IPP

[dA82] Alejandro de Acosta. Invariance principles in probability for triangular arrays of B-valued random vectors and some applications. Annals of Probability, 10(2):346–373, May 1982. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993862.

deAcosta:1983:NPH

[dA83a] Alejandro de Acosta. A new proof of the Hartman–Wintner law of the iterated logarithm. Annals of Probability, 11(2):270–276, May 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993596.

deAcosta:1983:SDF

[dA83b] Alejandro de Acosta. Small deviations in the functional Central Limit Theorem with applications to functional laws of the iterated logarithm. Annals of Probability, 11(1):78–101, February 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993661.

deAcosta:1988:LDV

[dA88] A. de Acosta. Large deviations for vector-valued functionals of a Markov chain: Lower bounds. Annals of Probability, 16(3):925– 960, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991670.

deAcosta:1983:SRC

[dAK83] A. de Acosta and J. Kuelbs. Some results on the cluster set

C(\{\frac{S_{n}}{a_{n}} \}) and the LIL. Annals of Probability, 11(1):102–122,

February 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993662.

Daley:1980:TBR

[Dal80] D. J. Daley. Tight bounds for the renewal function of a random walk. Annals of Probability, 8(3):615–621, June 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994732.

Daley:1988:TBE

[Dal88] D. J. Daley. Tight bounds on the exponential approximation of some aging distributions. Annals of Probability, 16(1):414–423, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991912.

Darling:1983:SCG

[Dar83] D. A. Darling. On the supremum of a certain Gaussian process. Annals of Probability, 11(3):803–806, August 1983. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993527.

Davis:1983:SLP

[Dav83] Richard A. Davis. Stable limits for partial sums of dependent random variables. Annals of Probability, 11(2):262–269, May 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993595.

Deheuvels:1987:LLE

[DD87] Paul Deheuvels and Luc Devroye. Limit laws of Erdős–Rényi– Shepp type. Annals of Probability, 15(4):1363–1386, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991982.

Deheuvels:1986:ECR

[DDL86] Paul Deheuvels, Luc Devroye, and James Lynch. Exact convergence rate in the limit theorems of Erdős–Rényi and Shepp. Annals of Probability, 14(1):209–223, January 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992623.

Dehling:1986:CLT

[DDP86] Herold Dehling, Manfred Denker, and Walter Philipp. Central Limit Theorems for mixing sequences of random variables under minimal conditions. Annals of Probability, 14(4):1359–1370, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992376.

DeMare:1980:OPC

[De 80] Jacques De Mare. Optimal prediction of catastrophes with applications to Gaussian processes. Annals of Probability, 8(4):841–850, August 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994670.

DeAcosta:1981:IVR

[De 81] Alejandro De Acosta. Inequalities for B-valued random vectors with applications to the Strong Law of Large Numbers. Annals of Probability, 9(1):157–161, February 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994517.

DeHaan:1984:SRM

[De 84] L. De Haan. A spectral representation for max-stable processes. Annals of Probability, 12(4):1194–1204, November 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993148.

DeBlassie:1987:STB

[DeB87] R. Dante DeBlassie. Stopping times of Bessel processes. Annals of Probability, 15(3):1044–1051, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992079.

DeBlassie:1988:DCD

[DeB88] R. Dante DeBlassie. Doob’s conditioned diffusions and their lifetimes. Annals of Probability, 16(3):1063–1083, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991678.

Deheuvels:1982:SLB

[Deh82] Paul Deheuvels. Strong limiting bounds for maximal uniform spacings. Annals of Probability, 10(4):1058–1065, November 1982. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993728.

Deheuvels:1984:SLT

[Deh84] Paul Deheuvels. Strong limit theorems for maximal spacings from a general univariate distribution. Annals of Probability, 12(4): 1181–1193, November 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993147.

Deheuvels:1986:IED

[Deh86] Paul Deheuvels. On the influence of the extremes of an i.i.d. sequence on the maximal spacings. Annals of Probability, 14(1): 194–208, January 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992622.

Derriennic:1983:TEP

[Der83] Yves Derriennic. Un theorème ergodique presque sous-additif. (French) [An almost-subadditive ergodic theorem]. Annals of Probability, 11(3):669–677, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993511.

Deuschel:1988:CLT

[Deu88] Jean-Dominique Deuschel. Central Limit Theorem for an infinite lattice system of interacting diffusion processes. Annals of Probability, 16(2):700–716, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991781.

Deuschel:1989:IPE

[Deu89] Jean-Dominique Deuschel. Invariance principle and empirical mean large deviations of the critical Ornstein–Uhlenbeck process. Annals of Probability, 17(1):74–90, January 1989. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991495.

Devroye:1981:LIL

[Dev81] Luc Devroye. Laws of the iterated logarithm for order statistics of uniform spacings. Annals of Probability, 9(5):860–867, October 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994313.

Devroye:1982:LLL

[Dev82] Luc Devroye. A log log law for maximal uniform spacings. Annals of Probability, 10(3):863–868, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993799.

Diaconis:1980:FTM

[DF80a] P. Diaconis and D. Freedman. De Finetti’s theorem for Markov chains. Annals of Probability, 8(1):115–130, February 1980. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994828.

Diaconis:1980:FES

[DF80b] P. Diaconis and D. Freedman. Finite exchangeable sequences. Annals of Probability, 8(4):745–764, August 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994663.

Dawson:1989:SHL

[DFG89] Donald A. Dawson, Klaus Fleischmann, and Luis G. Gorostiza. Stable hydrodynamic limit fluctuations of a critical branching particle system in a random medium. Annals of Probability, 17 (3):1083–1117, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991258.

Durrett:1983:SCP

[DG83] Richard Durrett and David Griffeath. Supercritical contact processes on Z. Annals of Probability, 11(1):1–15, February 1983. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993655.

Demko:1981:DPT

[DH81] Stephen Demko and Theodore P. Hill. Decision processes with total-cost criteria. Annals of Probability, 9(2):293–301, April 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994470.

Dawson:1982:WRM

[DH82] Donald A. Dawson and Kenneth J. Hochberg. Wandering random measures in the Fleming–Viot model. Annals of Probability, 10 (3):554–580, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993767.

Daley:1984:LLM

[DH84] D. J. Daley and Peter Hall. Limit laws for the maximum of weighted and shifted i.i.d. random variables. Annals of Probability, 12(2):571–587, May 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993306.

deHaan:1982:LLT

[dHR82] L. de Haan and S. I. Resnick. Local limit theorems for sample extremes. Annals of Probability, 10(2):396–413, May 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993865.

deHaan:1989:ERC

[dHR89] L. de Haan and S. T. Rachev. Estimates of the rate of convergence for max-stable processes. Annals of Probability, 17(2):651–677, April 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991420.

Durrett:1981:SLS

[DL81] Richard Durrett and Thomas M. Liggett. The shape of the limit set in Richardson’s growth model. Annals of Probability, 9(2):186–193, April 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994460.

Durrett:1988:CPFa

[DL88] Richard Durrett and Xiu-Fang Liu. The contact process on a finite set. Annals of Probability, 16(3):1158–1173, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991682.

delaCal:1989:TST

[dlC89] Jesus de la Cal. On the three series theorem in number theory. Annals of Probability, 17(1):357–361, January 1989. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991514.

Davis:1984:RSS

[DM84] Burgess Davis and Itrel Monroe. Randomly started signals with white noise. Annals of Probability, 12(3):922–925, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993243.

Does:1982:BET

[Doe82] Ronald J. M. M. Does. Berry–Esseen theorems for simple linear rank statistics under the null- hypothesis. Annals of Probability, 10 (4):982–991, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993719.

Doney:1987:WHF

[Don87a] R. A. Doney. On Wiener–Hopf factorisation and the distribution of extrema for certain stable processes. Annals of Probability, 15(4): 1352–1362, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991981.

Donnelly:1987:RAF

[Don87b] Peter Donnelly. Review: Alan F. Karr, Point Processes and their Statistical Inference. Annals of Probability, 15(3):1226–1227, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992094.

Doob:1989:KEW

[Doo89] J. L. Doob. Kolmogorov’s early work on convergence theory and foundation. Annals of Probability, 17(3):815–821, July 1989. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991247.

Dehling:1982:ASI

[DP82] Herold Dehling and Walter Philipp. Almost sure invariance principles for weakly dependent vector-valued random variables. Annals of Probability, 10(3):689–701, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993777.

Davis:1985:BSP

[DP85] Burgess Davis and Edwin Perkins. Brownian slow points: The critical case. Annals of Probability, 13(3):779–803, August 1985. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992908.

Deheuvels:1986:SAP

[DP86] P. Deheuvels and D. Pfeifer. A semigroup approach to Poisson approximation. Annals of Probability, 14(2):663–676, April 1986. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992536.

DeMasi:1986:AEF

[DPSW86] A. De Masi, E. Presutti, H. Spohn, and W. D. Wick. Asymptotic equivalence of fluctuation fields for reversible exclusion processes with speed change. Annals of Probability, 14(2):409–423, April 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992524.

DeHaan:1984:ABF

[DR84] L. De Haan and S. I. Resnick. Asymptotically balanced functions and stochastic compactness of sample extremes. Annals of Probability, 12(2):588–608, May 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993307.

Davis:1985:LTM

[DR85] Richard Davis and Sidney Resnick. Limit theory for moving averages of random variables with regularly varying tail probabilities. Annals of Probability, 13(1):179–195, February 1985. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993074.

Drobot:1982:PVC

[Dro82] Vladimir Drobot. Probabilistic version of a curvature formula. Annals of Probability, 10(3):860–862, August 1982. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993798.

Davis:1988:CBP

[DS88a] Burgess Davis and Thomas S. Salisbury. Connecting Brownian paths. Annals of Probability, 16(4):1428–1457, October 1988. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991577.

Durrett:1988:CPFb

[DS88b] Richard Durrett and Roberto H. Schonmann. The contact process on a finite set. II. Annals of Probability, 16(4):1570–1583, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991584.

Deheuvels:1989:SRI

[DS89] Paul Deheuvels and Josef Steinebach. Sharp rates for increments of renewal processes. Annals of Probability, 17(2):700–722, April 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991422.

Durrett:1989:CPF

[DST89] Richard Durrett, Roberto H. Schonmann, and Nelson I. Tanaka. The contact process on a finite set. III: The critical case. Annals of Probability, 17(4):1303–1321, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991156.

Dudley:1987:UDC

[Dud87] R. M. Dudley. Universal Donsker classes and metric entropy. Annals of Probability, 15(4):1306–1326, October 1987. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991978.

Dunham:1980:AAS

[Dun80] James G. Dunham. Abstract alphabet sliding-block entropy compression coding with a fidelity criterion. Annals of Probability, 8(6): 1085–1092, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994570.

Dupuis:1988:LDA

[Dup88] Paul Dupuis. Large deviations analysis of some recursive algorithms with state dependent noise. Annals of Probability, 16(4): 1509–1536, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991581.

Durrett:1980:GOD

[Dur80] Richard Durrett. On the growth of one dimensional contact processes. Annals of Probability, 8(5):890–907, October 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994619.

Durrett:1982:NPS

[Dur82] Richard Durrett. A new proof of Spitzer’s result on the winding of two dimensional Brownian motion. Annals of Probability, 10 (1):244–246, February 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993928.

Durrett:1984:OPT

[Dur84] Richard Durrett. Oriented percolation in two dimensions. Annals of Probability, 12(4):999–1040, November 1984. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993140.

Dutko:1989:CLT

[Dut89] Michael Dutko. Central Limit Theorems for infinite urn models. Annals of Probability, 17(3):1255–1263, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991268.

Dykstra:1985:IPO

[Dyk85] Richard L. Dykstra. An iterative procedure for obtaining Iprojections onto the intersection of convex sets. Annals of Probability, 13(3):975–984, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992918.

Dynkin:1988:RSI

[Dyn88a] E. B. Dynkin. Regularized self-intersection local times of planar Brownian motion. Annals of Probability, 16(1):58–74, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991885.

Dynkin:1988:SIG

[Dyn88b] E. B. Dynkin. Self-intersection gauge for random walks and for Brownian motion. Annals of Probability, 16(1):1–57, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991884.

Dynkin:1989:KTM

[Dyn89] E. B. Dynkin. Kolmogorov and the theory of Markov processes. Annals of Probability, 17(3):822–832, July 1989. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991248.

Eagleson:1981:EDT

[Eag81] G. K. Eagleson. An extended dichotomy theorem for sequences of pairs of Gaussian measures. Annals of Probability, 9(3):453–459, June 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994417.

Eberlein:1986:SIP

[Ebe86] Ernst Eberlein. On strong invariance principles under dependence assumptions. Annals of Probability, 14(1):260–270, January 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992626.

Edgar:1982:AA

[Edg82] G. A. Edgar. Additive amarts. Annals of Probability, 10(1): 199–206, February 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993923.

Embrechts:1981:CTI

[EG81] Paul Embrechts and Charles M. Goldie. Comparing the tail of an infinitely divisible distribution with integrals of its Lévy measure. Annals of Probability, 9(3):468–481, June 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994419.

Ethier:1987:IMS

[EG87] S. N. Ethier and R. C. Griffiths. The infinitely-many-sites model as a measure-valued diffusion. Annals of Probability, 15(2):515–545, April 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992157.

Einmahl:1987:SIP

[Ein87] Uwe Einmahl. Strong invariance principles for partial sums of independent random vectors. Annals of Probability, 15(4):1419–1440, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991985.

Einmahl:1989:SRS

[Ein89] Uwe Einmahl. Stability results and strong invariance principles for partial sums of Banach space valued random variables. Annals of Probability, 17(1):333–352, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991512.

Elliott:1989:IPH

[EK89] Robert J. Elliott and Michael Kohlmann. Integration by parts, homogeneous chaos expansions and smooth densities. Annals of Probability, 17(1):194–207, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991504.

Ellis:1980:CAM

[Ell80a] Martin H. Ellis. Conditions for attaining d̄ by a Markovian joining. Annals of Probability, 8(3):431–440, June 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994718.

Ellis:1980:KCC

[Ell80b] Martin H. Ellis. On Kamae’s conjecture concerning the \overline{d}-distance between two- state Markov processes. Annals of Probability, 8 (2):372–376, April 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994784.

Ellis:1984:LDG

[Ell84] Richard S. Ellis. Large deviations for a general class of random vectors. Annals of Probability, 12(1):1–12, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993370.

Elliott:1986:RTD

[Ell86] Robert J. Elliott. Reverse time differentiation and smoothing formulae for a finite state Markov process. Annals of Probability, 14(2):480–489, April 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992527.

Ellis:1988:IMC

[Ell88a] Richard S. Ellis. Inequalities for multivariate compound Poisson distributions. Annals of Probability, 16(2):658–661, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991778.

Ellis:1988:LDE

[Ell88b] Richard S. Ellis. Large deviations for the empirical measure of a Markov chain with an application to the multivariate empirical measure. Annals of Probability, 16(4):1496–1508, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991580.

Elton:1981:LLN

[Elt81] John Elton. A Law of Large Numbers for identically distributed martingale differences. Annals of Probability, 9(3):405– 412, June 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994414.

Einmahl:1988:LIL

[EM88a] John H. J. Einmahl and David M. Mason. Laws of the iterated logarithm in the tails for weighted uniform empirical processes. Annals of Probability, 16(1):126–141, January 1988. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991889.

Einmahl:1988:SLT

[EM88b] John H. J. Einmahl and David M. Mason. Strong limit theorems for weighted quantile processes. Annals of Probability, 16(4): 1623–1643, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991588.

Emery:1982:GSI

[Eme82] M. Emery. A generalization of stochastic integration with respect to semimartingales. Annals of Probability, 10(3):709–727, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993779.

Embrechts:1984:RTB

[EMO84] Paul Embrechts, Makoto Maejima, and Edward Omey. A renewal theorem of Blackwell type. Annals of Probability, 12(2):561–570, May 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993305.

Englund:1981:RTE

[Eng81] Gunnar Englund. A remainder term estimate for the normal approximation in classical occupancy. Annals of Probability, 9 (4):684–692, August 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994376.

El-Neweihi:1987:CIS

[ENS87] Emad El-Neweihi and Thomas H. Savits. Convolution of the IFRA scaled-mins class. Annals of Probability, 15(1):423–427, January 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992280.

Epstein:1989:SLT

[Eps89] Raisa Epstein. Some limit theorems for functionals of the Brownian sheet. Annals of Probability, 17(2):538–558, April 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991414.

Ellis:1982:LMG

[ER82] Richard S. Ellis and Jay S. Rosen. Laplace’s method for Gaussian integrals with an application to statistical mechanics. Annals of Probability, 10(1):47–66, February 1982. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993913. See cor-

rection [ER83].

Ellis:1983:CCB

[ER83] Richard S. Ellis and Jay S. Rosen. Correction: Correction to “Laplace’s Method for Gaussian Integrals with an Application to Statistical Mechanics”. Annals of Probability, 11(2):456, May 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993611. See [ER82].

Erickson:1980:REI

[Eri80] K. Bruce Erickson. Rates of escape of infinite dimensional Brownian motion. Annals of Probability, 8(2):325–338, April 1980. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994780.

Erickson:1981:LSC

[Eri81] Roy V. Erickson. Lipschitz smoothness and convergence with applications to the Central Limit Theorem for summation processes. Annals of Probability, 9(5):831–851, October 1981. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994311.

Escriba:1987:SBM

[Esc87] L. Barba Escriba. A stopped Brownian motion formula with two sloping line boundaries. Annals of Probability, 15(4):1524–1526, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991991.

Emery:1983:VPP

[ESY83] M. Emery, C. Stricker, and J. A. Yan. Valeurs prises par les martingales locales continues à un instant donné. (French) [Values taken by local continuous martingales at a given instant]. Annals of Probability, 11(3):635–641, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993507.

Etemadi:1980:FLT

[Ete80] Nasrollah Etemadi. Further limit theorems for the range of a two-parameter random walk in space. Annals of Probability, 8 (5):917–927, October 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994621.

Einmahl:1988:SBW

[EvZ88] John H. J. Einmahl and Martien C. A. van Zuijlen. Strong bounds for weighted empirical distribution functions based on uniform spacings. Annals of Probability, 16(1):108–125, January 1988. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991888.

Ellis:1989:ULD

[EW89] Richard S. Ellis and Aaron D. Wyner. Uniform large deviation property of the empirical process of a Markov chain. Annals of Probability, 17(3):1147–1151, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991261.

Faden:1985:ERC

[Fad85] Arnold M. Faden. The existence of regular conditional probabilities: Necessary and sufficient conditions. Annals of Probability, 13 (1):288–298, February 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993081.

Falk:1984:CSD

[Fal84] M. Falk. On the convergence of spectral densities of arrays of weakly stationary processes. Annals of Probability, 12(3):918–921, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993242.

Falkner:1987:CBM

[Fal87] Neil Falkner. Conditional Brownian motion in rapidly exhaustible domains. Annals of Probability, 15(4):1501–1514, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991989.

Farahmand:1986:ANR

[Far86] Kambiz Farahmand. On the average number of real roots of a random algebraic equation. Annals of Probability, 14(2):702–709, April 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.

aop/1176992539. See correction [Far87].

Farahmand:1987:CAN

[Far87] Kambiz Farahmand. Correction: On the average number of real roots of a random algebraic equation. Annals of Probability, 15(3): 1230, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992097. See [Far86].

Fishburn:1988:MSR

[FDS88] Peter C. Fishburn, Peter G. Doyle, and L. A. Shepp. The match set of a random permutation has the FKG property. Annals of Probability, 16(3):1194–1214, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991685.

Fernholz:1980:LPA

[Fer80] Luisa Turrin Fernholz. Limit points associated with generalized iterated logarithm laws. Annals of Probability, 8(2):390–392, April 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994787.

Ferrari:1986:SEP

[Fer86] Pablo A. Ferrari. The simple exclusion process as seen from a tagged particle. Annals of Probability, 14(4):1277–1290, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992369.

Frangos:1987:QVC

[FI87] Nikos E. Frangos and Peter Imkeller. Quadratic variation for a class of L log^{+} L-bounded two-parameter martingales. Annals of Probability, 15(3):1097–1111, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992083.

Fill:1983:CRR

[Fil83] James Allen Fill. Convergence rates related to the Strong Law of Large Numbers. Annals of Probability, 11(1):123–142, February 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993663.

Fill:1988:BCR

[Fil88] James Allen Fill. Bounds on the coarseness of random sums. Annals of Probability, 16(4):1644–1664, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991589.

Finster:1982:MTF

[Fin82a] Mark Finster. The maximum term and first passage times for autoregressions. Annals of Probability, 10(3):737–744, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993781.

Finster:1982:OSA

[Fin82b] Mark Finster. Optimal stopping on autoregressive schemes. Annals of Probability, 10(3):745–753, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993782.

Fitzsimmons:1987:CTP

[Fit87] P. J. Fitzsimmons. A converse to a theorem of P. Lévy. Annals of Probability, 15(4):1515–1523, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991990.

Falk:1986:RWS

[FK86] M. Falk and W. Kohne. On the rate at which the sample extremes become independent. Annals of Probability, 14(4):1339–1346, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992374.

Flatto:1982:LTS

[Fla82] L. Flatto. Limit theorems for some random variables associated with urn models. Annals of Probability, 10(4):927–934, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993714.

Follmer:1988:LDE

[FO88] Hans Follmer and Steven Orey. Large deviations for the empirical field of a Gibbs measure. Annals of Probability, 16(3):961–977, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991671.

Follmer:1984:ASC

[Fol84] Hans Follmer. Almost sure convergence of multiparameter martingales for Markov random fields. Annals of Probability, 12(1): 133–140, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993378.

Flatto:1985:RSG

[FOW85] L. Flatto, A. M. Odlyzko, and D. B. Wales. Random shuffles and group representations. Annals of Probability, 13(1):154–178, February 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993073.

Feigin:1981:GPD

[FP81] Paul D. Feigin and Ury Passy. The geometric programming dual to the extinction probability problem in simple branching processes. Annals of Probability, 9(3):498–503, June 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994422.

Falk:1988:IOS

[FR88] M. Falk and R.-D. Reiss. Independence of order statistics. Annals of Probability, 16(2):854–862, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991791.

Falk:1989:WCS

[FR89] M. Falk and R.-D. Reiss. Weak convergence of smoothed and nonsmoothed bootstrap quantile estimates. Annals of Probability, 17 (1):362–371, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991515.

Frangos:1985:RBV

[Fra85] Nikos E. Frangos. On regularity of Banach-valued processes. Annals of Probability, 13(3):985–990, August 1985. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992919.

Freed:1981:PSA

[Fre81] Karl F. Freed. Polymers as self-avoiding walks. Annals of Probability, 9(4):537–554, August 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994359.

Freidlin:1985:LTL

[Fre85] Mark Freidlin. Limit theorems for large deviations and reactiondiffusion equations. Annals of Probability, 13(3):639–675, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992901.

Fritz:1987:GDI

[Fri87] J. Fritz. Gradient dynamics of infinite point systems. Annals of Probability, 15(2):478–514, April 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992156.

Fernholz:1980:SRV

[FT80a] Luisa Turrin Fernholz and Henry Teicher. Stability of random variables and iterated logarithm laws for martingales and quadratic forms. Annals of Probability, 8(4):765–774, August 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994664.

Fremlin:1980:GM

[FT80b] D. H. Fremlin and M. Talagrand. A Gaussian measure on l^{∞}. Annals of Probability, 8(6):1192–1193, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994583.

Fox:1985:NLT

[FT85] Robert Fox and Murad S. Taqqu. Noncentral limit theorems for quadratic forms in random variables having long-range dependence. Annals of Probability, 13(2):428–446, May 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993001.

Fitzsimmons:1988:SRS

[FT88] P. J. Fitzsimmons and Michael Taksar. Stationary regenerative sets and subordinators. Annals of Probability, 16(3):1299–1305, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991692.

Gaenssler:1988:RGR

[Gae88] Peter Gaenssler. Review: Galen R. Shorack, Jon A. Wellner, Empirical Processes with Applications to Statistics. Annals of Probability, 16(3):1372–1388, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991697.

Gandolfi:1989:UIC

[Gan89] Alberto Gandolfi. Uniqueness of the infinite cluster for stationary Gibbs states. Annals of Probability, 17(4):1403–1415, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991161.

Gardner:1982:NCD

[Gar82] R. J. Gardner. A note on conditional distributions and orthogonal measures. Annals of Probability, 10(3):877–878, August 1982. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993802.

Gawronski:1984:BSS

[Gaw84] Wolfgang Gawronski. On the Bell-shape of stable densities. Annals of Probability, 12(1):230–242, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993386.

Gawronski:1988:AFD

[Gaw88] Wolfgang Gawronski. Asymptotic forms for the derivatives of one-sided stable laws. Annals of Probability, 16(3):1348–1364, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991695.

Geman:1980:LTN

[Gem80] Stuart Geman. A limit theorem for the norm of random matrices. Annals of Probability, 8(2):252–261, April 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994775.

Geman:1986:SRL

[Gem86] Stuart Geman. The spectral radius of large random matrices. Annals of Probability, 14(4):1318–1328, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992372.

Getoor:1988:KMP

[Get88] R. K. Getoor. Killing a Markov process under a stationary measure involves creation. Annals of Probability, 16(2):564–585, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991773.

Grize:1987:CLP

[GF87] Yves L. Grize and Terrence L. Fine. Continuous lower probabilitybased models for stationary processes with bounded and divergent time averages. Annals of Probability, 15(2):783–803, April 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992172.

Guijing:1986:ASQ

[GFZ86] Chen Guijing, Kong Fanchao, and Lin Zhengyan. Answers to some questions about increments of a Wiener process. Annals of Probability, 14(4):1252–1261, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992366.

Gray:1982:SCA

[GG82] Lawrence Gray and David Griffeath. A stability criterion for attractive nearest neighbor spin systems on Z. Annals of Probability, 10(1):67–85, February 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993914.

Goldie:1986:CSI

[GG86a] Charles M. Goldie and Priscilla E. Greenwood. Characterisations of set-indexed Brownian motion and associated conditions for finite-dimensional convergence. Annals of Probability, 14(3):802– 816, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992439.

Goldie:1986:VSI

[GG86b] Charles M. Goldie and Priscilla E. Greenwood. Variance of setindexed sums of mixing random variables and weak convergence of set-indexed processes. Annals of Probability, 14(3):817–839, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992440.

Golden:1988:NEA

[GGL88] K. Golden, S. Goldstein, and J. L. Lebowitz. Nash estimates and the asymptotic behavior of diffusions. Annals of Probability, 16 (3):1127–1146, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991680.

Geman:1980:OD

[GH80] Donald Geman and Joseph Horowitz. Occupation densities. Annals of Probability, 8(1):1–67, February 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994824.

Goldstein:1981:NSL

[GH81] Martin I. Goldstein and Fred M. Hoppe. Necessary and sufficient lifetime conditions for normed convergence of critical agedependent processes with infinite variance. Annals of Probability, 9(3):490–497, June 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994421.

Gine:1985:CDA

[GH85] Evarist Gine and Marjorie G. Hahn. Characterization and domains of attraction of p-stable random compact sets. Annals of Probability, 13(2):447–468, May 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993002.

Geman:1984:LTA

[GHR84] Donald Geman, Joseph Horowitz, and Jay Rosen. A local time analysis of intersections of Brownian paths in the plane. Annals of Probability, 12(1):86–107, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993375.

Gut:1986:CRE

[GJ86] Allan Gut and Svante Janson. Converse results for existence of moments and uniform integrability for stopped random walks. Annals of Probability, 14(4):1296–1317, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992371.

Griffin:1984:ALL

[GJP84] Philip S. Griffin, Naresh C. Jain, and William E. Pruitt. Approximate local limit theorems for laws outside domains of attraction. Annals of Probability, 12(1):45–63, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993373.

Gray:1980:AMS

[GK80] Robert M. Gray and J. C. Kieffer. Asymptotically mean stationary measures. Annals of Probability, 8(5):962–973, October 1980. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994624.

Goodman:1989:RCF

[GK89a] Victor Goodman and James Kuelbs. Rates of convergence for the functional LIL. Annals of Probability, 17(1):301–316, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991510.

Griffin:1989:SNL

[GK89b] Philip S. Griffin and James D. Kuelbs. Self-normalized laws of the iterated logarithm. Annals of Probability, 17(4):1571–1601, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991175.

Gandolfi:1988:UIO

[GKR88] A. Gandolfi, M. Keane, and L. Russo. On the uniqueness of the infinite occupied cluster in dependent two- dimensional site percolation. Annals of Probability, 16(3):1147–1157, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991681.

Goodman:1981:SRL

[GKZ81] V. Goodman, J. Kuelbs, and J. Zinn. Some results on the LIL in Banach space with applications to weighted empirical processes. Annals of Probability, 9(5):713–752, October 1981. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994305.

Griffeath:1982:CPS

[GL82] David Griffeath and Thomas M. Liggett. Critical phenomena for Spitzer’s reversible nearest particle systems. Annals of Probability, 10(4):881–895, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993711.

Glover:1980:CDP

[Glo80] Joseph Glover. Compactifications for dual processes. Annals of Probability, 8(6):1119–1134, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994573.

Glover:1981:ART

[Glo81a] Joseph Glover. Applications of raw time-changes to Markov processes. Annals of Probability, 9(6):1019–1029, December 1981. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994272.

Glover:1981:RTC

[Glo81b] Joseph Glover. Raw time changes of Markov processes. Annals of Probability, 9(1):90–102, February 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994510.

Gine:1983:CLT

[GM83a] Evarist Gine and Michel B. Marcus. The Central Limit Theorem for stochastic integrals with respect to Lévy processes. Annals of Probability, 11(1):58–77, February 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993660.

Goldberg:1983:BMG

[GM83b] S. I. Goldberg and C. Mueller. Brownian motion, geometry, and generalizations of Picard’s little theorem. Annals of Probability, 11 (4):833–846, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993435.

Gray:1980:BSS

[GOD80] R. M. Gray, D. S. Ornstein, and R. L. Dobrushin. Block synchronization, sliding-block coding, invulnerable sources and zero error codes for discrete noisy channels. Annals of Probability, 8 (4):639–674, August 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994658.

Goldman:1984:TSO

[Gol84] A. Goldman. Temps de séjour et oscillation du mouvement brownien au voisinage de la sphère euclidienne. (French) [Residence time and oscillation of Brownian movement in the neighborhood of the Euclidean sphere]. Annals of Probability, 12(3):829–842, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993231.

Goodman:1988:CNS

[Goo88] Victor Goodman. Characteristics of normal samples. Annals of Probability, 16(3):1281–1290, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991690.

Gorostiza:1983:HDL

[Gor83] Luis G. Gorostiza. High density limit theorems for infinite systems of unscaled branching Brownian motions. Annals of Probability, 11(2):374–392, May 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993603. See correction [Gor84].

Gorostiza:1984:CCB

[Gor84] Luis G. Gorostiza. Correction: Correction to “High Density Limit Theorems for Infinite Systems of Unscaled Branching Brownian Motions”. Annals of Probability, 12(3):926–927, August 1984. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993244. See

[Gor83].

Gordon:1988:GPA

[Gor88] Yehoram Gordon. Gaussian processes and almost spherical sections of convex bodies. Annals of Probability, 16(1):180–188, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991893.

Gotze:1981:EEB

[Got81] F. Gotze. On Edgeworth expansions in Banach spaces. Annals of Probability, 9(5):852–859, October 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994312.

Gotze:1986:RCC

[Got86] F. Gotze. On the rate of convergence in the Central Limit Theorem in Banach spaces. Annals of Probability, 14(3):922–942, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992448.

Gotze:1989:EEF

[Got89] F. Gotze. Edgeworth expansions in functional limit theorems. Annals of Probability, 17(4):1602–1634, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991176.

Galves:1987:MCD

[GOV87] Antonio Galves, Enzo Olivieri, and Maria Eulalia Vares. Metastability for a class of dynamical systems subject to small random perturbations. Annals of Probability, 15(4):1288–1305, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991977.

Greenwood:1983:CLT

[GP83a] Priscilla Greenwood and Edwin Perkins. A conditioned limit theorem for random walk and Brownian local time on square root boundaries. Annals of Probability, 11(2):227–261, May 1983. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993594.

Groeneboom:1983:ANS

[GP83b] Piet Groeneboom and Ronald Pyke. Asymptotic normality of statistics based on the convex minorants of empirical distribution functions. Annals of Probability, 11(2):328–345, May 1983. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993599.

Galves:1987:EFO

[GP87] Antonio Galves and Errico Presutti. Edge fluctuations for the one dimensional supercritical contact process. Annals of Probability, 15(3):1131–1145, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992086.

Griffin:1989:ANS

[GP89] Philip S. Griffin and William E. Pruitt. Asymptotic normality and subsequential limits of trimmed sums. Annals of Probability, 17 (3):1186–1219, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991264.

Glover:1986:HHH

[GR86] Joseph Glover and Murali Rao. Hunt’s hypothesis (h) and Getoor’s conjecture. Annals of Probability, 14(3):1085–1087, July 1986. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992463.

Goldie:1989:RPO

[GR89] Charles M. Goldie and Sidney Resnick. Records in a partially ordered set. Annals of Probability, 17(2):678–699, April 1989. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991421.

Gray:1986:DGA

[Gra86] Lawrence Gray. Duality for general attractive spin systems with applications in one dimension. Annals of Probability, 14(2):371– 396, April 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992522.

Grey:1980:NLC

[Gre80] D. R. Grey. A new look at convergence of branching processes. Annals of Probability, 8(2):377–380, April 1980. CODEN APBYAE.

ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994785.

Greven:1984:HBC

[Gre84] Andreas Greven. The hydrodynamical behavior of the coupled branching process. Annals of Probability, 12(3):760–767, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993226.

Greven:1985:CBP

[Gre85] A. Greven. The coupled branching process in random environment. Annals of Probability, 13(4):1133–1147, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992800.

Griffin:1982:RER

[Gri82] Philip S. Griffin. Rates of escape of random walks. Annals of Probability, 10(3):869–871, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993800.

Griffeath:1983:BCP

[Gri83a] David Griffeath. The binary contact path process. Annals of Probability, 11(3):692–705, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993514.

Griffin:1983:ITR

[Gri83b] Philip S. Griffin. An integral test for the rate of escape of ddimensional random walk. Annals of Probability, 11(4):953–961, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993444.

Griffin:1983:PES

[Gri83c] Philip S. Griffin. Probability estimates for the small deviations of d-dimensional random walk. Annals of Probability, 11(4):939–952, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993443.

Griffin:1986:MNS

[Gri86] Philip S. Griffin. Matrix normalized sums of independent identically distributed random vectors. Annals of Probability, 14(1): 224–246, January 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992624.

Grimmett:1987:TBR

[Gri87] G. R. Grimmett. Two books on random graphs. Annals of Probability, 15(4):1612–1617, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991998.

Groeneboom:1983:CMB

[Gro83] Piet Groeneboom. The concave majorant of Brownian motion. Annals of Probability, 11(4):1016–1027, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993450.

Grubel:1987:SDG

[Gru87] Rudolf Grubel. On subordinated distributions and generalized renewal measures. Annals of Probability, 15(1):394–415, January 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992278.

Ghoussoub:1980:VVS

[GS80] N. Ghoussoub and J. Michael Steele. Vector valued subadditive processes and applications in probability. Annals of Probability, 8 (1):83–95, February 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994826.

Groeneboom:1981:LDG

[GS81] Piet Groeneboom and Galen R. Shorack. Large deviations of goodness of fit statistics and linear combinations of order statistics. Annals of Probability, 9(6):971–987, December 1981. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994268.

Gilula:1985:MMD

[GS85] Zvi Gilula and Gideon Schwarz. On the maximum of a measure of deviation from independence between discrete random variables.

Annals of Probability, 13(1):314–317, February 1985. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993085.

Gundy:1980:LCC

[Gun80] Richard F. Gundy. Local convergence of a class of martingales in multidimensional time. Annals of Probability, 8(3):607–614, June 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994731.

Gut:1980:CRP

[Gut80] Allan Gut. Convergence rates for probabilities of moderate deviations for sums of random variables with multidimensional indices. Annals of Probability, 8(2):298–313, April 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994778.

Gut:1983:SLI

[Gut83] Allan Gut. Strong Laws for independent identically distributed random variables indexed by a sector. Annals of Probability, 11 (3):569–577, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993501.

Gut:1985:CCL

[Gut85] Allan Gut. On complete convergence in the Law of Large Numbers for subsequences. Annals of Probability, 13(4):1286–1291, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992812.

Glotzl:1982:BEE

[GW82] E. Glotzl and A. Wakolbinger. Bayes estimators and ergodic decomposability with an application to Cox processes. Annals of Probability, 10(3):872–876, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993801.

Gine:1984:SLT

[GZ84] Evarist Gine and Joel Zinn. Some limit theorems for empirical processes. Annals of Probability, 12(4):929–989, November 1984. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993138.

Gine:1986:EPI

[GZ86] Evarist Gine and Joel Zinn. Empirical processes indexed by Lipschitz functions. Annals of Probability, 14(4):1329–1338, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992373.

Gine:1987:LLN

[GZ87] Evarist Gine and Joel Zinn. The Law of Large Numbers for partial sum processes indexed by sets. Annals of Probability, 15(1): 154–163, January 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992261.

Gzyl:1980:IGT

[Gzy80] Henryk Gzyl. Infinitesimal generators of time changed processes. Annals of Probability, 8(4):716–726, August 1980. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994661.

Haeusler:1984:NRC

[Hae84] Erich Haeusler. A note on the rate of convergence in the martingale Central Limit Theorem. Annals of Probability, 12(2):635–639, May 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993311.

Haeusler:1988:RCC

[Hae88] Erich Haeusler. On the rate of convergence in the Central Limit Theorem for martingales with discrete and continuous time. Annals of Probability, 16(1):275–299, January 1988. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991901.

Haigh:1985:RRS

[Hai85] John Haigh. Rotational representations of stochastic matrices. Annals of Probability, 13(3):1024–1027, August 1985. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992926.

Hajek:1982:SEH

[Haj82] Bruce Hajek. Stochastic equations of hyperbolic type and a twoparameter Stratonovich calculus. Annals of Probability, 10(2):451– 463, May 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993869.

Hall:1980:CRC

[Hal80a] Peter Hall. Characterizing the rate of convergence in the Central Limit Theorem. Annals of Probability, 8(6):1037–1048, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994566.

Hall:1980:LBM

[Hal80b] Peter Hall. On the limiting behaviour of the mode and median of a sum of independent random variables. Annals of Probability, 8(3):419–430, June 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994717.

Hall:1981:CSR

[Hal81] Peter Hall. A converse to the Spitzer–Rosen theorem. Annals of Probability, 9(4):633–641, August 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994368.

Hall:1982:BRC

[Hal82a] Peter Hall. Bounds on the rate of convergence of moments in the Central Limit Theorem. Annals of Probability, 10(4):1004–1018, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993721.

Hall:1982:LTE

[Hal82b] Peter Hall. Limit theorems for estimators based on inverses of spacings of order statistics. Annals of Probability, 10(4):992–1003, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993720.

Hall:1982:RCW

[Hal82c] Peter Hall. On the rate of convergence in the weak law of large numbers. Annals of Probability, 10(2):374–381, May 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993863.

Hall:1983:CSA

[Hal83a] Peter Hall. Chi squared approximations to the distribution of a sum of independent random variables. Annals of Probability, 11(4): 1028–1036, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993451.

Hall:1983:SWD

[Hal83b] Peter Hall. Sets which determine the rate of convergence in the Central Limit Theorem. Annals of Probability, 11(2):355–361, May 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993601.

Hall:1984:IER

[Hal84] Peter Hall. On the influence of extremes on the rate of convergence in the Central Limit Theorem. Annals of Probability, 12 (1):154–172, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993380.

Hall:1985:CP

[Hal85a] Peter Hall. On continuum percolation. Annals of Probability, 13(4): 1250–1266, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992809.

Hall:1985:CDS

[Hal85b] Peter Hall. On the coverage of k-dimensional space by kdimensional spheres. Annals of Probability, 13(3):991–1002, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992920.

Hall:1986:CCM

[Hal86] Peter Hall. Clump counts in a Mosaic. Annals of Probability, 14(2):424–458, April 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992525.

Hall:1987:EES

[Hal87] Peter Hall. Edgeworth expansion for Student’s t statistic under minimal moment conditions. Annals of Probability, 15(3):920–931, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992073.

Hall:1988:ERN

[Hal88a] Peter Hall. On the effect of random norming on the rate of convergence in the Central Limit Theorem. Annals of Probability, 16 (3):1265–1280, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991689.

Hall:1988:RCB

[Hal88b] Peter Hall. Rate of convergence in bootstrap approximations. Annals of Probability, 16(4):1665–1684, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991590.

Hansen:1988:LCL

[Han88] Bjorn G. Hansen. On log-concave and log-convex infinitely divisible sequences and densities. Annals of Probability, 16(4):1832–1839, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991600.

Hansen:1989:FCL

[Han89] Jennie C. Hansen. A functional Central Limit Theorem for random mappings. Annals of Probability, 17(1):317–332, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.

aop/1176991511. See correction [Han91].

Hansen:1991:CFC

[Han91] Jennie C. Hansen. Correction: A functional Central Limit Theorem for random mappings. Annals of Probability, 19(3):1393–1396, July 1991. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176990350. See [Han89].

Harris:1981:BMH

[Har81] Theodore E. Harris. Brownian motions on the homeomorphisms of the plane. Annals of Probability, 9(2):232–254, April 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994465.

Hajek:1987:DTB

[HB87] Bruce Hajek and Toby Berger. A decomposition theorem for binary Markov random fields. Annals of Probability, 15(3):1112–1125, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992084.

Hebbar:1980:ASL

[Heb80] H. Vishnu Hebbar. Almost sure limit points of maxima of stationary Gaussian sequences. Annals of Probability, 8(2):393–399, April 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994788.

Heinkel:1984:LLN

[Hei84] Bernard Heinkel. On the Law of Large Numbers in 2-uniformly smooth Banach spaces. Annals of Probability, 12(3):851–857, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993233.

Helland:1981:MCW

[Hel81a] Inge S. Helland. Minimal conditions for weak convergence to a diffusion process on the line. Annals of Probability, 9(3):429–452, June 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994416.

Helmers:1981:BET

[Hel81b] R. Helmers. A Berry–Esseen theorem for linear combinations of order statistics. Annals of Probability, 9(2):342–347, April 1981. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994478.

Hendricks:1983:ULB

[Hen83] W. J. Hendricks. A uniform lower bound for Hausdorff dimension for transient symmetric Lévy processes. Annals of Probability, 11 (3):589–592, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993503.

Herrndorf:1983:SSM

[Her83] Norbert Herrndorf. Stationary strongly mixing sequences not satisfying the Central Limit Theorem. Annals of Probability, 11 (3):809–813, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993529.

Herrndorf:1984:ECL

[Her84a] Norbert Herrndorf. An example on the Central Limit Theorem for associated sequences. Annals of Probability, 12(3):912–917, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993241.

Herrndorf:1984:FCL

[Her84b] Norbert Herrndorf. A functional Central Limit Theorem for weakly dependent sequences of random variables. Annals of Probability, 12 (1):141–153, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993379.

Heunis:1986:PSD

[Heu86] Andrew J. Heunis. On the prevalence of stochastic differential equations with unique strong solutions. Annals of Probability, 14(2):653–662, April 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992535.

Hall:1981:RCM

[HH81] Peter Hall and C. C. Heyde. Rates of convergence in the martingale Central Limit Theorem. Annals of Probability, 9(3):395–404, June 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994413.

Hahn:1983:PTR

[HHK83] Marjorie G. Hahn, Peter Hahan, and Michael J. Klass. Pointwise translation of the Radon transform and the general central limit problem. Annals of Probability, 11(2):277–301, May 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993597.

Hannum:1981:DRR

[HHL81] Robert C. Hannum, Myles Hollander, and Naftali A. Langberg. Distributional results for random functionals of a Dirichlet process. Annals of Probability, 9(4):665–670, August 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994373.

Holmes:1982:OSL

[HHM82] J. P. Holmes, William N. Hudson, and J. David Mason. Operatorstable laws: Multiple exponents and elliptical symmetry. Annals of Probability, 10(3):602–612, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993770.

Hijab:1984:ABE

[Hij84] Omar Hijab. Asymptotic Bayesian estimation of a first order equation with small diffusion. Annals of Probability, 12(3):890–902, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993238.

Hill:1982:CGS

[Hil82] T. P. Hill. Conditional generalizations of Strong Laws which conclude the partial sums converge almost surely. Annals of Probability, 10(3):828–830, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993792.

Hill:1987:PGP

[Hil87] Theodore P. Hill. Partitioning general probability measures. Annals of Probability, 15(2):804–813, April 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992173.

Haeusler:1988:NBR

[HJ88] Erich Haeusler and Konrad Joos. A nonuniform bound on the rate of convergence in the martingale Central Limit Theorem. Annals of Probability, 16(4):1699–1720, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991592.

Hudson:1986:SGE

[HJV86] William N. Hudson, Zbigniew J. Jurek, and Jerry Alan Veeh. The symmetry group and exponents of operator stable probability measures. Annals of Probability, 14(3):1014–1023, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992455.

Hahn:1980:MNS

[HK80] Marjorie G. Hahn and Michael J. Klass. Matrix normalization of sums of random vectors in the domain of attraction of the multivariate normal. Annals of Probability, 8(2):262–280, April 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994776.

Hahn:1981:MCL

[HK81] Marjorie G. Hahn and Michael J. Klass. The multidimensional Central Limit Theorem for arrays normed by affine transformations. Annals of Probability, 9(4):611–623, August 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994366.

Hill:1982:CSR

[HK82] T. P. Hill and Robert P. Kertz. Comparisons of stop rule and supremum expectations of i.i.d. random variables. Annals of Probability, 10(2):336–345, May 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993861.

Hooghiemstra:1984:CLT

[HK84] G. Hooghiemstra and D. P. Kennedy. Conditioned limit theorems and heavy traffic. Annals of Probability, 12(2):631–634, May 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993310.

Hall:1980:IPC

[HKP80] Richard L. Hall, Marek Kanter, and Michael D. Perlman. Inequalities for the probability content of a rotated square and related convolutions. Annals of Probability, 8(4):802–813, August 1980. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994667.

Hahn:1987:ANT

[HKS87] M. G. Hahn, J. Kuelbs, and J. D. Samur. Asymptotic normality of trimmed sums of Φ-mixing random variables. Annals of Probability, 15(4):1395–1418, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991984.

Hill:1980:SLS

[HLS80] Bruce M. Hill, David Lane, and William Sudderth. A Strong Law for some generalized urn processes. Annals of Probability, 8(2):214–226, April 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994772.

Harrison:1985:SDR

[HLS85] J. M. Harrison, H. J. Landau, and L. A. Shepp. The stationary distribution of reflected Brownian motion in a planar region. Annals of Probability, 13(3):744–757, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992906.

Hill:1987:EUP

[HLS87] Bruce M. Hill, David Lane, and William Sudderth. Exchangeable urn processes. Annals of Probability, 15(4):1586–1592, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991995.

Hudson:1981:OSD

[HM81] William N. Hudson and J. David Mason. Operator-stable distribution on R^{2} with multiple exponents. Annals of Probability, 9(3):482–489, June 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994420.

Haeusler:1987:LIL

[HM87] Erich Haeusler and David M. Mason. A law of the iterated logarithm for sums of extreme values from a distribution with a regularly varying upper tail. Annals of Probability, 15(3):932–953, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992074.

Hudson:1983:DNA

[HMV83] William N. Hudson, J. David Mason, and Jerry Alan Veeh. The domain of normal attraction of an operator-stable law. Annals of Probability, 11(1):178–184, February 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993667.

Hogan:1986:CPC

[Hog86] Michael L. Hogan. Comments on a problem of Chernoff and Petkau. Annals of Probability, 14(3):1058–1063, July 1986. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992458.

Holst:1981:CCR

[Hol81a] Lars Holst. On convergence of the coverage by random arcs on a circle and the largest spacing. Annals of Probability, 9(4):648–655, August 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994370.

Holst:1981:SCL

[Hol81b] Lars Holst. Some conditional limit theorems in exponential families. Annals of Probability, 9(5):818–830, October 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994310.

Holte:1982:CMB

[Hol82] John M. Holte. Critical multitype branching processes. Annals of Probability, 10(2):482–495, May 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993871.

Holgate:1983:NBP

[Hol83] P. Holgate. The natural boundary problem for random power series with degenerate tail fields. Annals of Probability, 11(3):814–816, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993530.

Holley:1985:RCE

[Hol85] Richard Holley. Rapid convergence to equilibrium in one dimensional stochastic Ising models. Annals of Probability, 13(1):72–89, February 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993067.

Hollander:1988:MPR

[Hol88] W. Th. F. Den Hollander. Mixing properties for random walk in random scenery. Annals of Probability, 16(4):1788–1802, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991597.

Hoover:1984:SGM

[Hoo84] Douglas N. Hoover. Synonymity, generalized martingales, and subfiltration. Annals of Probability, 12(3):703–713, August 1984. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993221.

Hoover:1987:CAD

[Hoo87] Douglas N. Hoover. A characterization of adapted distribution. Annals of Probability, 15(4):1600–1611, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991997.

Horvath:1984:SAE

[Hor84] Lajos Horváth. Strong approximation of extended renewal processes. Annals of Probability, 12(4):1149–1166, November 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993145.

Hill:1983:AUN

[HP83] Theodore P. Hill and Victor C. Pestien. The advantage of using non-measurable stop rules. Annals of Probability, 11(2):442–450, May 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993609.

Haussmann:1986:TRD

[HP86] U. G. Haussmann and E. Pardoux. Time reversal of diffusions. Annals of Probability, 14(4):1188–1205, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992362.

Hanson:1981:LLN

[HR81a] D. L. Hanson and Ralph P. Russo. On the law of large numbers. Annals of Probability, 9(3):513–519, June 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994425.

Harrison:1981:RBM

[HR81b] J. Michael Harrison and Martin I. Reiman. Reflected Brownian motion on an orthant. Annals of Probability, 9(2):302–308, April 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994471.

Hanson:1983:SMR

[HR83a] D. L. Hanson and Ralph P. Russo. Some more results on increments of the Wiener process. Annals of Probability, 11(4): 1009–1015, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993449.

Hanson:1983:SRI

[HR83b] D. L. Hanson and Ralph P. Russo. Some results on increments of the Wiener process with applications to lag sums of i.i.d. random variables. Annals of Probability, 11(3):609–623, August 1983. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993505.

Hanson:1989:SLI

[HR89] D. L. Hanson and Ralph P. Russo. Some “LIM INF” results for increments of a Wiener process. Annals of Probability, 17 (3):1063–1082, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991257.

Harrison:1981:SBM

[HS81] J. M. Harrison and L. A. Shepp. On skew Brownian motion. Annals of Probability, 9(2):309–313, April 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994472.

Hsu:1989:HSC

[Hsu89] Pei Hsu. Heat semigroup on a complete Riemannian manifold. Annals of Probability, 17(3):1248–1254, July 1989. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991267.

Hagwood:1983:MCM

[HT83] Charles Hagwood and Henry Teicher. A multidimensional CLT for maxima of normed sums. Annals of Probability, 11(4):1048–1050, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993454.

Hu:1988:GKE

[Hu88] K. Y. Hu. A generalization of Kolmogorov’s extension theorem and an application to the construction of stochastic processes with random time domains. Annals of Probability, 16(1):222–230, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991896.

Huggins:1985:LIL

[Hug85] R. M. Huggins. Laws of the iterated logarithm for time changed Brownian motion with an application to branching processes. Annals of Probability, 13(4):1148–1156, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992801.

Hurzeler:1985:OST

[Hur85] Harry E. Hurzeler. The optional sampling theorem for processes indexed by a partially ordered set. Annals of Probability, 13(4): 1224–1235, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992807.

Hagwood:1982:EES

[HW82] Charles Hagwood and Michael Woodroofe. On the expansion for expected sample size in non-linear renewal theory. Annals of Probability, 10(3):844–848, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993795.

Harrison:1987:MRB

[HW87] J. M. Harrison and R. J. Williams. Multidimensional reflected Brownian motions having exponential stationary distributions.

Annals of Probability, 15(1):115–137, January 1987. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992259.

Hwang:1980:LMR

[Hwa80] Chii-Ruey Hwang. Laplace’s method revisited: Weak convergence of probability measures. Annals of Probability, 8(6):1177–1182, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994579.

Hwang:1987:PKA

[Hwa87] J. S. Hwang. On a problem of Kahane about the image of Gaussian Taylor series. Annals of Probability, 15(3):1203–1209, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992091.

Iscoe:1989:LDV

[IM89] I. Iscoe and D. McDonald. Large deviations for l^{2}-valued Ornstein– Uhlenbeck processes. Annals of Probability, 17(1):58–73, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991494.

Imhof:1985:RBM

[Imh85] J.-P. Imhof. On the range of Brownian motion and its inverse process. Annals of Probability, 13(3):1011–1017, August 1985. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992923.

Imkeller:1986:CTT

[Imk86] Peter Imkeller. On changing time for two-parameter strong martingales: A counterexample. Annals of Probability, 14(3):1080–1084, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992462.

Irle:1981:TPO

[Irl81] Albrecht Irle. Transitivity in problems of optimal stopping. Annals of Probability, 9(4):642–647, August 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994369.

Iscoe:1988:SMV

[Isc88] I. Iscoe. On the supports of measure-valued critical branching Brownian motion. Annals of Probability, 16(1):200–221, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991895.

Iscoe:1989:ASM

[Isc89] I. Iscoe. Addendum: On the supports of measure-valued critical branching Brownian motion. Annals of Probability, 17(2):813, April 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991429.

Iyengar:1986:LBM

[Iye86] Satish Iyengar. On a lower bound for the multivariate normal Mills’ ratio. Annals of Probability, 14(4):1399–1403, October 1986. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992380.

Jacobsen:1984:TOC

[Jac84] Martin Jacobsen. Two operational characterizations of cooptional times. Annals of Probability, 12(3):714–725, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993222.

Jajte:1987:EEH

[Jaj87] R. Jajte. On the existence of the ergodic Hilbert transform. Annals of Probability, 15(2):831–835, April 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992176.

Jakubowski:1986:PCL

[Jak86] Adam Jakubowski. Principle of conditioning in limit theorems for sums of random variables. Annals of Probability, 14(3):902–915, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992446.

Janson:1983:LTC

[Jan83a] Svante Janson. Limit theorems for certain branching random walks on compact groups and homogeneous spaces. Annals of Probability, 11(4):909–930, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993441.

Janson:1983:RTD

[Jan83b] Svante Janson. Renewal theory for M-dependent variables. Annals of Probability, 11(3):558–568, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993500.

Janson:1984:RDS

[Jan84] Svante Janson. Runs in m-dependent sequences. Annals of Probability, 12(3):805–818, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993229.

Janke:1985:RST

[Jan85] Steven J. Janke. Recurrent sets for transient Lévy processes with bounded kernels. Annals of Probability, 13(4):1204–1218, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992805.

Janson:1987:MSS

[Jan87] Svante Janson. Maximal spacings in several dimensions. Annals of Probability, 15(1):274–280, January 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992269.

Janson:1988:NCH

[Jan88] Svante Janson. Normal convergence by higher semiinvariants with applications to sums of dependent random variables and random graphs. Annals of Probability, 16(1):305–312, January 1988. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991903.

Joag-Dev:1983:IUU

[JD83] Kumar Joag-Dev. Independence via uncorrelatedness under certain dependence structures. Annals of Probability, 11(4):1037– 1041, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993452.

Joag-Dev:1983:ANR

[JDPP83] Kumar Joag-Dev, Michael D. Perlman, and Loren D. Pitt. Association of normal random variables and Slepian’s inequality. Annals of Probability, 11(2):451–455, May 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993610.

Jennen:1985:SOA

[Jen85] Christel Jennen. Second-order approximations to the density, mean and variance of Brownian first-exit times. Annals of Probability, 13 (1):126–144, February 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993071.

Jammalamadaka:1986:LTT

[JJ86] S. Rao Jammalamadaka and Svante Janson. Limit theorems for a triangular scheme of U-statistics with applications to interpoint distances. Annals of Probability, 14(4):1347–1358, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992375.

James:1988:CBC

[JJS88] Barry James, Kang Ling James, and David Siegmund. Conditional boundary crossing probabilities, with applications to change-point problems. Annals of Probability, 16(2):825–839, April 1988. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991789.

Johnson:1983:EFC

[JK83] B. McK. Johnson and T. Killeen. An explicit formula for the C.D.F. of the L_{1} norm of the Brownian bridge. Annals of Probability, 11(3):807–808, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993528.

Jain:1983:GM

[JM83] Naresh C. Jain and Ditlev Monrad. Gaussian measures in B_{p}. Annals of Probability, 11(1):46–57, February 1983. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993659.

Jagers:1985:BPP

[JN85] Peter Jagers and Olle Nerman. Branching processes in periodically varying environment. Annals of Probability, 13(1):254–268, February 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993079.

Jain:1980:DPA

[JO80] Naresh C. Jain and Steven Orey. Domains of partial attraction and tightness conditions. Annals of Probability, 8(3):584–599, June 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994728.

Joe:1987:MRD

[Joe87] Harry Joe. Majorization, randomness and dependence for multivariate distributions. Annals of Probability, 15(3):1217–1225, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992093.

Johnson:1985:BCM

[Joh85] W. B. Johnson. Best constants in moment inequalities for linear combinations of independent and exchangeable random variables. Annals of Probability, 13(1):234–253, February 1985. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993078.

Jain:1984:ABL

[JP84] Naresh C. Jain and William E. Pruitt. Asymptotic behavior of the local time of a recurrent random walk. Annals of Probability, 12 (1):64–85, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993374.

Jain:1987:LTP

[JP87a] Naresh C. Jain and William E. Pruitt. Lower tail probability estimates for subordinators and nondecreasing random walks. Annals of Probability, 15(1):75–101, January 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992257.

Jain:1987:MIL

[JP87b] Naresh C. Jain and William E. Pruitt. Maximal increments of local time of a random walk. Annals of Probability, 15(4):1461–1490, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991987.

Jacod:1988:TRL

[JP88] Jean Jacod and Philip Protter. Time reversal on Lévy processes. Annals of Probability, 16(2):620–641, April 1988. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991776.

Jureckova:1987:SOA

[JS87] Jana Jureckova and Pranab Kumar Sen. A second-order asymptotic distributional representation of M-estimators with discontinuous score functions. Annals of Probability, 15(2):814–823, April 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992174.

Johnson:1989:SIR

[JS89] William B. Johnson and G. Schechtman. Sums of independent random variables in rearrangement invariant function spaces. Annals of Probability, 17(2):789–808, April 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991427.

Jurek:1982:SCO

[Jur82] Zbigniew J. Jurek. Structure of a class of operator-selfdecomposable probability measures. Annals of Probability, 10(3):849–856, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993796.

Jurek:1985:RBS

[Jur85] Zbigniew J. Jurek. Relations between the s-selfdecomposable and selfdecomposable measures. Annals of Probability, 13(2):592–608, May 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993012.

Kalikow:1981:GRW

[Kal81a] Steven A. Kalikow. Generalized random walk in a random environment. Annals of Probability, 9(5):753–768, October 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994306.

Kallenberg:1981:SBT

[Kal81b] Olav Kallenberg. Splitting at backward times in regenerative sets. Annals of Probability, 9(5):781–799, October 1981. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994308.

Kallenberg:1987:HSM

[Kal87] Olav Kallenberg. Homogeneity and the strong Markov property. Annals of Probability, 15(1):213–240, January 1987. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992265.

Kallenberg:1988:SPS

[Kal88] Olav Kallenberg. Spreading and predictable sampling in exchangeable sequences and processes. Annals of Probability, 16(2):508–534, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991771.

Kanda:1989:NCM

[Kan89] Mamoru Kanda. A note on capacitary measures of semipolar sets. Annals of Probability, 17(1):379–384, January 1989. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991517.

Karandikar:1982:MDN

[Kar82] Rajeeva L. Karandikar. Multiplicative decomposition of nonsingular matrix valued continuous semimartingales. Annals of Probability, 10(4):1088–1091, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993734.

Karatzas:1984:GID

[Kar84] Ioannis Karatzas. Gittins indices in the dynamic allocation problem for diffusion processes. Annals of Probability, 12(1):173–192, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993381.

Kasahara:1982:LTS

[Kas82] Yuji Kasahara. A limit theorem for slowly increasing occupation times. Annals of Probability, 10(3):728–736, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993780.

Kaspi:1985:ELM

[Kas85] H. Kaspi. Excursion laws of Markov processes in classical duality. Annals of Probability, 13(2):492–518, May 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993005.

Kaspi:1988:RTC

[Kas88] H. Kaspi. Random time changes for processes with random birth and death. Annals of Probability, 16(2):586–599, April 1988. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991774.

Kaufman:1989:DPO

[Kau89] Robert Kaufman. Dimensional properties of one-dimensional Brownian motion. Annals of Probability, 17(1):189–193, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991503.

Kuelbs:1980:LLL

[KD80] J. Kuelbs and R. M. Dudley. Log log laws for empirical measures. Annals of Probability, 8(3):405–418, June 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994716.

Keener:1982:RTM

[Kee82] Robert W. Keener. Renewal theory for Markov chains on the real line. Annals of Probability, 10(4):942–954, November 1982. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993716.

Kelly:1987:ODC

[Kel87] F. P. Kelly. One-dimensional circuit-switched networks. Annals of Probability, 15(3):1166–1179, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992089.

Kent:1982:SDD

[Ken82] John T. Kent. The spectral decomposition of a diffusion hitting time. Annals of Probability, 10(1):207–219, February 1982. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993924.

Kennedy:1985:OSI

[Ken85] D. P. Kennedy. Optimal stopping of independent random variables and maximizing prophets. Annals of Probability, 13(2):566–571, May 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993009.

Kendall:1987:RPB

[Ken87] Wilfrid S. Kendall. The radial part of Brownian motion on a manifold: A semimartingale property. Annals of Probability, 15(4): 1491–1500, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991988.

Kent:1989:CPR

[Ken89] John T. Kent. Continuity properties for random fields. Annals of Probability, 17(4):1432–1440, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991163.

Kesten:1986:IMK

[Kes86] Harry Kesten. The influence of Mark Kac on probability theory. Annals of Probability, 14(4):1103–1128, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992356. See cor-

rection [Kes87a].

Kesten:1987:CIM

[Kes87a] Harry Kesten. Correction: The influence of Mark Kac on probability theory. Annals of Probability, 15(3):1228, July 1987. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992095. See

[Kes86].

Kesten:1987:PTF

[Kes87b] Harry Kesten. Percolation theory and first-passage percolation. Annals of Probability, 15(4):1231–1271, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991975.

Key:1984:RTC

[Key84] Eric S. Key. Recurrence and transience criteria for random walk in a random environment. Annals of Probability, 12(2):529–560, May 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993304.

Key:1987:LDR

[Key87] Eric S. Key. Limiting distributions and regeneration times for multitype branching processes with immigration in a random environment. Annals of Probability, 15(1):344–353, January 1987. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992273.

Key:1988:LEM

[Key88] Eric S. Key. Lyapunov exponents for matrices with invariant subspaces. Annals of Probability, 16(4):1721–1728, October 1988. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991593.

Kieffer:1980:CSP

[Kie80a] John C. Kieffer. On coding a stationary process to achieve a given marginal distribution. Annals of Probability, 8(1):131–141, February 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994829.

Kieffer:1980:TBS

[Kie80b] John C. Kieffer. On the transmission of Bernoulli sources over stationary channels. Annals of Probability, 8(5):942–961, October 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994623.

Kieffer:1984:SDT

[Kie84] John C. Kieffer. A simple development of the Thouvenot relative isomorphism theory. Annals of Probability, 12(1):204–211, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993383.

Kindermann:1980:ACF

[Kin80] Ross P. Kindermann. Asymptotic comparisons of functionals of Brownian motion and random walk. Annals of Probability, 8(6): 1135–1147, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994574.

Kipnis:1986:CLT

[Kip86] Claude Kipnis. Central Limit Theorems for infinite series of queues and applications to simple exclusion. Annals of Probability, 14 (2):397–408, April 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992523.

Kallianpur:1985:WNC

[KK85] G. Kallianpur and R. L. Karandikar. White noise calculus and nonlinear filtering theory. Annals of Probability, 13(4):1033–1107, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992798.

Keller:1987:ABD

[KKR87] G. Keller, G. Kersting, and U. Rosler. On the asymptotic behaviour of discrete time stochastic growth processes. Annals of Probability, 15(1):305–343, January 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992272.

Klass:1980:PBR

[Kla80] Michael J. Klass. Precision bounds for the relative error in the approximation of E|S_{n}| and extensions. Annals of Probability, 8(2):350–367, April 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994782.

Klass:1981:MAE

[Kla81] Michael J. Klass. A method of approximating expectations of functions of sums of independent random variables. Annals of Probability, 9(3):413–428, June 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994415.

Klass:1983:MRW

[Kla83] Michael J. Klass. On the maximum of a random walk with small negative drift. Annals of Probability, 11(3):491–505, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993498.

Klass:1984:FML

[Kla84a] Michael J. Klass. The finite mean LIL bounds are sharp. Annals of Probability, 12(3):907–911, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993240.

Klass:1984:MGR

[Kla84b] Michael J. Klass. The minimal growth rate of partial maxima. Annals of Probability, 12(2):380–389, May 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993296.

Klaassen:1985:IC

[Kla85a] Chris A. J. Klaassen. On an inequality of Chernoff. Annals of Probability, 13(3):966–974, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992917.

Klass:1985:RSS

[Kla85b] Michael J. Klass. The Robbins–Siegmund series criterion for partial maxima. Annals of Probability, 13(4):1369–1370, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992820.

Klass:1988:BPI

[Kla88] Michael J. Klass. A best possible improvement of Wald’s equation. Annals of Probability, 16(2):840–853, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991790.

Klass:1989:MPI

[Kla89] Michael J. Klass. Maximizing E max_{1≤k≤n} S_{k}^{+}/E_{n}^{+}: A prophet inequality for sums of i.i.d. mean zero variates. Annals of Probability, 17(3):1243–1247, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991266.

Klebaner:1989:GGN

[Kle89a] Fima C. Klebaner. Geometric growth in near-supercritical population size dependent multitype Galton–Watson processes. Annals of Probability, 17(4):1466–1477, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991167.

Klebaner:1989:SDE

[Kle89b] Fima C. Klebaner. Stochastic difference equations and generalized gamma distributions. Annals of Probability, 17(1):178–188, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991502.

Kliemann:1987:RIM

[Kli87] Wolfgang Kliemann. Recurrence and invariant measures for degenerate diffusions. Annals of Probability, 15(2):690–707, April 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992166.

Klein:1982:DIS

[KLS82] Abel Klein, Lawrence J. Landau, and David S. Shucker. Decoupling inequalities for stationary Gaussian processes. Annals of Probability, 10(3):702–708, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993778.

Kelly:1987:NPT

[KM87] F. P. Kelly and I. M. MacPhee. The number of packets transmitted by collision detect random access schemes. Annals of Probability, 15(4):1557–1568, October 1987. CODEN APBYAE.

ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991993.

Kaspi:1988:RSR

[KM88] H. Kaspi and B. Maisonneuve. Regenerative systems on the real line. Annals of Probability, 16(3):1306–1332, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991693.

Karlin:1988:MLC

[KO88a] Samuel Karlin and Friedemann Ost. Maximal length of common words among random letter sequences. Annals of Probability, 16 (2):535–563, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991772.

Kurtz:1988:UCC

[KO88b] T. G. Kurtz and D. L. Ocone. Unique characterization of conditional distributions in nonlinear filtering. Annals of Probability, 16 (1):80–107, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991887.

Kono:1982:AVS

[Kon82] Norio Kono. Another version of Strassen’s log log law with an application to approximate upper functions of a Gaussian process with a positive index. Annals of Probability, 10(2):303–319, May 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993859.

Korzeniowski:1984:MSB

[Kor84] Andrzej Korzeniowski. On Marcinkiewicz SLLN in Banach spaces. Annals of Probability, 12(1):279–280, February 1984. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993393.

Kotelenez:1986:LLN

[Kot86] Peter Kotelenez. Law of large numbers and Central Limit Theorem for linear chemical reactions with diffusion. Annals of Probability, 14(1):173–193, January 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992621.

Kuelbs:1980:ASI

[KP80] J. Kuelbs and Walter Philipp. Almost sure invariance principles for partial sums of mixing B-valued random variables. Annals of Probability, 8(6):1003–1036, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994565.

Kennedy:1989:TVD

[KQ89] J. E. Kennedy and M. P. Quine. The total variation distance between the binomial and Poisson distributions. Annals of Probability, 17(1):396–400, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991519.

Kieffer:1981:SUP

[KR81] John C. Kieffer and Maurice Rahe. Selecting universal partitions in ergodic theory. Annals of Probability, 9(4):705–709, August 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994379.

Kallianpur:1983:SM

[KR83] G. Kallianpur and D. Ramachandran. On the splicing of measures. Annals of Probability, 11(3):819–822, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993532.

Karatzas:1984:TDB

[KS84] Ioannis Karatzas and Steven E. Shreve. Trivariate density of Brownian motion, its local and occupation times, with application to stochastic control. Annals of Probability, 12(3):819–828, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993230.

Krakowiak:1986:RMF

[KS86] Wieslaw Krakowiak and Jerzy Szulga. Random multilinear forms. Annals of Probability, 14(3):955–973, July 1986. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992450.

Krengel:1987:PCG

[KS87] Ulrich Krengel and Louis Sucheston. Prophet compared to gambler: An inequality for transforms of processes. Annals of Probability, 15(4):1593–1599, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991996.

Krakowiak:1988:MSI

[KS88] Wieslaw Krakowiak and Jerzy Szulga. A multiple stochastic integral with respect to a strictly p-stable random measure. Annals of Probability, 16(2):764–777, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991786.

Klass:1987:CLT

[KT87] Michael Klass and Henry Teicher. The Central Limit Theorem for exchangeable random variables without moments. Annals of Probability, 15(1):138–153, January 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992260.

Kuczek:1982:CEA

[Kuc82] Thomas Kuczek. On the convergence of the empiric age distribution for one dimensional supercritical age dependent branching processes. Annals of Probability, 10(1):252–258, February 1982. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993930.

Kuczek:1989:CLT

[Kuc89] Thomas Kuczek. The Central Limit Theorem for the right edge of supercritical oriented percolation. Annals of Probability, 17(4): 1322–1332, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991157.

Kuelbs:1981:WCS

[Kue81] J. Kuelbs. When is the cluster set of S_{n}/a_{n} empty? Annals of Probability, 9(3):377–394, June 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994412.

Kuelbs:1985:LWD

[Kue85] J. Kuelbs. The LIL when X is in the domain of attraction of a Gaussian law. Annals of Probability, 13(3):825–859, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992910.

Kurtz:1980:OST

[Kur80a] Thomas G. Kurtz. The optional sampling theorem for martingales indexed by directed sets. Annals of Probability, 8(4):675–681, August 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994659.

Kurtz:1980:RMP

[Kur80b] Thomas G. Kurtz. Representations of Markov processes as multiparameter time changes. Annals of Probability, 8(4):682–715, August 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994660.

Kurtz:1981:CLT

[Kur81] Thomas G. Kurtz. The Central Limit Theorem for Markov chains. Annals of Probability, 9(4):557–560, August 1981. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994361.

Kurtz:1984:APC

[Kur84] Thomas G. Kurtz. Acknowledgment of priority: The Central Limit Theorem for Markov chains. Annals of Probability, 12(1):282, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993395.

Kuster:1985:AGC

[Kus85] Petra Kuster. Asymptotic growth of controlled Galton–Watson processes. Annals of Probability, 13(4):1157–1178, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992802.

Kaimanovich:1983:RWD

[KV83] V. A. Kaimanovich and A. M. Vershik. Random walks on discrete groups: Boundary and entropy. Annals of Probability, 11 (3):457–490, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993497.

Kwapien:1987:DSI

[KW87] Stanislaw Kwapien and Wojbor A. Woyczynski. Double stochastic integrals, random quadratic forms and random series in Orlicz spaces. Annals of Probability, 15(3):1072–1096, July 1987. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992082.

Kaufman:1988:EHS

[KW88] Robert Kaufman and Jang-Mei Wu. An example on highly singular parabolic measure. Annals of Probability, 16(4):1821–1831, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991599.

Kwapien:1987:DIP

[Kwa87] Stanislaw Kwapien. Decoupling inequalities for polynomial chaos. Annals of Probability, 15(3):1062–1071, July 1987. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992081.

Kuelbs:1983:SRL

[KZ83] J. Kuelbs and J. Zinn. Some results on LIL behavior. Annals of Probability, 11(3):506–557, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993499.

Lacey:1989:LIL

[Lac89] Michael T. Lacey. Laws of the iterated logarithm for the empirical characteristic function. Annals of Probability, 17(1):292–300, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991509.

Lai:1988:BCP

[Lai88] Tze Leung Lai. Boundary crossing problems for sample means. Annals of Probability, 16(1):375–396, January 1988. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991909.

Lalley:1984:CMR

[Lal84] S. P. Lalley. Conditional Markov renewal theory i. finite and denumerable state space. Annals of Probability, 12(4):1113–1148, November 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993144.

Lalley:1986:RRO

[Lal86] S. P. Lalley. Regenerative representation for one-dimensional Gibbs states. Annals of Probability, 14(4):1262–1271, October 1986. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992367.

Lawler:1989:ISA

[Law89] Gregory F. Lawler. The infinite self-avoiding walk in high dimensions. Annals of Probability, 17(4):1367–1376, October 1989. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991159.

LeJan:1981:JTC

[Le 81] Yves Le Jan. On the jumps of time changed diffusions. Annals of Probability, 9(6):1043–1044, December 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994276.

LeGall:1986:SWP

[Le 86] Jean-François Le Gall. Sur la saucisse de Wiener et les points multiples du mouvement brownien. (French) [On the Wiener sausage and the multiple points of Brownian movement]. Annals of Probability, 14(4):1219–1244, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992364.

LeGall:1988:FRW

[Le 88] Jean-François Le Gall. Fluctuation results for the Wiener sausage. Annals of Probability, 16(3):991–1018, July 1988. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991673.

Lee:1985:DRR

[Lee85a] Mei-Ling Ting Lee. Dependence by reverse regular rule. Annals of Probability, 13(2):583–591, May 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993011.

Lee:1985:DTP

[Lee85b] Mei-Ling Ting Lee. Dependence by total positivity. Annals of Probability, 13(2):572–582, May 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993010.

Lee:1988:LDN

[Lee88] Tzong-Yow Lee. Large deviations for noninteracting infinite particle systems. Annals of Probability, 16(4):1537–1558, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991582.

Lee:1989:LDS

[Lee89] Tzong-Yow Lee. Large deviations for systems of noninteracting recurrent particles. Annals of Probability, 17(1):46–57, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991493.

Li:1980:MAS

[Li80] Shuo-Yen Robert Li. A martingale approach to the study of occurrence of sequence patterns in repeated experiments. Annals of Probability, 8(6):1171–1176, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994578.

Liao:1985:DLE

[Lia85] Ming Liao. Domination of last exit distributions. Annals of Probability, 13(4):1358–1363, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992818.

Liao:1988:HDS

[Lia88] Ming Liao. Hitting distributions of small geodesic spheres. Annals of Probability, 16(3):1039–1050, July 1988. CODEN APBYAE.

ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991676.

Liggett:1980:LRE

[Lig80] Thomas M. Liggett. Long range exclusion processes. Annals of Probability, 8(5):861–889, October 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994618.

Liggett:1983:ANP

[Lig83a] Thomas M. Liggett. Attractive nearest particle systems. Annals of Probability, 11(1):16–33, February 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993656.

Liggett:1983:TCE

[Lig83b] Thomas M. Liggett. Two critical exponents for finite reversible nearest particle systems. Annals of Probability, 11(3):714–725, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993516.

Liggett:1985:ISE

[Lig85] Thomas M. Liggett. An improved subadditive ergodic theorem. Annals of Probability, 13(4):1279–1285, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992811.

Liggett:1989:ECA

[Lig89] Thomas M. Liggett. Exponential L_{2} convergence of attractive reversible nearest particle systems. Annals of Probability, 17(2):403– 432, April 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991408.

Lindgren:1980:MPN

[Lin80] Georg Lindgren. Model processes in nonlinear prediction with application to detection and alarm. Annals of Probability, 8(4): 775–792, August 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994665.

Lin:1981:MIH

[Lin81] T. F. Lin. Multiple integrals of a homogeneous process with independent increments. Annals of Probability, 9(3):529–532, June 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994427.

Lindgren:1985:OPL

[Lin85] Georg Lindgren. Optimal prediction of level crossings in Gaussian processes and sequences. Annals of Probability, 13(3):804–824, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992909.

Lalley:1986:CPA

[LL86] S. P. Lalley and G. Lorden. A control problem arising in the sequential design of experiments. Annals of Probability, 14(1): 136–172, January 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992620.

Lloyd:1984:OPD

[Llo84] Stuart P. Lloyd. Ordered prime divisors of a random integer. Annals of Probability, 12(4):1205–1212, November 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993149.

Langberg:1980:CNC

[LLP80] Naftali A. Langberg, Ramon V. Leon, and Frank Proschan. Characterization of nonparametric classes of life distributions. Annals of Probability, 8(6):1163–1170, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994577.

Liu:1989:EMV

[LM89] Li Liu and Carl Mueller. On the extinction of measure-valued critical branching Brownian motion. Annals of Probability, 17(4): 1463–1465, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991166.

Lorenzen:1981:OSS

[Lor81] Thomas J. Lorenzen. Optimal stopping with sampling cost: The secretary problem. Annals of Probability, 9(1):167–172, February 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994519.

Landers:1981:NM

[LR81] Dieter Landers and Lothar Rogge. The natural median. Annals of Probability, 9(6):1041–1042, December 1981. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994275. See cor-

rection [LR82].

Landers:1982:CNM

[LR82] D. Landers and L. Rogge. Correction: The natural median. Annals of Probability, 10(4):1092, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993735. See [LR81].

Leha:1984:DPT

[LR84] G. Leha and G. Ritter. On diffusion processes and their semigroups in Hilbert spaces with an application to interacting stochastic systems. Annals of Probability, 12(4):1077–1112, November 1984. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993143.

Landers:1986:IHM

[LR86a] D. Landers and L. Rogge. An inequality for the Hausdorff-metric of σ-fields. Annals of Probability, 14(2):724–730, April 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992541.

Landers:1986:SOA

[LR86b] Dieter Landers and Lothar Rogge. Second-order approximation in the conditional Central Limit Theorem. Annals of Probability, 14 (1):313–325, January 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992630.

Lindvall:1986:CMD

[LR86c] Torgny Lindvall and L. C. G. Rogers. Coupling of multidimensional diffusions by reflection. Annals of Probability, 14(3):860– 872, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992442.

Landers:1987:NEC

[LR87] Dieter Landers and Lothar Rogge. Nonuniform estimates in the conditional Central Limit Theorem. Annals of Probability, 15 (2):776–782, April 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992171.

Leadbetter:1988:ETS

[LR88] M. R. Leadbetter and Holger Rootzen. Extremal theory for stochastic processes. Annals of Probability, 16(2):431–478, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991767.

LeGall:1989:MPL

[LRS89] Jean-François Le Gall, Jay S. Rosen, and Narn-Rueih Shieh. Multiple points of Lévy processes. Annals of Probability, 17(2):503–515, April 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991412.

Langberg:1982:IML

[LS82] Naftali A. Langberg and Moshe Shaked. On the identifiability of multivariate life distribution functions. Annals of Probability, 10 (3):773–779, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993785.

Lalley:1987:CLT

[LS87a] S. P. Lalley and T. Sellke. A conditional limit theorem for the frontier of a branching Brownian motion. Annals of Probability, 15 (3):1052–1061, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992080.

Lynch:1987:LDP

[LS87b] James Lynch and Jayaram Sethuraman. Large deviations for processes with independent increments. Annals of Probability, 15 (2):610–627, April 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992161.

Lalley:1988:TWI

[LS88] S. Lalley and T. Sellke. Traveling waves in inhomogeneous branching Brownian motions. I. Annals of Probability, 16(3):1051–1062, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991677.

Lalley:1989:TWI

[LS89] S. Lalley and T. Sellke. Travelling waves in inhomogeneous branching Brownian motions. II. Annals of Probability, 17(1):116–127, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991498.

Ledoux:1986:CIP

[LT86] M. Ledoux and M. Talagrand. Conditions d’intégrabilité pour les multiplicateurs dans le TLC banachique. (French) [Integrability conditions for the multipliers in the Banach TLC]. Annals of Probability, 14(3):916–921, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992447.

Ledoux:1988:CLI

[LT88] M. Ledoux and M. Talagrand. Characterization of the law of the iterated logarithm in Banach spaces. Annals of Probability, 16 (3):1242–1264, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991688.

Ledoux:1989:CTR

[LT89] M. Ledoux and M. Talagrand. Comparison theorems, random geometry and some limit theorems for empirical processes. Annals of Probability, 17(2):596–631, April 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991418.

Luczak:1984:ESC

[Luc84] A. Luczak. Elliptical symmetry and characterization of operatorstable and operator semi-stable measures. Annals of Probability, 12(4):1217–1223, November 1984. CODEN APBYAE.

ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993151.

Lawler:1983:MSO

[LV83] G. F. Lawler and R. J. Vanderbei. Markov strategies for optimal control problems indexed by a partially ordered set. Annals of Probability, 11(3):642–647, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993508.

Lai:1982:LIL

[LW82] Tze Leung Lai and Ching Zong Wei. A law of the iterated logarithm for double arrays of independent random variables with applications to regression and time series models. Annals of Probability, 10(2):320–335, May 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993860.

Li:1989:LIL

[LW89] Deli Li and Zhiquan Wu. The law of the iterated logarithm for B-valued random variables with multidimensional indices. Annals of Probability, 17(2):760–774, April 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991425.

LePage:1981:CSD

[LWZ81] Raoul LePage, Michael Woodroofe, and Joel Zinn. Convergence to a stable distribution via order statistics. Annals of Probability, 9 (4):624–632, August 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994367.

Lynch:1983:SCE

[Lyn83] James Lynch. Some comments on the Erdős–Rényi Law and a theorem of Shepp. Annals of Probability, 11(3):801–802, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993526.

Lyons:1983:SCT

[Lyo83] Terry Lyons. A simple criterion for transience of a reversible Markov chain. Annals of Probability, 11(2):393–402, May 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993604.

MacQueen:1981:CP

[Mac81] J. MacQueen. Circuit processes. Annals of Probability, 9(4): 604–610, August 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994365.

Madras:1986:PRF

[Mad86] Neal Madras. A process in a randomly fluctuating environment. Annals of Probability, 14(1):119–135, January 1986. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992619.

Maharam:1982:OME

[Mah82] Dorothy Maharam. Orthogonal measures: An example. Annals of Probability, 10(3):879–880, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993803.

Maller:1980:NDP

[Mal80] R. A. Maller. A note on domains of partial attraction. Annals of Probability, 8(3):576–583, June 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994727.

Maller:1988:ANT

[Mal88] R. A. Maller. Asymptotic normality of trimmed means in higher dimensions. Annals of Probability, 16(4):1608–1622, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991587.

Mandelbaum:1987:CMA

[Man87] Avi Mandelbaum. Continuous multi-armed bandits and multiparameter processes. Annals of Probability, 15(4):1527–1556, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991992.

Marcus:1981:WCE

[Mar81] Michael B. Marcus. Weak convergence of the empirical characteristic function. Annals of Probability, 9(2):194–201, April 1981. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994461.

Martinsek:1982:MER

[Mar82] Adam T. Martinsek. Moments and error rates of two-sided stopping rules. Annals of Probability, 10(4):935–941, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993715.

Martinsek:1984:AOS

[Mar84] Adam T. Martinsek. Approximations to optimal stopping rules for exponential random variables. Annals of Probability, 12(3): 876–881, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993236.

March:1986:BMH

[Mar86] Peter March. Brownian motion and harmonic functions on rotationally symmetric manifolds. Annals of Probability, 14(3):793– 801, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992438.

Mason:1981:BWE

[Mas81] David M. Mason. Bounds for weighted empirical distribution functions. Annals of Probability, 9(5):881–884, October 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994315.

Mason:1982:LLN

[Mas82a] David M. Mason. Laws of large numbers for sums of extreme values. Annals of Probability, 10(3):754–764, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993783.

Mason:1982:SCS

[Mas82b] David M. Mason. Some characterizations of Strong Laws for linear functions of order statistics. Annals of Probability, 10(4):

1051–1057, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993727.

Mason:1984:WCW

[Mas84] David M. Mason. Weak convergence of the weighted empirical quantile process in L^{2}(0, 1). Annals of Probability, 12(1):243–255, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993387.

Mason:1989:EVE

[Mas89a] David M. Mason. An extended version of the Erdős–Rényi Strong Law of Large Numbers. Annals of Probability, 17(1):257–265, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991507.

Massart:1989:SAM

[Mas89b] Pascal Massart. Strong approximation for multivariate empirical and related processes, via KMT constructions. Annals of Probability, 17(1):266–291, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991508.

Mathai:1982:CGP

[Mat82] A. M. Mathai. On a conjecture in geometric probability regarding asymptotic normality of a random simplex. Annals of Probability, 10(1):247–251, February 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993929.

Matthews:1988:CPB

[Mat88a] Peter Matthews. Covering problems for Brownian motion on spheres. Annals of Probability, 16(1):189–199, January 1988. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991894.

Matthews:1988:CPM

[Mat88b] Peter Matthews. Covering problems for Markov chains. Annals of Probability, 16(3):1215–1228, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991686.

Mazziotto:1988:TPH

[Maz88] G. Mazziotto. Two-parameter hunt processes and a potential theory. Annals of Probability, 16(2):600–619, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991775.

McCormick:1980:ESL

[McC80a] William P. McCormick. An extension to a Strong Law result of Mittal and Ylvisaker for the maxima of stationary Gaussian processes. Annals of Probability, 8(3):498–510, June 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994724.

McCormick:1980:WCM

[McC80b] William P. McCormick. Weak convergence for the maxima of stationary Gaussian processes using random normalization. Annals of Probability, 8(3):483–497, June 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994723.

McConnell:1985:SAF

[McC85] Terry R. McConnell. The size of an analytic function as measured by Lévy’s time change. Annals of Probability, 13(3):1003–1005, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992921.

McConnell:1987:TPM

[McC87] Terry R. McConnell. A two-parameter maximal ergodic theorem with dependence. Annals of Probability, 15(4):1569–1585, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991994.

Miamee:1988:WMA

[Mia88] A. G. Miamee. On the Weiner–Masani algorithm for finding the generating function of multivariate stochastic processes. Annals of Probability, 16(4):1854–1859, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991602.

Michel:1982:GAR

[Mic82] R. Michel. Generalization and application of a result of C. C. Heyde. Annals of Probability, 10(4):1066–1068, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993729.

Mijnheer:1982:LP

[Mij82] Joop Mijnheer. Limit points of {n^{−1/α}S_{n}}. Annals of Probability, 10(2):382–395, May 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993864.

Millar:1981:CTS

[Mil81] P. W. Millar. Comparison theorems for sample function growth. Annals of Probability, 9(2):330–334, April 1981. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994476.

Millet:1985:RTO

[Mil85] Annie Millet. On randomized tactics and optimal stopping in the plane. Annals of Probability, 13(3):946–965, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992916.

Mitra:1981:DSS

[Mit81] Shashanka S. Mitra. Distribution of symmetric stable laws of index 2^{−n}. Annals of Probability, 9(4):710–711, August 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994380.

Mitra:1982:SLI

[Mit82] S. S. Mitra. Stable laws of index 2^{−n}. Annals of Probability, 10 (3):857–859, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993797.

Mitoma:1983:TP

[Mit83a] Itaru Mitoma. Tightness of probabilities on C([0, 1]; Y′) and D([0, 1]; Y′). Annals of Probability, 11(4):989–999, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993447.

Mitra:1983:APB

[Mit83b] S. S. Mitra. Acknowledgment of priority: “Distribution of Symmetric Stable Laws of Index 2^{−n}” and “Stable Laws of Index 2^{−n}”. Annals of Probability, 11(2):456, May 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993612.

Mitro:1983:ATR

[Mit83c] Joanna B. Mitro. An application of time reversal to Brownian local time. Annals of Probability, 11(1):225–226, February 1983. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993672.

Merzbach:1986:CSP

[MN86] Ely Merzbach and David Nualart. A characterization of the spatial Poisson process and changing time. Annals of Probability, 14(4): 1380–1390, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992378.

Merzbach:1988:MAP

[MN88] Ely Merzbach and David Nualart. A martingale approach to point processes in the plane. Annals of Probability, 16(1):265–274, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991900.

Millet:1989:IPT

[MNS89] A. Millet, D. Nualart, and M. Sanz. Integration by parts and time reversal for diffusion processes. Annals of Probability, 17 (1):208–238, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991505.

McLeish:1982:ERS

[MO82] D. L. McLeish and G. L. O’Brien. The expected ratio of the sum of squares to the square of the sum. Annals of Probability, 10(4): 1019–1028, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993722.

Marshall:1983:DAM

[MO83] Albert W. Marshall and Ingram Olkin. Domains of attraction of multivariate extreme value distributions. Annals of Probability, 11 (1):168–177, February 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993666.

Moricz:1983:CMO

[Mor83] Ferenc Moricz. On the Cesaro means of orthogonal sequences of random variables. Annals of Probability, 11(3):827–832, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993534.

Minami:1985:ABE

[MOT85] Nariyuki Minami, Yukio Ogura, and Matsuyo Tomisaki. Asymptotic behavior of elementary solutions of one-dimensional generalized diffusion equations. Annals of Probability, 13(3):698–715, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992904.

Mountford:1989:TIM

[Mou89a] T. S. Mountford. Time inhomogeneous Markov processes and the polarity of single points. Annals of Probability, 17(2):573–585, April 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991416.

Mountford:1989:UDR

[Mou89b] T. S. Mountford. Uniform dimension results for the Brownian sheet. Annals of Probability, 17(4):1454–1462, October 1989. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176991165.

Metivier:1980:SDI

[MP80] M. Metivier and J. Pellaumail. On a stopped Doob’s inequality and general stochastic equations. Annals of Probability, 8 (1):96–114, February 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994827.

Milman:1987:GPM

[MP87] V. D. Milman and G. Pisier. Gaussian processes and mixed volumes. Annals of Probability, 15(1):292–304, January 1987. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992271.

Mauldin:1983:OTK

[MPvW83] R. Daniel Mauldin, David Preiss, and Heinrich v. Weizsacker. Orthogonal transition kernels. Annals of Probability, 11(4):970–988, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993446.

Menaldi:1985:RDP

[MR85] Jose-Luis Menaldi and Maurice Robin. Reflected diffusion processes with jumps. Annals of Probability, 13(2):319–341, May 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992994.

Maejima:1987:IMR

[MR87] Makoto Maejima and Svetlozar T. Rachev. An ideal metric and the rate of convergence to a self-similar process. Annals of Probability, 15(2):708–727, April 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992167.

Major:1980:ECM

[MS80a] Peter Major and Domokos Szasz. On the effect of collisions on the motion of an atom in R^{1}. Annals of Probability, 8(6):1068–1078, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994568.

Millet:1980:CVC

[MS80b] Annie Millet and Louis Sucheston. A characterization of Vitali conditions in terms of maximal inequalities. Annals of Probability, 8(2):339–349, April 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994781.

Marshall:1982:CMN

[MS82a] Albert W. Marshall and Moshe Shaked. A class of multivariate new better than used distributions. Annals of Probability, 10(1): 259–264, February 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993931.

Mori:1982:ABS

[MS82b] T. F. Mori and G. J. Szekely. Asymptotic behaviour of symmetric polynomial statistics. Annals of Probability, 10(1):124–131, February 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993917.

Morrow:1989:LDR

[MS89] Gregory J. Morrow and Stanley Sawyer. Large deviation results for a class of Markov chains arising from population genetics. Annals of Probability, 17(3):1124–1146, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991260.

Moricz:1982:MPB

[MSS82] F. A. Moricz, R. J. Serfling, and W. F. Stout. Moment and probability bounds with quasi-superadditive structure for the maximum partial sum. Annals of Probability, 10(4):1032–1040, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993724.

McConnell:1986:DIM

[MT86] Terry R. McConnell and Murad S. Taqqu. Decoupling inequalities for multilinear forms in independent symmetric random variables. Annals of Probability, 14(3):943–954, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992449.

Mukerjee:1984:ASE

[Muk84] H. G. Mukerjee. Almost sure equiconvergence of conditional expectations. Annals of Probability, 12(3):733–741, August 1984. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993224.

Mason:1987:RKI

[MV87] David M. Mason and Willem R. Van Zwet. A refinement of the KMT inequality for the uniform empirical process. Annals of Probability, 15(3):871–884, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992070.

Marcus:1984:BLI

[MZ84] Michael B. Marcus and Joel Zinn. The bounded law of the iterated logarithm for the weighted empirical distribution process in the non-i.i.d. case. Annals of Probability, 12(2):335–360, May 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993294.

Nair:1981:ADM

[Nai81] K. Aiyappan Nair. Asymptotic distribution and moments of normal extremes. Annals of Probability, 9(1):150–153, February 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994515.

Nawrotzki:1982:FMC

[Naw82] Kurt Nawrotzki. Finite Markov chains in stationary random environments. Annals of Probability, 10(4):1041–1046, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993725.

Newman:1987:DBR

[New87] Charles M. Newman. Decomposition of binary random fields and zeros of partition functions. Annals of Probability, 15(3):1126– 1130, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992085.

Neyman:1982:RTS

[Ney82] Abraham Neyman. Renewal theory for sampling without replacement. Annals of Probability, 10(2):464–481, May 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993870.

Ney:1983:DPA

[Ney83] Peter Ney. Dominating points and the asymptotics of large deviations for random walk on R^{d}. Annals of Probability, 11(1): 158–167, February 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993665.

Ney:1984:CLD

[Ney84] Peter Ney. Convexity and large deviations. Annals of Probability, 12(3):903–906, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993239.

Ney:1986:TBL

[Ney86] Peter Ney. Two books on large deviations. Annals of Probability, 14 (4):1428–1431, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992384.

Nishioka:1981:DNP

[Nis81] Kunio Nishioka. The degenerate Neumann problem and degenerate diffusions with Venttsel’s boundary conditions. Annals of Probability, 9(1):103–118, February 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994511.

Nemetz:1983:MIL

[NK83] T. Nemetz and N. Kusolitsch. A method of investigating the longest paths in certain random graphs. Annals of Probability, 11 (1):217–221, February 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993670.

Ney:1987:MAPa

[NN87a] P. Ney and E. Nummelin. Markov additive processes I. Eigenvalue properties and limit theorems. Annals of Probability, 15 (2):561–592, April 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992159.

Ney:1987:MAPb

[NN87b] P. Ney and E. Nummelin. Markov additive processes II. Large deviations. Annals of Probability, 15(2):593–609, April 1987. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992160.

Nolan:1988:PPI

[Nol88] John P. Nolan. Path properties of index-β stable fields. Annals of Probability, 16(4):1596–1607, October 1988. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991586. See cor-

rection [Nol92].

Nolan:1992:CPP

[Nol92] John P. Nolan. Correction: Path properties of index-β stable fields. Annals of Probability, 20(3):1601–1602, July 1992. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989709. See

[Nol88].

Norberg:1984:CER

[Nor84] Tommy Norberg. Convergence and existence of random set distributions. Annals of Probability, 12(3):726–732, August 1984. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993223.

Nowak:1985:UMS

[Now85] Andrzej S. Nowak. Universally measurable strategies in zero-sum stochastic games. Annals of Probability, 13(1):269–287, February 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993080.

Nolan:1988:FLT

[NP88] Deborah Nolan and David Pollard. Functional limit theorems for U-processes. Annals of Probability, 16(3):1291–1298, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991691.

Neuhoff:1982:CEP

[NS82] David L. Neuhoff and Paul C. Shields. Channel entropy and primitive approximation. Annals of Probability, 10(1):188–198, February 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993922.

Nair:1986:NCE

[NSK86] Vijayan N. Nair, Lawrence A. Shepp, and Michael J. Klass. On the number of crossings of empirical distribution functions. Annals of Probability, 14(3):877–890, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992444.

Nualart:1984:QVT

[Nua84] D. Nualart. On the quadratic variation of two-parameter continuous martingales. Annals of Probability, 12(2):445–457, May 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993300.

Newman:1981:IPC

[NW81] C. M. Newman and A. L. Wright. An invariance principle for certain dependent sequences. Annals of Probability, 9(4):671–675, August 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994374.

Nualart:1989:RBS

[NZ89] D. Nualart and M. Zakai. On the relation between the Stratonovich and ogawa integrals. Annals of Probability, 17(4):1536–1540, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991172.

OBrien:1980:PIR

[O’B80] G. L. O’Brien. Pairwise independent random variables. Annals of Probability, 8(1):170–175, February 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994834.

OBrien:1987:EVS

[O’B87] George L. O’Brien. Extreme values for stationary and Markov sequences. Annals of Probability, 15(1):281–291, January 1987. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992270.

Oelschlager:1984:MAL

[Oel84] Karl Oelschlager. A martingale approach to the Law of Large Numbers for weakly interacting stochastic processes. Annals of Probability, 12(2):458–479, May 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993301.

Oelschlager:1988:HDP

[Oel88] Karl Oelschlager. Homogenization of a diffusion process in a divergence-free random field. Annals of Probability, 16(3):1084– 1126, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991679.

Okazaki:1981:BTM

[Oka81] Yoshiaki Okazaki. Bochner’s theorem on measurable linear functionals of a Gaussian measure. Annals of Probability, 9(4):663–664, August 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994372.

Okada:1985:DPP

[Oka85a] Norio Okada. On differentiability preserving properties of semigroups associated with one-dimensional singular diffusions. Annals of Probability, 13(1):206–225, February 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993076.

Okazaki:1985:BTM

[Oka85b] Yoshiaki Okazaki. Bochner’s theorem in measurable dual of type 2 Banach space. Annals of Probability, 13(3):1022–1023, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992925.

Omey:1988:RCD

[Ome88] E. Omey. Rates of convergence for densities in extreme value theory. Annals of Probability, 16(2):479–486, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991768.

Orey:1988:LDP

[OP88] Steven Orey and Stephan Pelikan. Large deviation principles for stationary processes. Annals of Probability, 16(4):1481–1495, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991579.

Odlyzko:1985:UHC

[OR85] A. M. Odlyzko and L. B. Richmond. On the unimodality of high convolutions of discrete distributions. Annals of Probability, 13 (1):299–306, February 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993082.

Ossiander:1987:CLT

[Oss87] Mina Ossiander. A Central Limit Theorem under metric entropy with L_{2} bracketing. Annals of Probability, 15(3):897–919, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992072.

OBrien:1985:SSP

[OV85] George L. O’Brien and Wim Vervaat. Self-similar processes with stationary increments generated by point processes. Annals of Probability, 13(1):28–52, February 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993064.

Peligrad:1982:IPM

[Pel82] Magda Peligrad. Invariance principles for mixing sequences of random variables. Annals of Probability, 10(4):968–981, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993718.

Peligrad:1985:IPM

[Pel85] Magda Peligrad. An invariance principle for ϕ-mixing sequences. Annals of Probability, 13(4):1304–1313, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992814.

Peligrad:1987:CLT

[Pel87] Magda Peligrad. On the Central Limit Theorem for ρ-mixing sequences of random variables. Annals of Probability, 15(4): 1387–1394, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991983.

Pemantle:1988:PTR

[Pem88] Robin Pemantle. Phase transition in reinforced random walk and RWRE on trees. Annals of Probability, 16(3):1229–1241, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991687.

Penrose:1989:ESI

[Pen89] M. D. Penrose. On the existence of self-intersections for quasi-every Brownian path in space. Annals of Probability, 17(2):482–502, April 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991411.

Perkins:1981:GIC

[Per81] Edwin Perkins. A global intrinsic characterization of Brownian local time. Annals of Probability, 9(5):800–817, October 1981. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994309.

Papamarcou:1986:NUL

[PF86] Adrianos Papamarcou and Terrence L. Fine. A note on undominated lower probabilities. Annals of Probability, 14(2):710–723, April 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992540.

Philipp:1980:WIP

[Phi80] Walter Philipp. Weak and L^{p}-invariance principles for sums of B-valued random variables. Annals of Probability, 8(1):68–82, February 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.

aop/1176994825. See correction [Phi86].

Philipp:1986:CCB

[Phi86] Walter Philipp. Correction: Correction to “Weak and L^{p}- Invariance Principles for Sums of B-Valued Random Variables”. Annals of Probability, 14(3):1095–1101, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992465. See

[Phi80].

Pickands:1986:CDD

[Pic86] James Pickands III. The continuous and differentiable domains of attraction of the extreme value distributions. Annals of Probability, 14(3):996–1004, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992453.

Pincus:1983:CBR

[Pin83] Steve Pincus. A class of Bernoulli random matrices with continuous singular stationary measures. Annals of Probability, 11(4): 931–938, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993442.

Pinsky:1985:CDPb

[Pin85a] Ross Pinsky. A classification of diffusion processes with boundaries by their invariant measures. Annals of Probability, 13(3): 693–697, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992903.

Pinsky:1985:FDP

[Pin85b] Ross Pinsky. The I-function for diffusion processes with boundaries. Annals of Probability, 13(3):676–692, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992902.

Pinsky:1985:EDV

[Pin85c] Ross Pinsky. On evaluating the Donsker–Varadhan I-function. Annals of Probability, 13(2):342–362, May 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992995.

Pinsky:1985:CDPa

[Pin85d] Ross G. Pinsky. On the convergence of diffusion processes conditioned to remain in a bounded region for large time to limiting positive recurrent diffusion processes. Annals of Probability, 13(2):363–378, May 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992996.

Pinsky:1986:SCF

[Pin86] Ross Pinsky. A spectral criterion for the finiteness or infiniteness of stopped Feynman–Kac functionals of diffusion processes. Annals of Probability, 14(4):1180–1187, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992361.

Pinsky:1987:RTB

[Pin87] Ross G. Pinsky. Recurrence, transience and bounded harmonic functions for diffusions in the plane. Annals of Probability, 15 (3):954–984, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992075. See correction [Pin90].

Pinsky:1988:MMV

[Pin88] Ross G. Pinsky. A mini-max variational formula giving necessary and sufficient conditions for recurrence or transience of multidimensional diffusion processes. Annals of Probability, 16(2):662– 671, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991779.

Pinsky:1989:APD

[Pin89] Ross Pinsky. The averaging principle for diffusions with a small parameter in the case of a noncharacteristic boundary. Annals of Probability, 17(2):559–572, April 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991415.

Pinsky:1990:CRT

[Pin90] Ross G. Pinsky. Correction: Recurrence, transience and bounded harmonic functions for diffusions in the plane. Annals of Probability, 18(1):438–440, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990958. See [Pin87].

Pittel:1981:LBP

[Pit81a] B. G. Pittel. Limiting behavior of a process of runs. Annals of Probability, 9(1):119–129, February 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994512.

Pittenger:1981:RBT

[Pit81b] A. O. Pittenger. Regular birth times for Markov processes. Annals of Probability, 9(5):769–780, October 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994307.

Pitt:1982:PCN

[Pit82] Loren D. Pitt. Positively correlated normal variables are associated. Annals of Probability, 10(2):496–499, May 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993872.

Pittel:1983:DRT

[Pit83] Boris Pittel. On distributions related to transitive closures of random finite mappings. Annals of Probability, 11(2):428–441, May 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993608.

Pittel:1985:AGC

[Pit85] B. Pittel. Asymptotical growth of a class of random trees. Annals of Probability, 13(2):414–427, May 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993000.

Pittel:1987:PAC

[Pit87] B. Pittel. On probabilistic analysis of a coalesced hashing algorithm. Annals of Probability, 15(3):1180–1202, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992090.

Pitt:1989:PHP

[Pit89] Loren D. Pitt. On a problem of h. p. McKean: Independence of Brownian hitting times and places. Annals of Probability, 17(4): 1651–1657, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991179.

Pollard:1982:CLT

[Pol82] David Pollard. A Central Limit Theorem for k-means clustering. Annals of Probability, 10(4):919–926, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993713.

Poor:1982:MLS

[Poo82] H. Vincent Poor. Minimax linear smoothing for capacities. Annals of Probability, 10(2):504–507, May 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993874.

Pourahmadi:1984:MCB

[Pou84] Mohsen Pourahmadi. On the mean convergence of the best linear interpolator of multivariate stationary stochastic processes. Annals of Probability, 12(2):609–614, May 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993308.

Puri:1983:SLLa

[PR83a] Madan L. Puri and Dan A. Ralescu. Strong Law of Large Numbers for Banach space valued random sets. Annals of Probability, 11 (1):222–224, February 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993671.

Puri:1983:SLLb

[PR83b] Madan L. Puri and Dan A. Ralescu. Strong Law of Large Numbers with respect to a set-valued probability measure. Annals of Probability, 11(4):1051–1054, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993455.

Puri:1985:CNF

[PR85] Madan L. Puri and Dan A. Ralescu. The concept of normality for fuzzy random variables. Annals of Probability, 13(4):1373–1379, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992822.

Prinz:1982:MGC

[Pri82] P. Prinz. Martingales with given convex image. Annals of Probability, 10(4):1085–1087, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993733.

Protter:1980:EKR

[Pro80] Philip Protter. An extension of Kazamaki’s results on BMO differentials. Annals of Probability, 8(6):1107–1118, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994572.

Protter:1985:ASS

[Pro85a] Philip Protter. Approximations of solutions of stochastic differential equations driven by semimartingales. Annals of Probability, 13 (3):716–743, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992905.

Protter:1985:VED

[Pro85b] Philip Protter. Volterra equations driven by semimartingales. Annals of Probability, 13(2):519–530, May 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993006.

Protter:1986:TBS

[Pro86] Philip Protter. Three books on stochastic integration. Annals of Probability, 14(1):343–346, January 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992633.

Pruitt:1981:GOS

[Pru81a] William E. Pruitt. General one-sided laws of the iterated logarithm. Annals of Probability, 9(1):1–48, February 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994508.

Pruitt:1981:GRW

[Pru81b] William E. Pruitt. The growth of random walks and Lévy processes. Annals of Probability, 9(6):948–956, December 1981. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994266.

Pruitt:1983:CLL

[Pru83] William E. Pruitt. The class of limit laws for stochastically compact normed sums. Annals of Probability, 11(4):962–969, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993445.

Pruitt:1987:CSS

[Pru87] William E. Pruitt. The contribution to the sum of the summand of maximum modulus. Annals of Probability, 15(3):885–896, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992071.

Presutti:1983:HVM

[PS83] Errico Presutti and Herbert Spohn. Hydrodynamics of the voter model. Annals of Probability, 11(4):867–875, November 1983. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993437.

Pruitt:1983:BAC

[PT83] William E. Pruitt and S. James Taylor. The behavior of asymmetric Cauchy processes for large time. Annals of Probability, 11(2):302–327, May 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993598.

Perkins:1988:MCA

[PT88] Edwin A. Perkins and S. James Taylor. Measuring close approaches on a Brownian path. Annals of Probability, 16(4):1458–1480, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991578.

Pakshirajan:1981:FLI

[PV81] R. P. Pakshirajan and R. Vasudeva. A functional law of the iterated logarithm for a class of subordinators. Annals of Probability, 9(6): 1012–1018, December 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994271.

Pitman:1986:ALP

[PY86a] Jim Pitman and Marc Yor. Asymptotic laws of planar Brownian motion. Annals of Probability, 14(3):733–779, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992436.

Pitman:1986:LCC

[PY86b] Jim Pitman and Marc Yor. Level crossings of a Cauchy process. Annals of Probability, 14(3):780–792, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992437.

Pitman:1989:FAL

[PY89] Jim Pitman and Marc Yor. Further asymptotic laws of planar Brownian motion. Annals of Probability, 17(3):965–1011, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991253.

Pyke:1980:ABS

[Pyk80] Ronald Pyke. The asymptotic behavior of spacings under Kakutani’s model for interval subdivision. Annals of Probability, 8(1): 157–163, February 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994832.

Quine:1982:BEB

[QR82] M. P. Quine and J. Robinson. A Berry–Esseen bound for an occupancy problem. Annals of Probability, 10(3):663–671, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993775.

Quine:1984:NAS

[QR84] M. P. Quine and J. Robinson. Normal approximations to sums of scores based on occupancy numbers. Annals of Probability, 12 (3):794–804, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993228.

Quine:1980:TLT

[Qui80] M. P. Quine. Three limit theorems for scores based on occupancy numbers. Annals of Probability, 8(1):148–156, February 1980. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994831.

Ramachandran:1981:NRC

[Ram81] D. Ramachandran. A note on regular conditional probabilities in Doob’s sense. Annals of Probability, 9(5):907–908, October 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994321.

Rousseau-Egele:1983:TLL

[RE83] J. Rousseau-Egele. Un théorême de la limite locale pour une classe de transformations dilatantes et monotones par morceaux. (French) [A local limit theorem for a classes of piecewise-monotone dilatory transformations]. Annals of Probability, 11(3):772–788, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993522.

Reeds:1980:MIM

[Ree80] James Reeds. Monotonicity of an integral of m. klass. Annals of Probability, 8(2):368–371, April 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994783.

Reiss:1981:APM

[Rei81] R.-D. Reiss. Approximation of product measures with an application to order statistics. Annals of Probability, 9(2):335–341, April 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994477.

Reich:1982:SRD

[Rei82a] Jakob I. Reich. Some results on distributions arising from coin tossing. Annals of Probability, 10(3):780–786, August 1982. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993786.

Reich:1982:WDW

[Rei82b] Jakob I. Reich. When do weighted sums of independent random variables have a density — some results and examples. Annals of Probability, 10(3):787–798, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993787.

Reid:1983:EMS

[Rei83] J. Reid. Estimate on moments of the solutions to stochastic differential equations in the plane. Annals of Probability, 11(3):656–668, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993510.

Reich:1986:DWS

[Rei86] Jakob I. Reich. C^{∞} densities for weighted sums of independent random variables. Annals of Probability, 14(3):1005–1013, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992454.

Ressel:1985:FTT

[Res85] Paul Ressel. De Finetti-type theorems: An analytical approach. Annals of Probability, 13(3):898–922, August 1985. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992913.

Revesz:1982:IWR

[Rev82] P. Revesz. On the increments of Wiener and related processes. Annals of Probability, 10(3):613–622, August 1982. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993771.

Rice:1982:IAV

[Ric82] S. O. Rice. The integral of the absolute value of the pinned Wiener process– calculation of its probability density by numerical integration. Annals of Probability, 10(1):240–243, February 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993927.

Ritter:1981:GRW

[Rit81] Grant A. Ritter. Growth of random walks conditioned to stay positive. Annals of Probability, 9(4):699–704, August 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994378.

Ramaswami:1980:DTP

[RN80] V. Ramaswami and Marcel F. Neuts. A duality theorem for phase type queues. Annals of Probability, 8(5):974–985, October 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994625.

Robinson:1977:LDP

[Rob77] J. Robinson. Large deviation probabilities for samples from a finite population. Annals of Probability, 5(6):913–925, December 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.

aop/1176995660. See correction [Rob83].

Robinson:1983:CCB

[Rob83] J. Robinson. Corrections: Correction to “Large Deviation Probabilities for Samples from a Finite Population”. Annals of Probability, 11(4):1055, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993456. See [Rob77].

Rogers:1981:CAD

[Rog81] L. C. G. Rogers. Characterizing all diffusions with the 2M − X property. Annals of Probability, 9(4):561–572, August 1981. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994362.

Rootzen:1980:LDE

[Roo80] Holger Rootzen. Limit distributions for the error in approximations of stochastic integrals. Annals of Probability, 8(2):241–251, April 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994774.

Rootzen:1986:EVT

[Roo86] Holger Rootzen. Extreme value theory for moving average processes. Annals of Probability, 14(2):612–652, April 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992534.

Rootzen:1987:RLT

[Roo87] Holger Rootzen. A ratio limit theorem for the tails of weighted sums. Annals of Probability, 15(2):728–747, April 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992168.

Rosler:1980:UPT

[Ros80] Uwe Rosler. Unimodality of passage times for one-dimensional strong Markov processes. Annals of Probability, 8(4):853–859, August 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994672.

Ross:1981:GPS

[Ros81] Sheldon M. Ross. Generalized Poisson shock models. Annals of Probability, 9(5):896–898, October 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994318.

Rosenkrantz:1983:CLT

[Ros83] Walter A. Rosenkrantz. Calculation of the Laplace transform of the length of the busy period for the M—G—1 queue via martingales. Annals of Probability, 11(3):817–818, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993531.

Rosen:1984:SIR

[Ros84a] Jay Rosen. Self-intersections of random fields. Annals of Probability, 12(1):108–119, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993376.

Rosenblatt:1984:ANS

[Ros84b] M. Rosenblatt. Asymptotic normality, strong mixing and spectral density estimates. Annals of Probability, 12(4):1167–1180, November 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993146.

Rosen:1985:RIL

[Ros85] Jay Rosen. A representation for the intersection local time of Brownian motion in space. Annals of Probability, 13(1):145–153, February 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993072.

Rosen:1986:TFR

[Ros86a] Jay Rosen. Tanaka’s formula and renormalization for intersections of planar Brownian motion. Annals of Probability, 14(4): 1245–1251, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992365.

Ross:1986:RSS

[Ros86b] David Ross. Random sets without separability. Annals of Probability, 14(3):1064–1069, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992459.

Rosen:1987:JCI

[Ros87] Jay Rosen. Joint continuity of the intersection local times of Markov processes. Annals of Probability, 15(2):659–675, April 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992164.

Rosen:1988:CSI

[Ros88] Jay Rosen. Continuity and singularity of the intersection local time of stable processes in R^{2}. Annals of Probability, 16(1):75–79, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991886.

Rogers:1981:MF

[RP81] L. C. G. Rogers and J. W. Pitman. Markov functions. Annals of Probability, 9(4):573–582, August 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994363.

Ramakrishnan:1980:SSB

[RR80] S. Ramakrishnan and B. V. Rao. B-spaces are standard Borel. Annals of Probability, 8(6):1191, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994582.

Rudolph:1980:SOS

[RS80] Daniel Rudolph and J. Michael Steele. Sizes of order statistical events of stationary processes. Annals of Probability, 8(6): 1079–1084, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994569.

Ruzsa:1982:ITR

[RS82] I. Z. Ruzsa and G. J. Szekely. Intersections of traces of random walks with fixed sets. Annals of Probability, 10(1):132–136, February 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993918.

Ramakrishnan:1986:EVE

[RS86a] S. Ramakrishnan and W. D. Sudderth. The expected value of an everywhere stopped martingale. Annals of Probability, 14 (3):1075–1079, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992461.

Rubin:1986:DSS

[RS86b] Herman Rubin and Thomas Sellke. On the distributions of sums of symmetric random variables and vectors. Annals of Probability, 14(1):247–259, January 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992625.

Robertson:1988:FTC

[RS88] James B. Robertson and Stephen Simons. A De Finetti theorem for a class of pairwise independent stationary processes. Annals of Probability, 16(1):344–354, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991906.

Rosalsky:1981:LTD

[RT81] Andrew Rosalsky and Henry Teicher. A limit theorem for double arrays. Annals of Probability, 9(3):460–467, June 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994418.

Rhee:1984:BRC

[RT84] WanSoo Rhee and Michel Talagrand. Bad rates of convergence for the Central Limit Theorem in Hilbert space. Annals of Probability, 12(3):843–850, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993232. See correction [RT89a].

Rhee:1989:CBR

[RT89a] WanSoo Rhee and Michel Talagrand. Correction: Bad rates of convergence for the Central Limit Theorem in Hilbert space. Annals of Probability, 17(1):401, January 1989. CODEN APBYAE.

ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991520. See [RT84].

Rhee:1989:SDI

[RT89b] WanSoo T. Rhee and Michel Talagrand. A sharp deviation inequality for the stochastic traveling salesman problem. Annals of Probability, 17(1):1–8, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991490.

Ruppert:1982:ASA

[Rup82] David Ruppert. Almost sure approximations to the Robbins– Monro and Kiefer–Wolfowitz processes with dependent noise. Annals of Probability, 10(1):178–187, February 1982. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993921.

Ruschendorf:1981:ODR

[Rüs81] Ludger Rüschendorf. Ordering of distributions and rearrangement of functions. Annals of Probability, 9(2):276–283, April 1981. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994468.

Russo:1988:SLQ

[Rus88] Ralph P. Russo. Strong Laws for quantiles corresponding to moving blocks of random variables. Annals of Probability, 16(1):162–171, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991891.

Rosinski:1986:ISI

[RW86] J. Rosinski and W. A. Woyczynski. On Itô stochastic integration with respect to p-stable motion: Inner clock, integrability of sample paths, double and multiple integrals. Annals of Probability, 14 (1):271–286, January 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992627.

Rachev:1989:RCN

[RY89] S. T. Rachev and J. E. Yukich. Rates for the CLT via new ideal metrics. Annals of Probability, 17(2):775–788, April 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991426.

Ryznar:1986:ABS

[Ryz86] M. Ryznar. Asymptotic behaviour of stable measures near the origin. Annals of Probability, 14(1):287–298, January 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992628.

Saada:1987:LTP

[Saa87] Ellen Saada. A limit theorem for the position of a tagged particle in a simple exclusion process. Annals of Probability, 15(1): 375–381, January 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992275.

Salisbury:1987:TPT

[Sal87] Thomas S. Salisbury. Three problems from the theory of right processes. Annals of Probability, 15(1):263–267, January 1987. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992267.

Samur:1984:CSM

[Sam84] Jorge D. Samur. Convergence of sums of mixing triangular arrays of random vectors with stationary rows. Annals of Probability, 12(2):390–426, May 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993297.

Samorodnitsky:1988:CGP

[Sam88] Gennady Samorodnitsky. Continuity of Gaussian processes. Annals of Probability, 16(3):1019–1033, July 1988. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991674.

Sanz:1988:LTT

[San88] Marta Sanz. Local time for two-parameter continuous martingales with respect to the quadratic variation. Annals of Probability, 16(2):778–792, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991787.

Sato:1981:GMD

[Sat81] Hiroshi Sato. Gaussian measurable dual and Bochner’s theorem. Annals of Probability, 9(4):656–662, August 1981. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994371.

Samuel-Cahn:1984:CTS

[SC84] Ester Samuel-Cahn. Comparison of threshold stop rules and maximum for independent nonnegative random variables. Annals of Probability, 12(4):1213–1216, November 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993150.

Scheutzow:1985:SEN

[Sch85a] Michael Scheutzow. Some examples of nonlinear diffusion processes having a time-periodic law. Annals of Probability, 13(2):379–384, May 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992997.

Schwarz:1985:MDU

[Sch85b] Gideon Schwarz. Multivariate distributions with uniformly distributed projections. Annals of Probability, 13(4):1371–1372, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992821.

Schonmann:1986:CLT

[Sch86] Roberto Henrique Schonmann. Central Limit Theorem for the contact process. Annals of Probability, 14(4):1291–1295, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992370.

Schonmann:1987:ASD

[Sch87a] Roberto Henrique Schonmann. Absence of a stationary distribution for the edge process of subcritical oriented percolation in two dimensions. Annals of Probability, 15(3):1146–1147, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992087.

Schonmann:1987:NPC

[Sch87b] Roberto Henrique Schonmann. A new proof of the complete convergence theorem for contact processes in several dimensions with large infection parameter. Annals of Probability, 15(1):382–387, January 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992276.

Schmidt:1989:LPU

[Sch89] Klaus D. Schmidt. The lattice property of uniform amarts. Annals of Probability, 17(1):372–378, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991516.

Sen:1982:RTU

[Sen82] Pranab Kumar Sen. A renewal theorem for an urn model. Annals of Probability, 10(3):838–843, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993794.

Serfozo:1982:FLT

[Ser82] Richard Serfozo. Functional limit theorems for extreme values of arrays of independent random variables. Annals of Probability, 10 (1):172–177, February 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993920.

Serfozo:1986:CPA

[Ser86] Richard F. Serfozo. Compound Poisson approximations for sums of random variables. Annals of Probability, 14(4):1391–1398, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.

aop/1176992379. See correction [Ser88].

Serfozo:1988:CCP

[Ser88] Richard F. Serfozo. Correction: Compound Poisson approximations for sums of random variables. Annals of Probability, 16 (1):429–430, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991914. See [Ser86].

Sharpe:1980:STD

[Sha80] M. J. Sharpe. Some transformations of diffusions by time reversal. Annals of Probability, 8(6):1157–1162, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994576.

Shaked:1983:ELF

[Sha83] Moshe Shaked. Exponential life functions with NBU components. Annals of Probability, 11(3):752–759, August 1983. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993519.

Shanthikumar:1988:DPF

[Sha88] J. George Shanthikumar. DFR property of first-passage times and its preservation under geometric compounding. Annals of Probability, 16(1):397–406, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991910.

Shao:1989:PCR

[Sha89] Qi-Man Shao. On a problem of Csörgő and Revesz. Annals of Probability, 17(2):809–812, April 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991428.

Shenker:1981:MNR

[She81] M. Shenker. The mean number of real roots for one class of random polynomials. Annals of Probability, 9(3):510–512, June 1981. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994424.

Shepp:1982:IAV

[She82a] L. A. Shepp. On the integral of the absolute value of the pinned Wiener process. Annals of Probability, 10(1):234–239, February 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993926.

Shepp:1982:XCF

[She82b] L. A. Shepp. The XYZ conjecture and the FKG inequality. Annals of Probability, 10(3):824–827, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993791.

Shimura:1983:CCL

[Shi83] Michio Shimura. A class of conditional limit theorems related to ruin problem. Annals of Probability, 11(1):40–45, February 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993658.

Shiryaev:1989:KLC

[Shi89] A. N. Shiryaev. Kolmogorov: Life and creative activities. Annals of Probability, 17(3):866–944, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991251.

Siegmund:1982:LDB

[Sie82] D. Siegmund. Large deviations for boundary crossing probabilities. Annals of Probability, 10(3):581–588, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993768.

Siegrist:1983:HFD

[Sie83] Kyle Siegrist. Harmonic functions and the Dirichlet problem for revival Markov processes. Annals of Probability, 11(3):624–634, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993506.

Siegmund:1988:ATP

[Sie88] David Siegmund. Approximate tail probabilities for the maxima of some random fields. Annals of Probability, 16(2):487–501, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991769.

Silverstein:1980:CCC

[Sil80] Martin L. Silverstein. Classification of coharmonic and coinvariant functions for a Lévy process. Annals of Probability, 8(3):539–575, June 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994726.

Silverman:1983:CCE

[Sil83] B. W. Silverman. Convergence of a class of empirical distribution functions of dependent random variables. Annals of Probability, 11(3):745–751, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993518.

Silverstein:1985:SEL

[Sil85] Jack W. Silverstein. The smallest eigenvalue of a large dimensional Wishart matrix. Annals of Probability, 13(4):1364–1368, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992819.

Sine:1988:RUK

[Sin88] Robert Sine. Review: Ulrich Krengel, Ergodic Theorems. Annals of Probability, 16(1):424–428, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991913.

Sinai:1989:KWE

[Sin89] Ya. G. Sinai. Kolmogorov’s work on ergodic theory. Annals of Probability, 17(3):833–839, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991249.

Slaby:1988:UBL

[Sla88] M. Slaby. On the upper bound for large deviations of sums of i.i.d. random vectors. Annals of Probability, 16(3):978–990, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991672.

Slade:1989:SLS

[Sla89] Gordon Slade. The scaling limit of self-avoiding random walk in high dimensions. Annals of Probability, 17(1):91–107, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991496.

Smith:1982:ADS

[Smi82] Richard L. Smith. The asymptotic distribution of the strength of a series-parallel system with equal load-sharing. Annals of Probability, 10(1):137–171, February 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993919.

Smythe:1980:MPS

[Smy80] R. T. Smythe. Maxima of partial sums and a monotone regression estimator. Annals of Probability, 8(3):630–635, June 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994734.

Solomon:1980:RMR

[Sol80] Frederick Solomon. A renewal model with randomly selected parameters. Annals of Probability, 8(3):622–629, June 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994733.

Soltani:1984:EMA

[Sol84] A. Reza Soltani. Extrapolation and moving average representation for stationary random fields and Beurling’s theorem. Annals of Probability, 12(1):120–132, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993377.

Schaufele:1981:CNC

[SP81] Ronald A. Schaufele and Ronald Pyke. Correction note: Correction to “The Existence and Uniqueness of Stationary Measures for Markov-Renewal Processes”. Annals of Probability, 9(2):348, April 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994479.

Spitzer:1981:ISL

[Spi81] Frank Spitzer. Infinite systems with locally interacting components. Annals of Probability, 9(3):349–364, June 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994410.

Seoh:1985:CTL

[SRP85] Munsup Seoh, Stefan S. Ralescu, and Madan L. Puri. Cramér type large deviations for generalized rank statistics. Annals of Probability, 13(1):115–125, February 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993070.

Samuels:1981:OSS

[SS81] Stephen M. Samuels and J. Michael Steele. Optimal sequential selection of a monotone sequence from a random sample. Annals of Probability, 9(6):937–947, December 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994265.

Samorodnitsky:1989:AET

[SS89] Gennady Samorodnitsky and Jerzy Szulga. An asymptotic evaluation of the tail of a multiple symmetric α- stable integral. Annals of Probability, 17(4):1503–1520, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991170.

Stadtmuller:1984:NLI

[Sta84] Ulrich Stadtmuller. A note on the law of iterated logarithm for weighted sums of random variables. Annals of Probability, 12 (1):35–44, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993372.

Steele:1981:SEF

[Ste81] J. Michael Steele. Subadditive Euclidean functionals and nonlinear growth in geometric probability. Annals of Probability, 9 (3):365–376, June 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994411.

Steele:1982:OTR

[Ste82] J. Michael Steele. Optimal triangulation of random samples in the plane. Annals of Probability, 10(3):548–553, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993766.

Steele:1986:FID

[Ste86a] J. Michael Steele. Fisher information and detection of a Euclidean perturbation of an independent stationary process. Annals of Probability, 14(1):326–335, January 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992631.

Steinebach:1986:IER

[Ste86b] J. Steinebach. Improved Erdős–Rényi and strong approximation laws for increments of renewal processes. Annals of Probability, 14(2):547–559, April 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992530.

Steele:1988:GRE

[Ste88] J. Michael Steele. Growth rates of Euclidean minimal spanning trees with power weighted edges. Annals of Probability, 16(4): 1767–1787, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991596.

Stricker:1986:RPC

[Str86a] Christophe Stricker. Representation previsible et changement de temps. (French) [Predictable representation and time change]. Annals of Probability, 14(3):1070–1074, July 1986. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992460.

Stroock:1986:RWH

[Str86b] Daniel W. Stroock. On the rate at which a homogeneous diffusion approaches a limit, an application of large deviation theory to certain stochastic integrals. Annals of Probability, 14(3):840–859, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992441.

Stute:1982:LLK

[Stu82a] Winfried Stute. A law of the logarithm for kernel density estimators. Annals of Probability, 10(2):414–422, May 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993866.

Stute:1982:OBE

[Stu82b] Winfried Stute. The oscillation behavior of empirical processes. Annals of Probability, 10(1):86–107, February 1982. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993915.

Stute:1984:OBE

[Stu84] Winfried Stute. The oscillation behavior of empirical processes: The multivariate case. Annals of Probability, 12(2):361–379, May 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993295.

Stute:1986:ASC

[Stu86] Winfried Stute. On almost sure convergence of conditional empirical distribution functions. Annals of Probability, 14(3):891–901, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992445.

Shorack:1982:LTI

[SW82] Galen R. Shorack and Jon A. Wellner. Limit theorems and inequalities for the uniform empirical process indexed by intervals. Annals of Probability, 10(3):639–652, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993773.

Stout:1984:TAU

[SW84] Quentin F. Stout and Bette Warren. Tree algorithms for unbiased coin tossing with a biased coin. Annals of Probability, 12 (1):212–222, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993384.

Sweeting:1980:SCA

[Swe80] T. J. Sweeting. Speeds of convergence and asymptotic expansions in the Central Limit Theorem: A treatment by operators. Annals of Probability, 8(2):281–297, April 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994777.

Sweeting:1985:DUL

[Swe85] T. J. Sweeting. On domains of uniform local attraction in extreme value theory. Annals of Probability, 13(1):196–205, February 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993075.

Takahata:1981:BAN

[Tak81] Hiroshi Takahata. L_{∞}-Bound for asymptotic normality of weakly dependent summands using Stein’s result. Annals of Probability, 9 (4):676–683, August 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994375.

Taksar:1982:FHT

[Tak82] M. I. Taksar. First hitting time of curvilinear boundary by Wiener process. Annals of Probability, 10(4):1029–1031, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993723.

Talagrand:1985:SSR

[Tal85] Michel Talagrand. Some structure results for martingales in the limit and pramarts. Annals of Probability, 13(4):1192–1203, November 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992804.

Talagrand:1987:DCR

[Tal87a] Michel Talagrand. Donsker classes and random geometry. Annals of Probability, 15(4):1327–1338, October 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991979.

Talagrand:1987:GCP

[Tal87b] Michel Talagrand. The Glivenko–Cantelli problem. Annals of Probability, 15(3):837–870, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992069.

Talagrand:1987:MPE

[Tal87c] Michel Talagrand. Measurability problems for empirical processes. Annals of Probability, 15(1):204–212, January 1987. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992264.

Talagrand:1988:NCS

[Tal88a] Michel Talagrand. Necessary conditions for sample boundedness of p-stable processes. Annals of Probability, 16(4):1584–1595, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991585.

Talagrand:1988:SSI

[Tal88b] Michel Talagrand. The structure of sign-invariant GB-sets and of certain Gaussian measures. Annals of Probability, 16(1):172–179, January 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991892.

Talagrand:1989:IIS

[Tal89] Michel Talagrand. Isoperimetry and integrability of the sum of independent Banach-space valued random variables. Annals of Probability, 17(4):1546–1570, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991174.

Taylor:1989:MEC

[Tay89] J. C. Taylor. The minimal eigenfunctions characterize the Ornstein–Uhlenbeck process. Annals of Probability, 17(3):1055– 1062, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991256.

Taylor:1983:ASC

[TC83] R. L. Taylor and C. A. Calhoun. On the almost sure convergence of randomly weighted sums of random elements. Annals of Probability, 11(3):795–797, August 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993524.

Tchen:1980:IDG

[Tch80] Andre H. Tchen. Inequalities for distributions with given marginals. Annals of Probability, 8(4):814–827, August 1980. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994668.

Terrell:1983:CRD

[Ter83] George R. Terrell. A characterization of rectangular distributions. Annals of Probability, 11(3):823–826, August 1983. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993533.

Teugels:1981:LTO

[Teu81] Jozef L. Teugels. Limit theorems on order statistics. Annals of Probability, 9(5):868–880, October 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994314.

Thelen:1989:FID

[The89] Brian J. Thelen. Fisher information and dichotomies in equivalence/contiguity. Annals of Probability, 17(4):1664–1690, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991181.

Thompson:1986:CMK

[Tho86a] Colin J. Thompson. The contributions of mark Kac to mathematical physics. Annals of Probability, 14(4):1129–1138, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992357.

Thorisson:1986:MDC

[Tho86b] Hermann Thorisson. On maximal and distributional coupling. Annals of Probability, 14(3):873–876, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992443.

Thorisson:1988:BL

[Tho88] Hermann Thorisson. Backward limits. Annals of Probability, 16 (2):914–924, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991796.

Tomkins:1980:LTM

[Tom80] R. J. Tomkins. Limit theorems without moment hypotheses for sums of independent random variables. Annals of Probability, 8 (2):314–324, April 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994779.

Tomkins:1986:RVS

[Tom86] R. J. Tomkins. Regular variation and the stability of maxima. Annals of Probability, 14(3):984–995, July 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992452.

Tucker:1980:JLL

[Tuc80] Howard G. Tucker. Joint limit laws of sample moments of a symmetric distribution. Annals of Probability, 8(5):991–998, October 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994627.

Tudor:1989:CTS

[Tud89] Constantin Tudor. A comparison theorem for stochastic equations with Volterra drifts. Annals of Probability, 17(4):1541–1545, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991173.

Uchiyama:1982:PBS

[Uch82a] Kohei Uchiyama. The proportion of Brownian sojourn outside a moving boundary. Annals of Probability, 10(1):220–233, February 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993925.

Uchiyama:1982:SGB

[Uch82b] Kohei Uchiyama. Spatial growth of a branching process of particles living in R^{d}. Annals of Probability, 10(4):896–918, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993712.

Uhrin:1984:SRA

[Uhr84] Bela Uhrin. Some remarks about the convolution of unimodal functions. Annals of Probability, 12(2):640–645, May 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993312.

Ulbricht:1981:WSI

[Ulb81] Ralf Ulbricht. Weighted sums of independent identically distributed random variables. Annals of Probability, 9(4):693–698, August 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994377.

Ustunel:1984:APN

[Ust84] A. S. Ustunel. Additive processes on nuclear spaces. Annals of Probability, 12(3):858–868, August 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993234.

Ustunel:1989:ICW

[UZ89] Ali Suleyman Ustunel and Moshe Zakai. On independence and conditioning on Wiener space. Annals of Probability, 17(4):1441–1453, October 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991164.

VanZwet:1980:SLL

[Van80] W. R. Van Zwet. A Strong Law for linear functions of order statistics. Annals of Probability, 8(5):986–990, October 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994626.

Vanderbei:1984:PSD

[Van84] R. J. Vanderbei. Probabilistic solution of the Dirichlet problem for biharmonic functions in discrete space. Annals of Probability, 12(2):311–324, May 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993292.

Vasicek:1980:CLL

[Vas80] Oldrich Alfonso Vasicek. A conditional law of large numbers. Annals of Probability, 8(1):142–147, February 1980. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994830.

vanEs:1986:WLE

[vE86] Bert van Es. On the weak limits of elementary symmetric polynomials. Annals of Probability, 14(2):677–695, April 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176992537.

Vervaat:1985:SPP

[Ver85] Wim Vervaat. Sample path properties of self-similar processes with stationary increments. Annals of Probability, 13(1):1–27, February 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993063.

VanDenBerg:1987:CCC

[VF87] J. Van Den Berg and U. Fiebig. On a combinatorial conjecture concerning disjoint occurrences of events. Annals of Probability, 15(1):354–374, January 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992274.

Venkataraman:1982:LTS

[VN82a] K. N. Venkataraman and K. Nanthi. A limit theorem on a subcritical Galton–Watson process with immigration. Annals of Probability, 10(4):1069–1074, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993730.

Venkataraman:1982:SLT

[VN82b] K. N. Venkataraman and K. Nanthi. Some limit theorems on a supercritical simple Galton–Watson process. Annals of Probability, 10(4):1075–1078, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993731.

vanPutten:1985:IPC

[vPvS85] C. van Putten and J. H. van Schuppen. Invariance properties of the conditional independence relation. Annals of Probability, 13 (3):934–945, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992915.

Vandemaele:1982:CTL

[VV82] M. Vandemaele and N. Veraverbeke. Cramér type large deviations for linear combinations of order statistics. Annals of Probability, 10(2):423–434, May 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993867.

vanZuijlen:1982:PED

[vZ82] Martien C. A. van Zuijlen. Properties of the empirical distribution function for independent non- identically distributed random vectors. Annals of Probability, 10(1):108–123, February 1982. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993916.

Walsh:1982:PSB

[Wal82] John B. Walsh. Propagation of singularities in the Brownian sheet. Annals of Probability, 10(2):279–288, May 1982. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993857.

Watkins:1984:CLP

[Wat84] Joseph C. Watkins. A central limit problem in random evolutions. Annals of Probability, 12(2):480–513, May 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993302.

Watkins:1985:SIR

[Wat85] Joseph C. Watkins. A stochastic integral representation for random evolutions. Annals of Probability, 13(2):531–557, May 1985. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993007.

Watanabe:1987:AWF

[Wat87] Shinzo Watanabe. Analysis of Wiener functionals (Malliavin calculus) and its applications to heat kernels. Annals of Probability, 15(1):1–39, January 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992255.

Watkins:1989:DIP

[Wat89] Joseph C. Watkins. Donsker’s invariance principle for Lie groups. Annals of Probability, 17(3):1220–1242, July 1989. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991265.

Waymire:1980:ZRI

[Way80] Ed Waymire. Zero-range interaction at Bose–Einstein speeds under a positive recurrent single particle law. Annals of Probability, 8(3):441–450, June 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994719.

Weber:1988:SPV

[Web88] Michel Weber. Stochastic processes with value in exponential type Orlicz spaces. Annals of Probability, 16(3):1365–1371, July 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991696.

Weis:1984:COT

[Wei84] Lutz W. Weis. A characterization of orthogonal transition kernels. Annals of Probability, 12(4):1224–1227, November 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993152.

Weis:1985:TEC

[Wei85] Lutz W. Weis. Two examples concerning a theorem of Burgess and Mauldin. Annals of Probability, 13(3):1028–1031, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992927.

Wendel:1980:HSB

[Wen80a] J. G. Wendel. Hitting spheres with Brownian motion. Annals of Probability, 8(1):164–169, February 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994833.

Wendel:1980:IPB

[Wen80b] J. G. Wendel. An independence property of Brownian motion with drift. Annals of Probability, 8(3):600–601, June 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994729.

Whittle:1981:AAB

[Whi81] P. Whittle. Arm-acquiring bandits. Annals of Probability, 9 (2):284–292, April 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994469.

Wierman:1982:PT

[Wie82] John C. Wierman. Percolation theory. Annals of Probability, 10 (3):509–524, August 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993764.

Williams:1985:RCI

[Wil85] R. J. Williams. Recurrence classification and invariant measure for reflected Brownian motion in a wedge. Annals of Probability, 13 (3):758–778, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176992907.

Woodroofe:1987:AEB

[WK87] Michael Woodroofe and Robert Keener. Asymptotic expansions in boundary crossing problems. Annals of Probability, 15(1):102–114, January 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992258.

Wood:1983:BET

[Woo83] Thomas E. Wood. A Berry–Esseen theorem for associated random variables. Annals of Probability, 11(4):1042–1047, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993453.

Wood:1985:LLT

[Woo85] Thomas E. Wood. A local limit theorem for associated sequences. Annals of Probability, 13(2):625–629, May 1985. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993015.

Wright:1980:TM

[Wri80] A. Larry Wright. On a theorem of Maruyama. Annals of Probability, 8(4):851–852, August 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994671.

Wright:1981:EDL

[Wri81] F. T. Wright. The empirical discrepancy over lower layers and a related law of large numbers. Annals of Probability, 9(2):323–329, April 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994475.

Watts:1982:LDI

[WRL82] Vernon Watts, Holger Rootzen, and M. R. Leadbetter. On limiting distributions of intermediate order statistics from stationary sequences. Annals of Probability, 10(3):653–662, August 1982. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993774.

Wu:1982:SLT

[Wu82] Rong Wu. Some limit theorems on reversed Brownian motion. Annals of Probability, 10(4):1079–1084, November 1982. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993732.

Washburn:1981:OSS

[WW81] Robert B. Washburn and Alan S. Willsky. Optional sampling of submartingales indexed by partially ordered sets. Annals of Probability, 9(6):957–970, December 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994267.

Yamazato:1982:SUI

[Yam82] Makoto Yamazato. On strongly unimodal infinitely divisible distributions. Annals of Probability, 10(3):589–601, August 1982. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993769.

Yamato:1984:CFM

[Yam84] Hajime Yamato. Characteristic functions of means of distributions chosen from a Dirichlet process. Annals of Probability, 12 (1):262–267, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176993389.

Yamada:1985:LTN

[Yam85] Keigo Yamada. A limit theorem for nonnegative additive functionals of storage processes. Annals of Probability, 13(2):397–413, May 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992999.

Yanagihara:1986:SDM

[Yan86] Hiroshi Yanagihara. Stochastic determination of moduli of annular regions and tori. Annals of Probability, 14(4):1404–1410, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992381.

Yang:1988:GGW

[Yan88] Wei-Shih Yang. The generators of a Gaussian wave associated with the free Markov field. Annals of Probability, 16(2):752–763, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991785.

Yanev:1985:CBP

[YM85] N. M. Yanev and K. V. Mitov. Critical branching processes with nonhomogeneous migration. Annals of Probability, 13(3):923–933, August 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992914.

Yu:1981:NTW

[Yu81] Kai Fun Yu. Note on a theorem of Woodroofe’s. Annals of Probability, 9(6):1030–1033, December 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994273.

Zabell:1980:RCC

[Zab80] Sandy L. Zabell. Rates of convergence for conditional expectations. Annals of Probability, 8(5):928–941, October 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994622.

Zachary:1983:CSS

[Zac83] Stan Zachary. Countable state space Markov random fields and Markov chains on trees. Annals of Probability, 11(4):894–903, November 1983. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993439.

Zahle:1988:RCC

[Zah88] M. Zahle. Random Cell complexes and generalised sets. Annals of Probability, 16(4):1742–1766, October 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991595.

Zakai:1981:SCT

[Zak81] Moshe Zakai. Some classes of two-parameter martingales. Annals of Probability, 9(2):255–265, April 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994466.

Zakai:1985:MDD

[Zak85] Moshe Zakai. Malliavin derivatives and derivatives of functionals of the Wiener process with respect to a scale parameter. Annals of Probability, 13(2):609–615, May 1985. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993013.

Zaman:1983:SFS

[Zam83] Arif Zaman. Stationarity on finite strings and shift register sequences. Annals of Probability, 11(3):678–684, August 1983. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176993512.

Zaman:1984:NCP

[Zam84a] Arif Zaman. A non-clustering property of stationary sequences. Annals of Probability, 12(1):193–203, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176993382.

Zaman:1984:UMM

[Zam84b] Arif Zaman. Urn models for Markov exchangeability. Annals of Probability, 12(1):223–229, February 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176993385.

Zaman:1986:FFF

[Zam86] Arif Zaman. A finite form of De Finetti’s theorem for stationary Markov exchangeability. Annals of Probability, 14(4):1418–1427, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176992383.

Zeephongsekul:1981:LFA

[Zee81] P. Zeephongsekul. Laplace functional approach to point processes occurring in a traffic model. Annals of Probability, 9(6): 1034–1040, December 1981. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994274.

Zeitouni:1989:OMF

[Zei89] Ofer Zeitouni. On the Onsager–Machlup functional of diffusion processes around non C^{2} curves. Annals of Probability, 17(3):1037– 1054, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991255.

Zhang:1986:LLN

[Zha86] Cun-Hui Zhang. The lower limit of a normalized random walk. Annals of Probability, 14(2):560–581, April 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176992531.

Zhang:1988:NRT

[Zha88] Cun-Hui Zhang. A nonlinear renewal theory. Annals of Probability, 16(2):793–824, April 1988. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991788.

Zhang:1989:RTV

[Zha89] Cun-Hui Zhang. A renewal theory with varying drift. Annals of Probability, 17(2):723–736, April 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176991423.

Zhang:1984:LTF

[ZZ84] Yu Zhang and Yi Ci Zhang. A limit theorem for N_{0n}/n in first-passage percolation. Annals of Probability, 12(4):1068–1076, November 1984. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993142.